### v1

# CATArena: Evaluation of LLM Agents Through Iterative Tournament Competitions

Abstract
Large Language Model (LLM) agents have evolved from basic text generation to autonomously completing complex tasks through interaction with external tools. However, current benchmarks mainly assess end-to-end performance in fixed scenarios, restricting evaluation to specific skills and suffering from score saturation and growing dependence on expert annotation as agent capabilities improve. In this work, we emphasize the importance of learning ability, including both self-improvement and peer-learning, as a core driver for agent evolution toward human-level intelligence. We propose an iterative, competitive peer-learning framework, which allows agents to refine and optimize their strategies through repeated interactions and feedback, thereby systematically evaluating their learning capabilities. To address the score saturation issue in current benchmarks, we introduce CATArena, a tournament-style evaluation platform featuring four diverse board and card games with open-ended scoring. By providing tasks without explicit upper score limits, CATArena enables continuous and dynamic evaluation of rapidly advancing agent capabilities. Experimental results and analyses involving both minimal and commercial code agents demonstrate that CATArena provides reliable, stable, and scalable benchmarking for core agent abilities, particularly learning ability and strategy coding.
[…]
1 Introduction
With the rapid evolution of agents powered by large language models (LLMs), their capabilities have far surpassed simple text generation. By actively invoking external tools, LLM agents have significantly expanded the boundaries of artificial general intelligence (AGI). These agents are now able to autonomously complete complex, multi-step tasks that are previously considered beyond their reach, such as developing software project (Manish, 2024; Hu et al., 2025b), intelligently performing strategic planning (Belle et al., 2025), and learning user preference (Gao et al., 2024).
Existing benchmarks mainly focus on end-to-end performance in specific tasks, such as code generation (Yang et al., 2024), AI research (Nathani et al., 2025), and GUI automation (Wang et al., 2024). These benchmarks provide detailed observations and analyses of LLM agents’ abilities within particular scenarios and have driven significant progress in the field. However, there are important limitations to these approaches. First, the scores obtained in these end-to-end benchmarks only reflect performance on specific tasks, whereas an agent’s overall capability is composed of multiple fundamental skills working together. Second, the absolute scores in these benchmarks, which are typically based on objective correctness, have an upper bound. As agents become increasingly powerful, maintaining and updating these benchmarks requires additional expert-level annotation, and the level of required expertise continues to rise. In light of these challenges, there is an urgent need for a quantifiable and continuously evolving benchmark that systematically measures and analyzes the fundamental sub-abilities of agents.
Previous research has shown that self-learning is an essential ability for agents to achieve human-level intelligence (Gao et al., 2025; Zhu et al., 2025). Beyond self-learning, agents, similar to humans, also engage in peer learning, which enables collective evolution through interactions and shared experiences (Liu et al., 2024). During this evolutionary process, agents receive feedback from their environment and continually improve themselves. This capacity for learning and adaptation is indispensable for LLM agents, as it prepares them for ongoing evolution and more complex challenges. To systematically evaluate this crucial ability, we propose an iterative peer-learning-based competitive framework for LLM agents. In each iteration, agents are required to revise and update their strategies based on the outcomes and policies observed in previous rounds of competition. After every update, the agent policy codes are executed and competed against each other, generating dynamic performance rankings. Through this peer-learning architecture, we gain valuable insights into the learning abilities of LLM agents.
Building on this peer-learning framework, we introduce CATArena (Code Agent Tournament Arena) [^1] Code of CATArena is available in https://github.com/AGI-Eval-Official/CATArena., which utilizes four open-ended, rankable games. These games, including both board games and card games, provide LLM agents with a peer-learning environment and unlimited upper bound for improvement. They enable agents to continually improve and compete, ensuring that the evaluation framework remains challenging as agent capabilities grow. Furthermore, our competitive arena is inherently extensible and can be readily adapted to other types of open-ended, rankable tasks, facilitating the assessment of core agent abilities in new domains. As agent capabilities continue to advance, CATArena can evolve by incorporating tasks with greater complexity and discrimination, thereby supporting ongoing evaluation without the need for expert-level human annotation.
In our experiments, we conduct comparative performance evaluations and data analysis conducted on our self-developed minimal code agent and state-of-the-art commercial code agents. CATArena consistently provides stable and reliable benchmarks for assessing both agent capabilities and the agentic potential of the underlying LLMs. Within the peer-learning framework, we design general scoring metrics to systematically assess the fundamental abilities of participating agents, including their learning ability. Our experiments demonstrate that the strategy coding tasks applied in CATArena are fundamentally different from traditional LLM reasoning tasks. This represents a novel evaluation dimension that has not been addressed in previous work. Additionally, we analyze characteristics of CATArena, demonstrating its reliability and extensibility as a benchmarking platform.
In summary, our contributions are as follows:
  • Iterative Peer-learning-based Competitive Framework: We propose a novel framework that leverages iterative peer-learning and competition to evaluate the learning abilities of LLM agents. Agents continuously revise their strategies based on feedback and outcomes from previous rounds, aligning agent evolution with human evolution.
  • CATArena Benchmark: We introduce CATArena, a tournament-style benchmark for evaluating the basic capabilities of LLM agents using a diverse set of open-ended games, including board and card games. CATArena provides an unlimited upper bound for agent improvement and supports extensible evaluation across diverse, open-ended tasks.
  • Comprehensive Agent Evaluation: We design general and systematic evaluation matrices and conduct comparative experiments and analyses between our minimal code agent and state-of-the-art commercial agents, demonstrating the reliability, stability, and extensibility of the CATArena.
2 Related Work
Learning Ability.
Learning ability is crucial for LLM agents, as it enables continual adaptation and improvement in dynamic and complex environments. Recent studies have shown that self-learning methods, such as self-refinement (Madaan et al., 2023; Shinn et al., 2023), allow models to enhance their outputs through iterative feedback, while environmental feedback further supports continual learning (You et al., 2024). In addition to self-learning, peer-learning has also been increasingly recognized, with approaches encouraging agents to learn from others’ reasoning processes and shared experiences (Liang et al., 2024; Luo et al., 2025). These diverse learning mechanisms have led to notable advances in tasks such as code generation, complex reasoning, and collaborative problem-solving. In the context of LLM-driven agents, learning ability represents a critical capability that supports effective adaptation and enables agents to tackle increasingly complex tasks and evolving challenges (Zhu et al., 2025; Gao et al., 2025). Despite the progress, systematic evaluation of how agents learn from each other remains underexplored, highlighting the need for benchmarks that capture both self-learning and peer-learning abilities.
Evaluation on Agents.
Recent benchmarks primarily assess LLM-driven agents on end-to-end, task-specific abilities. Code-based evaluations such as GitTaskBench (Ni et al., 2025), SUPER (Bogin et al., 2024), ProjectEval (Liu et al., 2025), SWE-PolyBench (Rashid et al., 2025), RedCode (Guo et al., 2024), SWT-Bench (Mündler et al., 2025), InfiAgent-DABench (Hu et al., 2024), and DA-Code (Huang et al., 2024) focus on large-scale software development, code security, bug fixing, and data science tasks. Other works extend agent evaluation to research (Du et al., 2025), real-world tool use (Yao et al., 2024), and assistant scenarios (Mialon et al., 2024). While some benchmarks explore agent-vs-agent evaluation (Zhuge et al., 2024), most rely heavily on human annotation and objective correctness, leading to upper bounds and saturation as agent capabilities advance.
Open-ended Tasks.
To address these limitations, recent work has leveraged open-ended, rankable tasks. Benchmarks such as GameBench (Costarelli et al., 2024), lmgame-Bench (Hu et al., 2025a), GAMEBot (Lin et al., 2024), and card game evaluations (Wang et al., 2025) assess LLMs’ strategic reasoning through diverse games. Frameworks like Game Reasoning Arena (Cipolina-Kun et al., 2025), GVGAI-LLM (Li et al., 2025), ZeroSumEval (Alyahya et al., 2025), TextArena (Guertler et al., 2025), and MCU (Zheng et al., 2025) further extend evaluation to multi-turn reasoning, spatial adaptability, natural language interaction, and open-ended tasks. However, these benchmarks mainly focus on reasoning skills and do not systematically evaluate agents’ learning abilities or coding strategies. Our analysis shows that measuring learning ability and coding strategy is fundamentally distinct, and both are essential for advancing agent intelligence. Additionally, in human board game competitions, variant rules such as Chess960 (FIDE, 2023) and Six-plus Hold’em (sixplusholdem, 2025) are often introduced to reduce memorization and encourage creativity. Notably, these variant rules have received relatively less attention in model evaluation.
3 CATArena
Figure 1: Overview of the evaluation framework and CATArena. The evaluation framework adopts an iterative peer-learning based competitive process. In round 1, LLM agents develop initial strategies via coding. These strategies are matched in a tournament arena, producing rankings and logs. In each subsequent round, agents analyze previous codes and logs, refine their strategies, and compete again. CATArena includes four open-ended and rankable games to cover diverse settings. Based on the tournament results from all rounds, a carefully designed scoring matrix and evaluation scheme are used to robustly quantify various agent abilities.
3.1 Iterative Peer-Learning based Competitive Framework
As shown in Figure 1, we propose an iterative peer-learning-based competitive framework, where CATArena evaluates code agents through a two-phase workflow: initial strategy development and iterative improvement. The initial phase assesses each agent’s ability to independently implement a baseline strategy based on the game code, while the iterative phase focuses on the agent’s learning ability.
Initial Development (Round 1). In this stage, each agent receives the game code and a sample AI implementation. Without external guidance, each agent must develop its own strategy to participate in the tournament. This phase primarily examines the agent’s strategy coding ability and establishes a baseline for subsequent evaluation.
Iterative Improvement (Rounds $n>1$). After the first round, all strategies submitted are evaluated through a tournament (round-robin format for symmetric games, batch-based competition for asymmetric games). Comprehensive competition logs are generated, recording rankings, win counts, and move histories for all matches. In subsequent rounds, agents are provided with the game code, previous round submissions from all participants, and these detailed logs. Agents must analyze these resources from previous rounds to adapt and improve their own strategies. This phase assesses the agent’s learning ability through repeated cycles of analysis and refinement.
This iterative evaluation framework of CATArena enables a granular assessment of both basic coding skills and advanced learning capabilities, supporting a robust and scalable measurement of code agent performance.
Table 1: Overview of game arenas and representative variants in CATArena.
Game Symmetry Type Players Variant
Gomoku Board 22 Forbidden points; dual three-in-a-row
Texas Hold’em Card $\geq 8$ Card removal; swapped hand ranks
Chess Board 22 Chess960; forbidden/special moves
Bridge Card 44 Card exchange
    For Bridge, symmetry is defined by assigning identical agent strategies to both teammates.
3.2 Games and Variants
Building on the tournament-based evaluation framework, CATArena deploys four distinct game arenas, each selected to test the strategic reasoning and coding capabilities of code agents across varying levels of complexity and interaction patterns. These arenas include competitive and cooperative settings, as well as symmetric and asymmetric game structures, thus enabling a diverse analysis of agents’ strategy coding abilities and learning patterns.
In addition to standard rules, each game is extended with thoughtfully designed variants that introduce novel or altered mechanics, inspired by real-world adaptations such as Fischer Random Chess (Chess960) (FIDE, 2023). Like human competition, the variant rules encourage strategy generalization and penalize rote memorization, as most models are trained on card and board game data. Table 1 provides an overview of the selected games and their respective variants.
3.3 Tournament Format and Scoring System
After the completion of all $N$ development rounds, CATArena conducts a comprehensive tournament to quantitatively evaluate agent strategies and compute performance metrics. A total of $T$ agent models participate, each contributing strategies in every round of development. Tournament formats are tailored to game types: for symmetric games, all strategies engage in a round-robin cycle, ensuring exhaustive pairwise competition; for asymmetric games such as Texas Hold’em, strategies are grouped into batches and compete in multi-agent matches. To mitigate randomness, all matches are repeated multiple times, and results are averaged for robust evaluation.
Scores are recorded in a scoring matrix $W\in\mathbb{R}^{(TN)\times(TN)}$, where $W_{i,j}^{n,m}\in[0,1]$ denotes the score obtained by agent $i$’s strategy in round $n$ against agent $j$’s strategy in round $m$. When $n=m$, the notation simplifies to $W_{i,j}^{n}$; similarly, when $i=j$, it is denoted as $W_{i}^{m,n}$. This scoring system enables fine-grained, quantitative analysis of agent performance in both individual and iterative development stages. For asymmetric games, pairwise results are not feasible; instead, batch-based tournaments are used and the score matrix records the win rates of multi-agent matches. The tournament format is provided in the Appendix A.
3.4 Evaluation Metrics
Based on the scoring matrix $W$, we design a set of evaluation metrics to quantitatively assess the key capabilities of code agents. Specifically, our metrics are constructed to measure three core capabilities: strategy coding, learning, and generalizability. In the following sections, we define these metrics using symmetric games as examples. For asymmetric games, the evaluation principles remain consistent. The calculation of the scoring matrix $W$ is provided in Appendix B.
Strategy Coding.
Strategy coding measures the agent’s fundamental ability to abstract game strategies into reproducible algorithms and implement them as executable code, which is fundamentally different from general reasoning and strategic planning abilities. In CATArena, this metric evaluates how effectively an agent can independently develop a baseline strategy for the game environment and compete against other agents in the initial development stage.
For each agent $i$, strategy coding is quantified by the average score obtained against all other agents in the first round:
$S_{i}=\mathrm{avg}_{j\neq i}(W_{i,j}^{1}).$
This metric serves as the foundational benchmark for code agent evaluation in CATArena.
Learning Ability.
The learning capability of a code agent captures its ability to leverage historical information and opponent behaviors to improve its own performance.
Global Learning assesses an agent’s overall improvement in strategy quality. This metric evaluates the relative performance of agent $i$’s strategies against all strategies from all agents and rounds, and measures the average progress made compared to its initial baseline. It serves as the primary indicator of learning ability.
Formally, for agent $i$, global learning is defined as:
$L_{i}=\mathrm{average}_{n=2}^{N}\left(G_{i}^{n}-G_{i}^{1}\right),$
where $G_{i}^{n}$ represents the global performance of agent $i$’s strategy from round $n$:
$G_{i}^{n}=\mathrm{average}_{(i,n)\not=(j,m)}\left(W_{i,j}^{n,m}\right).$
This metric captures the agent’s ability to learn and adapt over multiple rounds, reflecting its progress in a comprehensive competitive landscape.
Counter-Adaptation measures an agent’s targeted learning ability, reflecting its capacity to achieve improved results against opponents in successive rounds. For agent $i$, the counter-adaptation score is defined as the average improvement in scores against other agents from round n1n-1 to round $n$ $(n\geq 2)$:
$C_{i}=\mathrm{average}_{n=2}^{N}\left(A_{i}^{n}-B_{i}^{n-1}\right),$
where the advance score $A_{i}^{n}$ and base score $B_{i}^{n-1}$ are defined as:
$A_{i}^{n}=\mathrm{average}_{j\neq i}\left(W_{i,j}^{n,n-1}\right),B_{i}^{n-1}=\mathrm{average}_{j\neq i}\left(W_{i,j}^{n-1}\right).$
Here, $A_{i}^{n}$ represents agent $i$’s average performance in round $n$ against the strategies submitted by other agents in the previous in round (n1n-1). The base score $B_{i}^{n-1}$ denotes agent $i$’s average performance in round n1n-1 against those same opponents. This comparison isolates the agent’s targeted adaptation from one round to the next.
Self-improvement evaluates an agent’s capacity to genuinely enhance its strategies over successive rounds of development. This metric reflects whether newly developed strategies can consistently outperform the agent’s own previous versions.
We quantify self-improvement by calculating the Pearson correlation Pearson (1896) between the round index and the agent’s average scores across rounds. For agent $i$, the self-improvement score is defined as
$\mathrm{SI}_{i}=\mathrm{Pearson}\left([1,\cdots,N],\ [S_{i}^{1},\cdots,S_{i}^{N}]\right).$
Here, $S_{i}^{n}$ denotes the average score of agent $i$’s strategy from round $n$ against its own strategies from other rounds
$S_{i}^{n}=\mathrm{average}_{m\neq n}\left(W_{i}^{n,m}\right).$
A higher self-improvement score indicates a stronger ability to iteratively refine and upgrade strategies throughout the development process.
Generalizability.
Generalizability measures an agent’s ability to comprehend and adapt to novel or altered game rules that differ from those encountered during training or prior experience. This metric specifically evaluates the agent’s capacity to generalize beyond previously seen environments, focusing on handling new or modified scenarios. For agent $i$, the generalizability score is defined as:
$U_{i}=B_{i}^{1;\mathrm{Variants}}-B_{i}^{1;\mathrm{Standard}},$
where $B_{i}^{1;\mathrm{Variants}}$ and $B_{i}^{1;\mathrm{Standard}}$ denote the base scores of agent $i$ in the first round under variant and standard rule settings, respectively. A higher value of $U_{i}$ indicates stronger generalizability, reflecting the agent’s ability to effectively develop and apply strategies for previously unseen tasks.
4 Experiments
4.1 Experimental Setups
Participants.
In our experiment, we employ three types of agents: (1) Minimal Agents (LLM + ADK Framework) [^2] https://github.com/AGI-Eval-Official/Minimal-CodeAgent: A baseline agent developed with the Agent Development Kit (ADK) Python toolkit. We provide essential tools, including file manipulation, bash scripting, and Python execution, for the ADK code agent. On this foundation, we integrate state-of-the-art LLMs to systematically compare their core competencies as code agents for strategy implementation. (2) Commercial Code Agents: State-of-the-art and commercial CLI-based agents (e.g., Claude Code, CodeX, Gemini-CLI, Qwen-Coder) are included for benchmarking. These agents feature advanced integration with various command-line interfaces, tools, and LLMs, resulting in enhanced overall capabilities. These agents serve as leading solutions in code agent development and provide valuable reference points for future research. (3) LLM-Player: In this control setting, LLMs directly output game moves without generating code. For each turn, the LLM receives the game rules, current state, and history, and returns the next action. This approach is specifically designed to assess the inherent strategic and reasoning capabilities of LLMs. Detailed agent parameter settings and model selections are presented in Appendix C.
Table 2: Agent Specifications and Open-Source Status.
Agent Type Agent Framework Model Agent OSS LLM OSS
Minimal basic code tools with ADK framework DeepSeek-3.1 (DeepSeek-AI, 2024)
Qwen3-Coder-480B (Team, 2025c)
Doubao-Seed-1.6 (Team, 2025d)
GPT-5 (OpenAI, 2025b)
Claude-4-Sonnet (Anthropic, 2025a)
Gemini-2.5-pro (Team, 2025a)
Commercial Gemini-CLI (Google, 2025) Gemini-2.5-pro (Team, 2025a)
Claude-Code (Anthropic, 2025b) Claude-4/3.7 Hybrid (Anthropic, 2025a)
CodeX (OpenAI, 2025a) GPT-5 (OpenAI, 2025b)
Qwen-Coder (Team, 2025b) Qwen3-Coder-480B (Team, 2025c)
Other LLM-Player Agents’s Corresponding LLM N/A N/A
Tournaments.
All experiments are conducted under two main tournament settings: (1) a comparison among minimal agents equipped with different LLMs ($T_{1}=6$), and (2) a comparison between the best-performing minimal agent and a set of commercial code agents ($T_{2}=5$). To reduce the impact of randomness on strategy generation, each tournament is repeated for four times, and all reported metrics are averaged over the four runs. Each tournament consists of $N=4$ rounds of iterative development. To further mitigate stochastic effects in competition outcomes, every entry in the scoring matrix $W$ is estimated by repeated matches. Detailed tournament prompts are listed in Appendix K.
We report the detailed scoring policy, generation configs, and repetition experiments in Appendix D. It is noteworthy that agents tend to generate different codes in repeated experiments, but their rankings are relatively stable.
4.2 Main Results
(a) Trends of global performance scores $G_{i}^{n}$.
(b) Quantitative comparison of learning abilities.
Figure 2: Visualization of agents’ learning patterns and scores.. For clarity, we use family names to represent LLM models instead of their full names. The results for other games are in Appendix E.
Table 3: Main LeaderBoard of CATArena. We conduct two groups of tournaments between minimal agents (rank from 1 to 6) and commercial agents (rank from 1 to 5), and report the average ranking across all four tasks. Metrics include S.C. (Strategy Coding), G.L. (Global Learning), and G.A. (Generalizability).
Agents Standard Variant G.A. $\downarrow$
S.C. $\downarrow$ G.L. $\downarrow$ S.C. $\downarrow$ G.L. $\downarrow$
Minimal
Claude-4-Sonnet 1.25 2.5 1.75 2.75 5.00
DeepSeek-Chat 5.75 2.75 4.25 2.75 2.75
Doubao-Seed 3.75 4.75 3.75 4.50 2.75
Gemini-2.5-Pro 3.25 3.75 3.25 2.75 3.25
GPT-5 3.75 3.50 3.00 3.75 2.25
Qwen3-Coder 2.25 3.75 3.00 4.5 4.75
Commercial
best ADK 3.25 2.25 2.00 3.75 2.50
Claude-Code 2.50 3.75 2.50 2.75 3.25
CodeX 2.25 2.75 3.00 3.00 3.25
Gemini-CLI 3.50 2.25 3.00 4.00 2.00
Qwen-Coder 3.00 3.75 4.00 1.25 3.25
Learning Ability.
Figure 2(a) visualizes the global performance scores $G_{i}^{n}$ revealing overall trends in agent strategies across iterations. Some agents, such as minimal agents driven by claude, exhibit a clear upward trajectory over multiple rounds, demonstrating strong learning capability. However, the performance of most agents remains unstable, and no obvious trend is observed. To explicitly illustrate the learning ability of each agent, we design a quantitative scoring method for global learning and introduce two additional learning modes: counter-adaption and self-improvement. The quantitative results for these three abilities are presented in Figure 2(b). Compared to Figure 2(a), these scores offer a more intuitive comparison of the agents’ learning strengths. This decomposition provides deeper insights into the mechanisms underlying agent learning. When both counter-adaption and self-improvement scores are positive, it indicates that the agent can effectively learn from both its opponents and itself, resulting in a positive global learning score. Compared to minimal agents, commercial agents exhibit stronger learning capabilities. Learning ability results for other games and a case study of strategy iteration mechanisms are provided in Appendix E. Through an analysis of the consistency of agent actions in endgame states, we observe that a majority of agents indeed learn from the code generated in the previous round, leading to increasing consistency. This phenomenon is most prominent in the Hold’em environment, possibly because its strategies are relatively simple and easier to learn.
Ability Evaluation.
We conduct two sets of tournaments: (1) minimal agents equipped with different LLMs, (2) the best-performing minimal agent against commercial code agents. For each tournament, we report the average ranking and scores for three core agent capabilities. The main leaderboard and main results of CATArena are summarized in Table 3 and Table 4. These results reveal the following key observations:
Observation 1: The performance gap among LLMs is more pronounced in minimal agents compared to commercial agents. Table 3 shows that Claude-4-Sonnet achieves the highest score among minimal agents, while the rankings of other LLMs are more dispersed. In contrast, commercial agents driven by the same LLMs exhibit much closer average rankings, with all agents scoring around 2.5 out of 5, indicating a reduced performance gap. Moreover, commercial agents demonstrate performance levels similar to the best-performing minimal agent. This suggests that the underlying agent framework can significantly influence how effectively an LLM’s capabilities are utilized, as commercial agents are often optimized for specific models.
Observation 2: The participating agents display different ranking orders across various capabilities. The tournament results reveal that the relative rankings of agents change depending on the specific core ability being tested. The ranking of these abilities provides a decomposition of end-to-end performance, offering insights for further optimization of both LLMs and agent frameworks.
Observation 3: Agents exhibit varied performance distributions across different tasks. The results indicate that agents’ performance is not uniformly distributed across all tasks, which is mainly attributed to the distinct nature and difficulty of the four tasks. In the variant tasks, the performance gap among agents is more pronounced, likely because the game rules and strategies are less familiar to agents.
[…]
Games Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant
[…]
Minimal
Claude-4-Sonnet 0.88 -0.447 0.78 -0.156 -0.14 0.58 0.118 0.13 0.110 -0.45 0.79 0.174 1.0 0.047 0.005 0.90 -0.170 0.65 0.018 -0.55
Deepseek-Chat 0.23 0.027 0.38 0.077 0.13 0.01 0.010 0.00 -0.022 -0.01 0 0 0.10 0 0.10 0 0 0.10 0.049 0.10
Doubao-Seed 0.33 -0.192 0.72 -0.302 0.46 0.04 -0.035 0 0 -0.04 0.2 -0.033 0.45 -0.516 0.40 0.58 -0.337 0.10 0.034 -0.46
Gemini-2.5-Pro 0.25 -0.066 0.00 0.173 -0.12 0.01 0.020 0.00 0.078 -0.01 0.90 -0.195 0.60 -0.049 -0.30 0.58 -0.147 0.90 0.003 0.46
GPT-5 0.48 0.062 0.76 -0.019 0.18 0.16 0.102 0.87 -0.050 0.71 0.47 -0.095 0.10 0.293 -0.02 0.38 -0.525 0.45 -0 0.24
Qwen3-Coder 0.85 -0.523 0.36 -0.089 -0.50 0.20 0.038 0.00 0.003 -0.20 0.65 0.032 0.76 -0.230 -0.19 0.58 -0.187 0.80 -0.532 0.21
Commercial
best ADK 0 0.075 0.75 -0.022 0.88 0.07 0.073 0.46 0.030 0.39 0.25 0.295 0 0.361 -0.25 0.91 -0.110 1.00 -0.342 -0.47
Claude-Code 0.78 -0.322 0.66 0.194 -0.28 0.01 0.100 0 0.105 -0.001 1.00 -0.240 0.93 -0.139 -0.04 0.56 -0.158 0.44 -0.226 0.03
CodeX 0.47 0.454 0.69 -0.095 0.34 0.72 0.050 0.17 0.067 -0.55 0.75 0.098 0.50 0.285 -0.25 0.38 0.033 0.34 0.064 0.09
Gemini-CLI 0.31 0.260 0.19 0.172 0.0 0.13 0.050 0.37 -0.007 0.24 0.01 0.254 0.83 -0.286 0.79 0.38 0.395 0.38 -0.154 0.16
Qwen-Coder 0.94 -0.054 0.22 -0.530 -0.94 0.07 0.058 0 0.105 -0.007 0.49 0.157 0.25 0.556 -0.25 0.28 -0.204 0.34 0.039 0.19
4.3 Effectiveness of CATArena
We design a series of experiments to demonstrate the effectiveness of CATArena.
Comparison between Agents and LLM-Players.
The primary task in CATArena is strategy coding, which relies on the underlying coding capabilities of LLMs. We posit that reasoning over code to develop strategies is fundamentally different from direct reasoning during gameplay. To validate this distinction, we compare agent-developed strategies with the LLM-Player baseline (see Appendix F). Our results show that current agents primarily implement simple rule-based algorithms, indicating substantial room for advancement in agents’ strategy coding abilities. CATArena fully leverages this non-saturation, enabling sustainable iterative peer-learning.
To further analyze the similarities and differences between agent-implemented code strategies and those of LLM-Players, we ask agents’ code and LLM-Player to select the next action on endgame states. Figure 3 illustrates the action consistency between agents’ code and LLM-Players in Chess. Surprisingly, the strategies encoded in agent code differ significantly from those inferred directly by the LLM, even if they are from the same model. Meanwhile, strategies produced by different agents and different LLMs also show notable similarities. This indicates that strategy coding and reasoning in LLMs are distinct capabilities. We report results of other tasks are demonstrated in Appendix G. CATArena evaluates the strategy coding ability of agents rather than their reasoning ability, thereby filling a gap in previous benchmarks. The relationship between strategy coding ability and LLM-based strategy reasoning remains unclear and requires further investigation.
Figure 3: Action consistency between agents’ code and LLM-Players on Chess endgames.
Collective Learning Trends Among Agents.
Table 5: Collective learning trends of agents across different tasks in CATArena. $\text{DIS}_{\text{range}}$ and $\text{DIS}_{\text{std}}$ represent the Pearson correlation coefficients of the range and standard deviation of agent performance scores over four rounds, reflecting the similarity and dispersion of agent strategies. $\text{Trend}_{\text{mean}}$ denotes the Pearson correlation between the mean agent performance and the round number, indicating the overall trend of group improvement.
Gomoku Hold’em Bridge Chess
StdV. VarV. StdV. VarV. StdV. VarV. StdV. VarV.
$\text{DIS}_{\text{std}}$ -0.05 0.15 -0.81 -0.80 -0.82 -0.57 0.55 -0.04
$\text{DIS}_{\text{range}}$ -0.16 0.44 -0.80 -0.76 -0.54 -0.33 -0.08 0.16
$\text{Trend}_{\text{mean}}$ 0.42 -0.02 0.75 0.67 0.24 -0.10 -0.74 -0.79
We analyze the collective learning dynamics of agents across tasks, as presented in Table 5. The metrics $\text{DIS}_{\text{range}}$ and $\text{DIS}_{\text{std}}$ report the Pearson correlation between the standard deviation and range of agent performance scores over four rounds ($B_{i}^{n},n=1,2,3,4$). The higher similarity in performance scores (i.e., $\text{DIS}>0$) suggests that agents can learn effective strategies more readily, indicating lower task difficulty. Based on these results, the relative difficulty ranking of tasks in CATArena is Chess >> Gomoku >> Bridge >> Hold’em.
[…]
Additional Results.
CATArena’s iterative peer-learning framework is easily extensible to new tasks for evaluating other fundamental agent abilities. We demonstrate this by introducing a Machine Learning (ML) track and multi-lingual track, with experimental results provided in Appendix H and Appendix I, respectively. Experimental results indicate that current agents still exhibit substantial potential for improvement. As agents continue to advance, the open-ended task design and peer-learning evaluation framework of CATArena ensure that systematic assessment can be sustained over time. We also report agent cost in terms of token usage, time consumption, and generated code statistics for each agent, in Appendix J. Notably, Claude-4-Sonnet utilizes the most tools and tokens, and also develops a significantly larger amount of code. In contrast, GPT-5 achieves the best balance between token usage and performance. Efficient utilization of tokens and external tools remains an important research direction to advance the capabilities of LLM agents.
5 Conclusion
In this work, we address two fundamental challenges in LLM agent evaluation: the need for systematic measurement of learning ability, and the tendency of traditional benchmarks to become saturated as agent capabilities improve. To this end, we propose an iterative peer-learning-based competitive framework, enabling agents to continually revise and enhance their strategies through dynamic interaction and feedback. Building on this, we introduce CATArena, a tournament-style benchmark featuring open-ended and rankable board and card games. CATArena provides an environment with unlimited potential for agent improvement potential and extensible evaluation across new domains. Experimental results demonstrate that our framework reliably assesses core agent abilities, particularly learning ability and strategy coding, while ensuring stability and scalability. The open and flexible architecture of CATArena supports ongoing research and benchmarking for future intelligent agents.
Limitations.
The current evaluation in CATArena is limited to four games, which primarily assess agents’ learning ability and strategy coding. These scenarios do not encompass the full spectrum of potential LLM agent capabilities. In future work, we plan to introduce a wider variety of more complex tasks to evaluate agents’ learning and other abilities from different perspectives.
Appendix A Tournament Format and Scoring System
Table 6: Configs of tournament on four games.
Environment Code Agents LLM-player
Gomoku
Board size: $\times15$
Number of pairwise matches: 4 $\times$ 2
Swap black and white pieces after each match
Maximum time per move: 10 s
Board size: $\times15$
Number of pairwise matches: 2 $\times$ 2
Swap black and white pieces after each match
Maximum time per move: 600 s
Texas Hold’em
Max players: 12
Rounds: 100
Random shuffle seat after each round
Initial chips: 2000
Blind increase every 24 hands
Max hands per round: 720 or until winner decided
Maximum time per move: 3 s
Max players: 12
Rounds: 10
Random shuffle seat after each round
Initial chips: 2000
Blind increase every 24 hands
Max hands per round: 720 or until winner decided
Maximum time per move: 1000 s
Bridge
Number of pairwise matches: 12 $\times$ 2
Swap directions of open/closed rooms
Use same deck for each pair of match
Maximum time per move: 10 s
Number of pairwise matches: 12 $\times$ 2
Swap directions of open/closed rooms
Use same deck for each pair of match
Maximum time per move: 200 s
Chess
Number of pairwise matches: 8 $\times$ 2
Swap black and white pieces after each match
Maximum moves per game: 200
Maximum time per move: 10 s
Number of pairwise matches: 2 $\times$ 2
Swap black and white pieces after each match
Maximum moves per game: 200
Maximum time per move: 600 s
We list the basic settings for each games in Table 6.
For each game, we ensure that the number of pairwise matches among code agents allows the final results to stabilize, i.e., for each game, the L1-norm fluctuation of the scoring matrix $W$ is less than 5%.
For LLM-players, since their reasoning time is relatively long for most games, we reduce the number of repeated experiments. Note that our comparison with LLM-players is only a qualitative analysis of the differences between LLM-player and corresponding coding agent. The exploration of LLM-players’ results is not the focus of this paper; refer to the main text for details.
Appendix B Evaluation Metric Calculation
Table 7: Scoring rules of tournament on four games.
Environment Scoring Metric
Gomoku Pairwise match scoring: Win = 1 point,  Draw = 0.5 point,  Lose = 0 point.
Texas Hold’em Multi-agent batches: score is the average win rate across all tournaments participated.
Bridge 20 VP system: Two opposing pairs’ scores sum to 20, Final score divided by 20, ensuring each pair’s score $\in[0,1]$.
Chess Pairwise match scoring: Win = 1 point,  Draw = 0.5 point,  Lose = 0 point.
To evaluate the basic capabilities of LLM agents, we define metrics for each applied game. Let $N$ be the number of rounds and $K$ be the number of participating agents. We construct a matchup matrix
$W\in\mathbb{R}^{(N\cdot K)\times(N\cdot K)},$
where the generic element $W_{i,j}^{n,m}$ denotes the score when agent $i$ from round $n$ plays against agent $j$ from round $m$. We abbreviate $W^{n}_{i,j}$ for same-round comparisons ($n=m$) and $W_{i}^{n,m}$ for self-comparisons across rounds ($i=j$). Diagonal entries $(n,i)=(m,j)$ are ignored.
For asymmetric games, pairwise results are not feasible; instead, batch-based tournaments are used and the score matrix records the win rates of multi-agent matches. For each batch, we obtain a single score group $W_{i_{1},i_{2},\ldots,i_{BS}}^{n_{1},n_{2},\ldots,n_{BS}}$, where $BS$ is the batch size. We conduct three types of experiments to accommodate different metric calculations: (1) $W_{i}^{1,2,\ldots,N}$, where the same agent’s strategies from $N$ rounds compete against each other, used to compute self-improvement metrics $S_{i}^{n}$; (2) $W_{i_{1},i_{2},\ldots,i_{T}}^{n}$, where all agents in the same round compete, used to calculate the base score $B$; (3) All $N\times T$ agent strategies are randomly shuffled and grouped for competition (with $BS=12$ in our experiments), used to compute both the global score $G_{i}^{n}$ and advanced score $A_{i}^{n}$.
Scoring rules for the four games are summarized in Table 7.
Appendix C Generation Configs
For all LLMs used in our work, we set temperature to be $0.1$, max token identical to their official APIs’ setting. We set top-p to $1.0$, Top-k to be $100$, and presence penalty to be default to the API.
Additionally, both Claude-4-Sonnet and DeepSeek-3.1 occasionally encounter tool call issues that result in no code being generated, as frequently reported by the community. If such errors occur three times in a row, we substitute Claude-4-Sonnet with Claude-3.7-Sonnet and DeepSeek-3.1 with DeepSeek v3.
For LLM-players, considering the uncertainty in model output formats, we allow up to three retries. The prompt of LLM-players are in arena’s code and not present in paper considering its excessive length.
Appendix D Repetition Experiments
Table 8: Standard deviation of ranking in Round 1 and Round 2 with repeating 4 times.
Games Average Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
Round 1
Minimal
Claude-4-Sonnet 0.80 0.91 1.58 0.43 1.12 0.71 0.50 1.64 0.00 0.87
Deepseek-Chat 0.72 0.81 0.83 0.87 0.83 1.22 1.22 0.71 0.00 0.43
Doubao-Seed 1.58 0.90 1.87 1.87 1.09 0.43 2.06 0.87 1.30 0.43
Gemini-2.5-Pro 1.24 1.23 1.50 1.30 1.12 0.83 1.12 1.66 1.22 1.12
GPT-5 0.75 1.18 0.71 1.50 0.71 1.30 1.09 1.50 0.50 0.43
Qwen3-Coder 1.16 0.84 1.48 1.50 1.22 0.71 1.09 0.71 0.83 0.43
Commercial
Claude-Code 1.27 1.01 1.12 1.66 0.43 0.00 1.79 1.09 1.73 1.30
CodeX 0.76 0.57 0.83 0.50 0.43 0.50 1.09 1.30 0.71 0.00
Gemini-CLI 1.14 0.93 0.83 0.43 1.12 1.00 1.30 1.79 1.30 0.50
Qwen-Coder 1.01 0.95 1.22 1.48 1.12 0.00 0.87 0.83 0.83 1.50
Round 2
Minimal
Claude-4-Sonnet 0.81 0.55 1.48 0.50 0.83 0.43 0.50 0.83 0.43 0.43
Deepseek-Chat 0.94 1.03 1.30 1.09 0.87 0.83 1.09 1.48 0.50 0.71
Doubao-Seed 0.79 0.88 1.09 1.12 0.87 0.71 0.71 0.87 0.50 0.83
Gemini-2.5-Pro 1.30 1.28 1.66 1.92 1.00 1.22 2.06 1.12 0.50 0.87
GPT-5 0.89 1.22 1.00 1.58 0.43 1.66 1.64 1.22 0.50 0.43
Qwen3-Coder 0.82 1.02 1.58 1.30 0.83 0.43 0.43 1.50 0.43 0.83
Commercial
Claude-Code 1.20 1.04 1.09 1.41 1.22 0.43 1.64 1.09 0.83 1.22
CodeX 0.71 0.87 0.43 0.87 0.50 1.09 0.83 1.09 1.09 0.43
Gemini-CLI 0.88 1.10 0.83 1.64 0.83 0.83 1.41 1.48 0.43 0.43
Qwen-Coder 1.13 0.90 1.66 0.83 0.43 0.83 1.00 1.12 1.41 0.83
We report the results of repetition experiments($N=4$) on first two tournament round in Table 8.
From table, we observe that 1. The rankings of most agents remain relatively stable ,with ranking standard deviation changes of less than one. However, a few agents, such as Gemini-2.5-Pro and Claude-Code, exhibit greater fluctuations; 2. The rankings for standard games are more stable than those for variant games; 3. The results of the open source model are more stable than those of the closed source model, and commercial agents are more stable than minimal agents; 4. Additionally, we observe that agents do not consistently generate runnable code repositories across multiple development attempts. Even commercial agents occasionally fail to produce successful builds, which suggests that current code agents still need to improve their development stability.
Appendix E Learning Ability
E.1 Global Learning Trend
(a) Gomoku.
(b) Hold’em
(c) Bridge.
(d) Chess.
Figure 4: Trend of global performance score $G_{i}^{n}$ in Gomoku, Hold’em, Bridge and Chess.
As shown in Figure 4, we present the trends of global performance scores $G_{i}^{n}$ across four games, revealing distinct performance patterns for different models. In many cases, agents experience a sharp decline in performance during an intermediate round, which we interpret as a learning failure. Typically, such failures are recovered in the following round.
E.2 Detailed Learning Score
We list the detailed score of global learning, counter-adpatation learning and self-improvement in table 91011 respectively.
Table 9: Global Learning with Group-wise Average Rankings.
Models Avg. Ranking Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
Minimal
Claude-4-Sonnet 2.50 2.75 -0.447 -0.156 0.118 0.110 0.174 0.047 -0.170 0.018
Deepseek-Chat 2.75 2.75 0.027 0.077 0.010 -0.022 0.000 0.000 0.000 0.049
Doubao-Seed 4.75 4.50 -0.192 -0.302 -0.035 0.000 -0.033 -0.516 -0.337 0.034
Gemini-2.5-Pro 3.75 2.75 -0.066 0.173 0.020 0.078 -0.195 -0.049 -0.147 0.003
GPT-5 3.50 3.75 0.062 -0.019 0.102 -0.050 -0.095 0.293 -0.525 -0.000
Qwen3-Coder 3.75 4.50 -0.523 -0.089 0.038 0.003 0.032 -0.230 -0.187 -0.532
Commercial
best ADK 2.25 3.75 0.075 -0.022 0.073 0.030 0.295 0.361 -0.110 -0.342
Claude-Code 3.75 2.75 -0.322 0.194 0.100 0.105 -0.240 -0.139 -0.158 -0.226
CodeX 2.75 3.00 0.454 -0.095 0.050 0.067 0.098 0.285 0.033 0.064
Gemini-CLI 2.25 4.00 0.260 0.172 0.050 -0.007 0.254 -0.286 0.395 -0.154
Qwen-Coder 3.75 1.25 -0.054 0.536 0.058 0.105 0.157 0.556 -0.204 0.039
Table 10: Counter-adaptation Score with Group-wise Average Rankings.
Games Avg. Ranking Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
Minimal
Claude-4-Sonnet 3.75 3.75 -0.096 -0.075 0.001 -0.023 0.005 0.128 -0.042 -0.075
Deepseek-Chat 2.75 2.00 0.008 0.088 0.004 0.021 0.000 0.000 0.000 0.100
Doubao-Seed 4.25 4.50 0.063 -0.196 -0.023 -0.061 -0.008 -0.133 -0.196 0.083
Gemini-2.5-Pro 2.75 2.75 0.354 0.192 0.014 0.097 -0.132 -0.238 -0.038 0.021
GPT-5 4.75 3.25 0.038 -0.012 -0.086 0.019 -0.098 0.052 -0.154 -0.025
Qwen3-Coder 2.75 4.75 -0.092 0.029 0.000 -0.080 0.037 -0.008 0.025 -0.167
Commercial
best ADK 2.50 3.00 0.260 0.000 0.018 0.039 0.127 0.238 -0.094 -0.104
Claude-Code 3.25 1.88 -0.042 0.083 0.091 0.194 0.060 0.081 -0.104 0.047
CodeX 3.50 3.50 0.104 -0.089 -0.034 0.033 0.023 -0.090 0.010 0.052
Gemini-CLI 2.75 4.25 0.120 -0.021 0.031 -0.083 -0.027 -0.169 0.188 0.010
Qwen-Coder 3.00 2.38 -0.130 0.193 0.041 0.077 0.092 0.081 -0.062 -0.021
Table 11: Self-improvement Score with Group-wise Average Rankings.
Games Avg. Ranking Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
Minimal
Claude-4-Sonnet 2.75 2.50 -0.103 -0.894 0.949 0.517 0.858 0.766 -0.848 0.478
Deepseek-Chat 3.25 3.38 0.000 0.893 -0.949 -0.747 0.000 0.000 0.000 0.000
Doubao-Seed 3.75 4.38 0.141 -0.686 0.075 0.000 -0.202 -0.775 -0.894 0.000
Gemini-2.5-Pro 3.63 2.00 0.400 0.775 -0.758 0.894 -0.598 -0.240 -0.775 0.913
GPT-5 4.88 4.50 -0.897 -0.949 -0.205 -0.050 -0.767 0.485 -0.775 -0.390
Qwen3-Coder 2.75 4.25 -0.400 0.161 0.668 0.202 0.473 -0.400 -0.258 -0.775
Commercial
best ADK 2.25 3.50 0.400 0.956 0.835 -0.614 0.738 0.546 -0.207 -0.726
Claude-Code 4.00 2.25 -0.230 0.969 0.346 0.904 0.113 0.537 -0.730 -0.225
CodeX 2.00 3.25 0.763 -0.763 -0.090 0.602 0.784 0.316 0.424 0.316
Gemini-CLI 3.75 4.00 -0.183 -0.356 -0.176 0.000 0.000 -0.995 0.811 -0.193
Qwen-Coder 3.00 2.00 0.632 0.717 0.826 0.705 0.641 0.677 -0.944 0.000
The trend on four games are rather different. In general, the commercial model group consistently demonstrates superior global learning capability, where the advantage is particularly evident in complex strategy games like Chess or Gomoku variant.
Despite most minimal agents fail to learn well on complex games, we still find that Claude-4-Sonnet significantly surpass the rivals on standard games. However, the Claude-4-Sonnet still lacks behind on some cases like Gomoku, indicating that current LLMs agentic ablilities are still limited by the framework, where the commericial agents optimize workflow for their specific models to achieve the best results.
In simple games such as Hold’em, a larger proportion of agents exhibit positive learning scores, whereas in complex games like Chess, the prevalence of negative scores increases markedly. This trend suggests that current agents still face significant limitations in learning complex strategies.
E.3 Behavioral Changes Induced by Learning
For each game, we randomly select 80-100 intermediate states from the agents’ rival history and require the agent or LLM-player to choose the next move for each state. To ensure clarity in our writing, we uniformly refer to these intermediate states as endgame throughout the paper. Please note that endgame here is not limited to the final stages of the game; samples are taken from early, middle, and late stages as well.
We visualize the action consistency among the first two rounds in four games’s endgame in Figure 5.
From the matrix, we observe that, in general, agents tend to learn the strategies of other agents in the first round (lower left part vs. upper left part). Specifically, Doubao-Seed and DeepSeek-Chat simply copy Claude-4-Sonnet’s strategy in Holdem. Additionally, the learning trend is more pronounced in simpler games (Holdem vs. Chess).
For simpler games like Holdem, agent strategies in the second round are more similar to those in the first round (lower right part vs. upper left part), while for more difficult games like Chess, the trend is reversed. This observation is consistent with our findings on $\text{Trend}_{\text{mean}}$ in Table 5.
(a) Holdem.
(b) Gomoku.
(c) Bridge.
(d) Chess.
Figure 5: Action consistency between round 1 and round 2 agents’ code on endgames.
Appendix F Comparison between Agent and LLM-Player
Table 12: Comparison of match outcomes between each agent and its corresponding LLM-Player. Each value indicates the agent’s win rate when competing against the LLM that powers it.
Agent VS LLM Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant
Claude-4-Sonnet 1.00 1.00 0.00 0.00 0.45 1.00 0.88 0.75
Deepseek-Chat 0.50 0.00 0.40 0.30 0.00 0.00 0.00 0.00
Doubao-Seed 0.50 0.25 0.00 0.00 0.00 0.10 0.00 0.00
Gemini-2.5-Pro 0.00 0.25 0.00 0.00 1.00 1.00 0.13 0.50
GPT-5 0.00 0.50 0.00 0.50 1.00 0.00 0.00 0.00
Qwen3-Coder 1.00 0.00 0.50 0.50 1.00 0.00 0.50 0.50
Claude-Code 1.00 1.00 0.00 0.00 0.70 0.85 0.75 0.63
CodeX 0.00 0.50 0.20 0.40 1.00 0.38 1.00 1.00
Gemini-CLI 0.00 0.00 0.10 0.40 0.00 1.00 0.38 0.50
Qwen-Coder 1.00 1.00 0.50 0.00 1.00 0.00 0.56 0.63
We compare the strategies of agents’ code and its corresponding LLM-Player on four games in table 12.
Interestingly, there is no strong correlation between the performance of the agents’ code and that of their underlying models.
In games with strong strategic elements, such as Gomoku and Chess, some agents’ code significantly outperforms their corresponding LLM-Player, indicating that the code implementation is able to better leverage game rules and strategies. For example, the agent developed by claude-4-sonnet achieves a 100% win rate against its LLM-Player in both standard and variant Gomoku, and also demonstrates a high win rate in Chess and Bridge. This suggests that the strategies implemented in the code are superior to the large model’s direct reasoning performance as a player in these games. In contrast, the agents developed by doubao-seed and deepseek-chat struggle to defeat their respective models. However, in Hold’em, agents generally have lower win rates than the LLM-Player, possibly because the LLM-Player performs better in games with more psychological tactics, which are difficult to simulate with code while can be summarized by context learning.
We further visualize the action consistency between agents’ code and LLM-palyers in Figure 6.
We find that, in most games, the actions of LLM-players tend to resemble each other, and the strategies implemented in agents’ code also exhibit high similarity among themselves. However, the code-based strategies and the plain reasoning of the same model often differ substantially. The only exception is Bridge, where we find numerous cases in which both LLM-players and code agents exhibit low consistency with human decisions. Considering that Bridge allows for a certain degree of decision freedom and its bidding rules are not strictly unified, we attribute this phenomenon to the intrinsic characteristics of the game. Similar observations are also reported in other studies (Kita et al., 2024).
These findings further demonstrate that the strategies generated by the agents and those employed by the LLM-players are based on different approaches. This difference merits additional investigation in subsequent studies.
Appendix G Case Study on Code
G.1 Strategy of Agents
In Gomoku, strategies display clear stratification. Gemini and DeepSeek rely on random or near-random moves, while Claude, Doubao, GPT-5, and Qwen3 employ similar pattern-based evaluation with candidate filtering and Minimax search. Differences mainly lie in threat recognition, search control, and opening play: Claude and Doubao handle openings and forced moves more effectively, GPT-5 is steadier under time limits, Qwen3 remains balanced, whereas Gemini and DeepSeek are notably weaker.
The code similarity among agents in Texas Hold’em is relatively high, only DeepSeek employs a fully random strategy, while other models calculate winning probabilities based on the hand. On one hand, this is because the available actions in Texas Hold’em are limited to fold, call/raise, and check. On the other hand, the strategies for Texas Hold’em are relatively straightforward to implement, as both reasoning and code are primarily based on hand strength. As a result, the code can closely simulate the reasoning process.
In the case of Chess, DeepSeek relies on an external library (Stockfish), but fails to configure it correctly, resulting in unsuccessful development. Even after multiple development iterations, DeepSeek continues to use this library without resolving the configuration issues. We also find that Claude, Doubao, Gemini, GPT, and Qwen3 utilize a similar combination of heuristic piece and board evaluation, Minimax search, and alpha-beta pruning, which leads to similar behavior. There are slight differences in how each model evaluates the value of Chess pieces and the actions in endgame scenarios. Notably, Claude incorporates an opening book, which distinguishes it from the others and leads to better performance.
For Bridge, bidding and play strategies also stratify clearly. Qwen3 and Gemini rely on minimal logic, following random choices or only basic rules on High Card Points (HCP). By contrast, GPT-5, Doubao, and Claude incorporate structured evaluation, moving from total point counting (GPT-5) to multi-layered systems with suit quality, competitive actions, and signaling (Doubao and Claude). Despite these differences, all models share reliance on HCP as a core metric. Overall, Claude achieves the most complete integration of evaluation and play, Doubao is comparably advanced, GPT-5 remains simpler, while Gemini and Qwen3 lag behind.
(a) Holdem.
(b) Gomoku.
(c) Bridge.
(d) Chess.
Figure 6: Action consistency between agents’ code and LLM-Players on endgames.
Appendix H Results of ML Track
Table 13: ML ability scores and average rankings of agents.
[…]
Minimal
Claude-4-Sonnet 0.787 0.360 0.600 0.700 2.25
DeepSeek-Chat 0.612 0.000 0.170 0.100 4.25
Doubao-Seed 0.375 0.110 0.290 0.100 4.25
Gemini-2.5-Pro 0.000 0.000 0.140 0.675 5.00
GPT-5 0.625 0.530 0.900 0.700 1.50
Qwen3-Coder 0.600 0.000 0.900 0.725 2.50
Commercial
best ADK 0.750 0.190 0.700 0.656 1.25
Claude-Code 0.578 0.170 0.000 0.406 4.00
CodeX 0.484 0.190 0.400 0.469 3.00
Gemini-CLI 0.187 0.280 0.200 0.438 3.50
Qwen-Coder 0.500 0.170 0.700 0.531 2.50
The detailed results of agents’ performance on machine learning devlopments is shown in table 13.
In ML track, agents autonomously generate data, design code, train models, and deliver ML-based strategies in a GPU-enabled environment. The results of ML ability is provided in Appendix H. Most agents only implement basic models with limited training, resulting in smaller performance gaps and different rankings compared to the strategy track.
Appendix I Results of Multi-lingual Track
We list the detailed results on different languages in  14.
Table 14: Scores of agents on games with different languages, with variance analysis.
[…]
Minimal
Claude-4-Sonnet 1.000 0.250 0.250 0.125 0.360 0.640 0.000 0.069 1.000 0.250 0.250 0.125 0.106
DeepSeek-Chat 1.000 0.250 0.250 0.125 1.000 0.000 0.000 0.222 0.500 0.500 0.500 0.000 0.116
Doubao-Seed 1.000 0.500 0.000 0.167 1.000 0.000 0.000 0.222 0.500 1.000 0.000 0.167 0.185
Gemini-2.5-Pro 0.750 0.750 0.000 0.125 0.010 0.990 0.000 0.216 1.000 0.250 0.250 0.125 0.155
GPT-5 0.500 0.000 1.000 0.167 1.000 0.000 0.000 0.222 1.000 0.500 0.000 0.167 0.185
Qwen3-Coder 1.000 0.250 0.250 0.125 0.610 0.290 0.100 0.044 0.688 0.000 0.812 0.128 0.099
Commercial
Claude-Code 1.000 0.250 0.250 0.125 0.200 0.020 0.780 0.105 1.000 0.250 0.250 0.125 0.118
CodeX 0.000 0.500 1.000 0.167 0.200 0.030 0.770 0.100 1.000 0.000 0.500 0.167 0.145
Gemini-CLI 0.750 0.750 0.000 0.125 1.000 0.000 0.000 0.222 1.000 0.250 0.250 0.125 0.157
Qwen-Coder 1.000 0.250 0.250 0.125 1.000 0.000 0.000 0.222 0.975 0.000 0.525 0.159 0.169
Most agents achieve their highest scores in Python, while several models exhibit significant performance fluctuations in JS and Go. Qwen3-Coder demonstrates the most consistent results across all languages, with the lowest average variance, indicating strong cross-language adaptability. In contrast, models such as GPT-5 and Doubao-Seed show considerable differences between languages, reflecting limited generalization ability. Commercial agents also exhibit score differences across different programming languages. Considering that board game strategies are inherently language-agnostic tasks, the performance gaps observed in the multi-language track of CATArena indicate that current agents are not yet able to effectively abstract strategies into unified algorithms and implement them consistently across languages. Such algorithmic abstraction should be a key direction for the future development of agents.
Appendix J Cost and Code Complexity of Participants
Table 15: Comparison of model cost on four games of 1st round.
Metric Claude-4-Sonnet Deepseek-Chat Doubao-Seed Gemini-2.5-Pro GPT-5 Qwen3-Coder Claude-Code CodeX Gemini-CLI Qwen-Coder
Gomoku
Total input tokens 825964 34266 83307 1074278 392378 251590 925166 - 169567 623393
Session tokens 31571 8256 20186 27481 25201 20013 29340 - 21590 26478
Total output tokens 13777 1400 9432 25952 14184 6338 11746 - 2303 8002
Total time (s) 441.916 29.021 206.159 858.027 280.399 135.770 323.200 - 519.100 607.200
Tools used 68 12 20 118 44 32 38 24 8 26
Valid lines of code 574 69 405 65 421 425 350 347 101 335
Avg thinking time (s) 0.0337 0.0026 2.9848 0.0026 1.9681 0.0903 0.0053 0.7026 0.1151 0.0036
Texas Hold’em Poker
Total input tokens 966609.5 44883 97274 168565 345444 326508 296329.5 - 840690.5 1268906.5
Session tokens 45384 9388 14413 14073 26418 16688 29663.5 - 27828 23646.5
Total output tokens 20974 985 10370 40762 24178 4472 8537 - 11833.5 12772
Total time (s) 489.150 25.385 256.145 481.315 473.910 131.630 143.420 - 365.485 474.075
Tools used 38 8 12 15 20 24 12 19 39 48
Valid lines of code 742 21 169 92 346 164 122 318 300 220
Avg thinking time (s) 0.0026 0.0023 0.0030 0.0025 0.1004 0.0025 0.0028 0.0021 0.0024 0.0023
Bridge
Total input tokens 1429350 18027 81206 266506 258633 114679 618991 - 204913 71696
Session tokens 38087 15077 24374 29486 35448 20358 34263 - 46945 34937
Total output tokens 17039 1131 19247 12713 3954 5394 12284 - 13222 20308
Total time (s) 727.550 3.523 333.361 191.775 128.895 334.394 274.200 - 188.700 613.000
Tools used 94 6 14 32 26 16 25 24 5 32
Valid lines of code 521 0 552 280 245 271 464 362 349 549
Avg thinking time (s) 0.0095 0.0000 0.0105 0.0094 0.0095 0.0093 0.0091 0.0091 0.0093 0.0093
Chess
Total input tokens 1612835 49655 86674 346890 484611 267156 709737 - 153825 23489
Session tokens 38729 9953 14799 93292 39962 20781 29818 - 25418 23620
Total output tokens 16584 1257 14411 80997 28695 6624 13017 - 4727 6845
Total time (s) 617.339 26.882 287.976 356.551 444.932 141.417 285.593 - 162.342 142.344
Tools used 110 14 16 40 44 34 32 14 6 18
Valid lines of code 734 0 292 359 437 360 393 283 262 325
Avg thinking time (s) 1.3500 0.0000 0.0030 0.0030 0.2690 0.0020 1.9730 0.0030 0.0030 0.0050
Average across games
Total input tokens 1208689.625 36707.75 87115.25 464059.75 370266.5 239983.25 637555.875 - 342248.875 496871.125
Session tokens 38442.75 10668.5 18443.0 41083.0 31757.25 19460.0 30771.125 - 30445.25 27170.375
Total output tokens 17093.5 1193.25 13365.0 40106.0 17752.75 5707.0 11396.0 - 8021.375 11981.75
Total time (s) 568.989 21.203 270.910 471.917 332.034 185.803 256.603 - 308.907 459.155
Tools used 77.5 10.0 15.5 51.25 33.5 26.5 26.75 20.25 14.375 30.875
Valid lines of code 642.75 22.5 354.5 199.0 362.25 305.0 332.25 327.5 253.0 357.25
Avg thinking time (s) 0.3489 0.0012 0.7503 0.0044 0.5867 0.0260 0.4975 0.1792 0.0324 0.0050
Table 16: Comparison of model cost on four games of 2nd round.
Metric Claude-4-Sonnet Deepseek-Chat Doubao-Seed Gemini-2.5-Pro GPT-5 Qwen3-Coder Claude-Code CodeX Gemini-CLI Qwen-Coder
Gomoku
Total input tokens 1924785 170294 625384 589934 1351710 4859091 761468 - 514517 573759
Session tokens 102767 3495 84577 38760 78092 151664 37085 - 53402 27637
Total output tokens 14154 1608 16828 17098 10126 8230 11705 - 5180 7006
Total time (s) 433.112 765.076 436.142 1195.109 387.641 918.349 265.7 - 104.4 273.5
Tools used 58 62 32 58 48 82 30 22 12 23
Valid lines of code 617 69 527 305 467 368 493 302 251 376
Avg thinking time (s) 1.9630 0.0026 2.1291 0.0000 1.5888 0.3499 0.7542 0.2172 0.0080 0.0047
Texas Hold’em Poker
Total input tokens 1497923 56126 225806 2076630 1328520 3671952 1143571 - 1048050 1071708
Session tokens 143575 9819 57890 240057 113649 323278 53929 - 40566 49161
Total output tokens 7309 1151.5 10263.5 46280.5 13269.5 5986.5 7349.5 - 12617 6954
Total time (s) 308.05 45.935 327.225 1171.58 691.74 1057.84 866.715 - 1453.08 163.275
Tools used 16 11 10 17 21 25 28 18 41 27
Valid lines of code 283 21 110 156 362 293 264 319 363 315
Avg thinking time (s) 0.0026 0.0024 0.0024 0.0019 0.0912 0.0026 0.0023 0.0024 0.0025 0.0025
Bridge
Total input tokens 740733 18388 184814 277299 900865 720015 940709 - 275340 105056
Session tokens 28130 15318 4076 131089 119593 62734 38388 - 49946 63756
Total output tokens 14200 1178 18169 7520 4983 10270 15530 - 8161 25125
Total time (s) 397.164 6.097 391.047 162.147 367.204 1863.073 365.9 - 169.0 1337.7
Tools used 48 8 28 26 68 44 33 17 6 38
Valid lines of code 650 0 731 271 245 288 700 367 280 1112
Avg thinking time (s) 0.0101 0.0000 0.0135 0.0099 0.0096 0.0099 0.0105 0.0105 0.0102 0.0106
Chess
Total input tokens 2118018 45051 154263 426659 390211 1014853 4129953 - 308116 31752
Session tokens 101195 3954 28503 26538 35482 36705 72658 - 33919 31877
Total output tokens 13067 412 10063 13435 16798 11018 25438 - 6283 7250
Total time (s) 1352.194 522.132 230.664 931.231 214.792 263.665 906.521 - 420.146 177.453
Tools used 44 26 28 44 34 80 78 14 13 28
Valid lines of code 647 0 305 347 559 499 736 351 299 335
Avg thinking time (s) 0.8240 0.0000 0.0020 0.0000 0.0030 0.0020 0.0030 0.0040 1.8470 0.0050
Average across games
Total input tokens 1570364.75 72464.75 297566.75 842630.5 992826.5 2566477.75 1743925.25 - 536505.75 445568.75
Session tokens 93916.75 8146.5 43761.5 109111.0 86704.0 143595.25 50515.0 - 44458.25 43107.75
Total output tokens 12182.5 1087.38 13830.88 21083.38 11294.13 8876.13 15005.63 - 8060.25 11583.75
Total time (s) 622.63 334.81 346.27 865.02 415.34 1025.73 601.21 - 536.66 487.98
Tools used 41.5 26.75 24.5 36.25 42.75 57.75 42.25 17.75 18.0 29.0
Valid lines of code 549.25 22.5 418.25 269.75 408.25 362.0 548.25 334.75 298.25 534.5
Avg thinking time (s) 0.7000 0.0013 0.5367 0.0029 0.4231 0.0911 0.1925 0.0585 0.4669 0.0057
Table 17: Comparison of model cost on four variant games of 1st round.
Metric Claude-4-Sonnet Deepseek-Chat Doubao-Seed Gemini-2.5-Pro GPT-5 Qwen3-Coder Claude-Code CodeX Gemini-CLI Qwen-Coder
Gomoku
Total input tokens 833531 35489 83792 342027 539388 68087 495982 - 256919 23991
Complete tokens 22521 8587 23512 27175 27291 11468 27275 - 39232 24082
Total output tokens 10900 1394 11740 13553 19676 1658 8153 - 3522 7436
Total time (s) 307.319 24.007 255.817 809.100 365.077 73.260 191.800 - 74.300 140.200
Tools used 64 14 20 36 52 14 23 13 7 18
Valid lines of code 859 64 746 414 301 298 254 313 156 431
Avg thinking time (s) 0.0473 0.0026 0.0333 6.7802 0.0904 0.0048 0.0106 0.0051 0.0124 0.7938
Texas Hold’em Poker
Total input tokens 1182600 39782 89270 2582120 191064 385246 160410 - 864568 775880
Complete tokens 39575 9174 14448 32547 16530 16013 24287 - 25654 25308
Total output tokens 13110 962 9897 52457 16407 4361 5801 - 10699 7568
Total time (s) 398.130 20.195 217.600 1337.345 558.055 124.640 84.920 - 319.745 270.315
Tools used 42 7 11 86 17 30 7 19 42 34
Valid lines of code 469 21 204 98 302 144 167 256 232 243
Avg thinking time (s) 0.0026 0.0022 0.0011 0.0486 0.0019 0.0011 0.0025 0.0801 0.0028 0.0021
Bridge
Total input tokens 475377 16805 122427 232284 110768 653245 1908005 - 173628 26896
Complete tokens 37640 13764 36846 22593 4433 4438 53465 - 33184 27023
Total output tokens 16704 1137 7583 10020 2378 22823 18489 - 7412 5191
Total time (s) 298.800 7.721 158.254 483.948 304.010 7111.237 462.300 - 229.200 129.600
Tools used 34 6 16 32 40 308 55 16 5 22
Valid lines of code 627 0 291 213 133 165 469 368 212 224
Avg thinking time (s) 0.0071 0.0000 0.0068 0.0068 0.0096 0.0067 0.0089 0.0089 0.0060 0.0092
Chess
Total input tokens 1000994 38977 84149 432950 229279 253674 979611 - 184178 115022
Complete tokens 38485 7761 17321 43002 20520 17592 34597 - 26984 19923
Total output tokens 16355 1125 10518 32332 16499 5948 13010 - 6014 8032
Total time (s) 473.286 23.186 241.221 729.121 283.023 144.435 357.587 - 139.114 368.952
Tools used 74 14 14 40 32 38 41 12 8 19
Valid lines of code 719 41 399 398 475 343 410 427 152 392
Avg thinking time (s) 0.7520 0.0000 0.0000 0.2560 1.1190 0.0020 1.9660 0.0040 0.0040 0.0050
Average across games
Total input tokens 873125.5 32763.25 94909.5 897345.25 267624.75 340063.0 886002.0 - 369823.25 235447.25
Complete tokens 34555.25 9821.5 23031.75 31329.25 17193.5 12377.75 34906.0 - 31263.5 24084.0
Total output tokens 14267.25 1154.5 9934.5 27090.5 13740.0 8697.5 11363.25 - 6911.75 7056.75
Total time (s) 369.3838 18.7773 218.2230 839.8785 377.5413 1863.3930 274.1518 - 190.5898 227.2668
Tools used 53.5 10.25 15.25 48.5 35.25 97.5 31.5 15.0 15.5 23.25
Valid lines of code 668.5 31.5 410.0 280.75 302.75 237.5 325.0 341.0 188.0 322.5
Avg thinking time (s) 0.2022 0.0012 0.0103 1.7729 0.3052 0.0036 0.4970 0.0245 0.0063 0.2025
Table 18: Comparison of model cost on four variant games of 2nd round.
Metric Claude-4-Sonnet Deepseek-Chat Doubao-Seed Gemini-2.5-Pro GPT-5 Qwen3-Coder Claude-Code CodeX Gemini-CLI Qwen-Coder
Gomoku
Total input tokens 1298765 111940 680676 545504 606704 1542046 805252 - 227722 873934
Complete tokens 132219 11100 121218 24003 82812 4309 38629 - 41093 32599
Total output tokens 7118 1902 26565 18241 4189 22014 11529 - 5554 8531
Total time (s) 261.323 76.973 538.087 325.892 179.801 25787.922 294.400 - 80.200 836.300
Tools used 46 30 28 70 38 538 30 23 6 30
Valid lines of code 264 64 535 111 276 43 328 302 210 384
Avg thinking time (s) 0.0212 0.0030 0.0855 0.0000 0.0904 0.0000 0.0884 0.0162 0.0070 0.0071
Texas Hold’em Poker
Total input tokens 1627500 50339 656196 675580 1452101 1822342 448924 - 1725678 2472353
Complete tokens 143514 11897 115933 103332 111407 118321 48631 - 48378 67453
Total output tokens 8180 1130 11609 18771 12897 6227 5880 - 16033 10296
Total time (s) 367.720 27.960 359.490 515.000 408.505 294.910 612.525 - 670.535 283.440
Tools used 18 10 16 20 22 27 16 23 53 48
Valid lines of code 289 22 272 105 324 295 167 256 232 243
Avg thinking time (s) 0.0025 0.0019 0.0019 0.0031 0.0019 0.0026 0.0025 0.0801 0.0028 0.0021
Bridge
Total input tokens 706814 17186 32819 775688 433901 866917 1110504 - 189659 48922
Complete tokens 41164 14021 14545 80161 $41$ 4944 45001 - 46073 49034
Total output tokens 15867 1188 13212 39919 1485 13928 17233 - 7966 7252
Total time (s) 359.864 6.074 189.645 279.231 1106.275 3238.660 411.500 - 77.700 220.000
Tools used 52 8 12 46 112 158 38 21 4 25
Valid lines of code 603 0 233 396 166 218 747 295 441 531
Avg thinking time (s) 0.0069 0.0000 0.0112 0.0067 0.0065 0.0101 0.0093 0.0091 0.0091 0.0092
Chess
Total input tokens 1475420 34542 172443 400275 473325 1682584 2213785 - 247641 109313
Complete tokens 56660 8283 31582 97638 32082 44129 43227 - 23389 8140
Total output tokens 22989 1473 16610 35371 16469 14241 24090 - 4834 9129
Total time (s) 696.717 32.412 366.723 1093.706 350.225 511.943 605.236 - 237.026 427.591
Tools used 82 14 30 62 44 116 67 16 13 29
Valid lines of code 1146 72 561 345 506 577 407 434 187 454
Avg thinking time (s) 1.6260 0.0000 0.0000 0.0440 0.7560 0.1780 0.0030 0.0040 1.7130 0.0050
Average across games
Total input tokens 1277124.75 53501.75 385533.5 599261.75 741507.75 1478472.25 1144616.25 - 597675.0 876130.5
Complete tokens 93389.25 11325.25 70819.5 76283.5 57610.5 42925.75 43872.0 - 39733.25 39306.5
Total output tokens 13538.5 1423.25 16999.0 28075.5 8760.0 14102.5 14683.0 - 8596.75 8802.0
Total time (s) 421.4060 35.8547 363.4863 553.4572 511.2015 7458.3587 480.9153 - 266.3653 441.8328
Tools used 49.5 15.5 21.5 49.5 54.0 209.75 37.75 20.75 19.0 33.0
Valid lines of code 575.5 39.5 400.25 239.25 318.0 283.25 412.25 321.75 267.5 403.0
Avg thinking time (s) 0.4141 0.0012 0.0247 0.0134 0.2137 0.0477 0.0258 0.0273 0.4330 0.0059
We list the agents’ cost and code statistics in table 15 for first round development of standard games,  16 for second round development of standard games, 17 for first round development of vairant games and  18 for second round development of variant games.
We can see that game development token costs show minimal variation, while differences is significant due to model changes. Claude (both minimal and code-based agents) consumes significantly more input tokens than competitors, exceeding the average by over 2 times, while Gemini generates notably more output tokens compared to other models. GPT-5 offers the best trade-off between cost. Among all agents, second-round game development require more input tokens, while output token growth remains marginal. In addition, commercial agents consistently use fewer tokens than their minimal-agent counterparts.
In terms of code complexity, agents driven by Claude-4 model consistently surpasses other agents in both the number of effective lines of code developed and the time spent considering development strategies. We observe that its development strategies are more sophisticated. Additionally, the complexity of its code increases with each iteration, which indirectly demonstrates the model’s exceptional learning capabilities.
Appendix K Full Prompts
The agent is instructed to develop a competitive game AI based on the provided game environment. The AI must be deployed as an HTTP service with a single-port startup script, follow the official development instructions, and be named with the model prefix. And the agent is encouraged to iteratively improve its strategy based on the tournament report and the previous codes. Full prompt of details are in Table 19 for main leaderboard,  LABEL:tab:ml_game_ai_prompt for ML track and  21 for multilingual track.
Table 19: Game AI Development Prompt
Game Development
There is a game project under game env path. You need to read its source code and develop a game AI. Your AI will compete against other AIs in a tournament, so please make your AI as strategic and competitive as possible.
The final AI should be provided as an HTTP service. You can refer to the guides in game env path / README.md and game env path / develop instruction.md for development instructions. The content in game env path / develop instruction.md is very important, please read it carefully!
Please develop your AI service directly under dir path.
Script Requirements
Please implement a script to start your AI service, with the name start_ai.sh in dir path. The script must accept exactly one argument, which is the port number game port to run the HTTP service. You should be able to start the AI service on a specified port by running:
bash start_ai.sh <port>
Your AI service should listen on the given port, and you can check its health status by running: curl -s http://localhost:<port>/health
Note: The script should not accept any other arguments except for the port number. Make sure your AI service uses this port for HTTP requests.
Other Requirements:
Use your model name as a prefix in the name of your AI service, i.e., model name_AI. Develop directly in dir path without repeatedly asking for the next step. Report to me only after you have completed the development.
Access the main server
You can play game of game env pathin at game server. You can play the games with your own AI or any other AI to improve your strategy. You can use bash tools to improve yourself.
Final Remind
You should write start_ai.sh in dir path and implement the AI service in dir path. DO NOT MODIFY THE CODE IN game env path
Condition (if round_num>1round\_num>1):
Tournament report of last round is in last round log dir. The historical records are quite large, please use tools start_interactive_shell and run_interactive_shell to analyze the data efficiently. The code corresponding to the log is stored in last round code dir. Please learn from it and improve your strategy.
Table 20: Machine Learning Game AI with MANDATORY Self-Play Training Prompt
Machine Learning Game AI with MANDATORY Self-Play Training
Develop a competitive game AI for game env path using REAL machine learning with actual training.
CRITICAL: NO PSEUDO-ML ALLOWED
MANDATORY: Implement real training with actual parameter updates.
Forbidden: Random weights, unused optimizers, no training loops
Required: Self-play training, loss.backward(), optimizer.step(), saved trained model
Training Requirements
1. Self-Play System: Generate training data by playing against itself
2. Training Loop: Real parameter updates with backpropagation
3. Model Saving: Save trained model weights (e.g., trained_model.pth)
4. Training Endpoint: /train HTTP endpoint to trigger training
Technical Implementation
The final AI should be provided as an HTTP service. You can refer to the guides in game env path/README.md and game env path/develop_instruction.md for development instructions. The content in game env path/develop_instruction.md is very important, please read it carefully!
Please develop your AI service directly under dir path.
Script Requirements
Implement a script to start your AI service, with the name start_ai.sh in dir path. The script must accept exactly one argument, the port number to run the HTTP service. Start the AI via:
bash start_ai.sh <port>
Check service health via: curl -s http://localhost:<port>/health
Additional Training Script: Also create train_ai.sh for self-play training:
bash train_ai.sh <num_episodes>
Note: The script should not accept any other arguments except for the port number. Ensure the AI service uses this port for HTTP requests.
Environment Requirements
MANDATORY: You MUST develop under the ml_env_py311 conda environment. PyTorch, TensorFlow, NumPy, pandas, and scikit-learn are already installed.
- DO NOT include PyTorch or TensorFlow in requirements.txt
- Create a requirements.txt file only for additional dependencies you need beyond the pre-installed packages
- All development and execution must be done in the ml_env_py311 environment
Available Game Port
We offer an interactive game port: http://localhost:game port to facilitate your testing and training. You can interact with the service and get feedback from the port.
Validation Requirements
Submission will be rejected if:
- The model uses only random weights without training
- No actual training loop with parameter updates
- No self-play data generation system
- Cannot demonstrate learning/improvement over time
- Training endpoints return fake responses without real training
Other Requirements
Use your model name as a prefix in the name of your AI service, i.e., model name_AI.
Develop directly in dir path without repeatedly asking for the next step. Report to me only after you have completed the development.
Final Remind
You should write start_ai.sh and train_ai.sh in dir path and implement the REAL ML-based AI service with actual training capabilities in dir path. DO NOT MODIFY THE CODE IN game env path. Please make sure to implement a genuine machine learning training workflow, including writing the code, performing training, tuning hyperparameters, and testing the model. After training, save the trained model parameters, and ensure that the AI service can be started with the specified script to perform inference.
Mandatory: Submission must include a trained model file (e.g., trained_model.pth, model_weights.pkl) proving real training.
Zero Tolerance for Pseudo-ML: Any submission without real training will be rejected.
Table 21: Multi-language Game AI Development Prompt
Game Development
There is a game project under game env path. You need to read its source code and develop a game AI. Your AI will compete against other AIs in a tournament, so please make your AI as strategic and competitive as possible.
The final AI should be provided as an HTTP service. You can refer to the guides in game env path / README.md and game env path / develop instruction.md for development instructions. The content in game env path / develop instruction.md is very important, please read it carefully!
Please develop your AI service directly under dir path.
Script Requirements
Please implement a script to start your AI service, with the name start_ai.sh in dir path. The script must accept exactly one argument, which is the port number game port to run the HTTP service. You should be able to start the AI service on a specified port by running:
bash start_ai.sh <port>
Your AI service should listen on the given port, and you can check its health status by running: curl -s http://localhost:<port>/health
Note: The script should not accept any other arguments except for the port number. Make sure your AI service uses this port for HTTP requests.
Other Requirements:
Use your model name as a prefix in the name of your AI service, i.e., model name_AI. Develop directly in dir path without repeatedly asking for the next step. Report to me only after you have completed the development.
Access the main server
You can play game of game env pathin at game server. You can play the games with your own AI or any other AI to improve your strategy. You can use bash tools to improve yourself.
Final Remind
You should write start_ai.sh in dir path and implement the AI service in dir path. DO NOT MODIFY THE CODE IN game env path
Condition (if language = JS)
JavaScript is the language you should use to develop your AI service. The version of Node.js is node version, the path of Node.js is node path, and it is already set in the PATH environment variable. You can use node to run the program.
Condition (if language = Go)
Go is the language you should use to develop your AI service. The version of Go is go version, the path of Go is go path, and it is already set in the PATH environment variable. You can use go to build the program.

### v2

@@
-CATArena: Evaluation of LLM Agents Through Iterative Tournament Competitions
+CATArena: Evaluating Evolutionary Capabilities of Code Agents via Iterative Tournaments
 Abstract
-Large Language Model (LLM) agents have evolved from basic text generation to autonomously completing complex tasks through interaction with external tools.
-However, current benchmarks mainly assess end-to-end performance in fixed scenarios, restricting evaluation to specific skills and suffering from score saturation and growing dependence on expert annotation as agent capabilities improve.
-In this work, we emphasize the importance of learning ability, including both self-improvement and peer-learning, as a core driver for agent evolution toward human-level intelligence.
-We propose an iterative, competitive peer-learning framework, which allows agents to refine and optimize their strategies through repeated interactions and feedback, thereby systematically evaluating their learning capabilities.
-To address the score saturation issue in current benchmarks, we introduce CATArena, a tournament-style evaluation platform featuring four diverse board and card games with open-ended scoring.
-By providing tasks without explicit upper score limits, CATArena enables continuous and dynamic evaluation of rapidly advancing agent capabilities.
-Experimental results and analyses involving both minimal and commercial code agents demonstrate that CATArena provides reliable, stable, and scalable benchmarking for core agent abilities, particularly learning ability and strategy coding.
-[…]
+Current evaluation for Large Language Model (LLM) code agents predominantly focus on generating functional code in single-turn scenarios, which fails to evaluate the agent’s capability for continuous code optimization and multi-turn iterative development.
+To bridge this gap, we introduce CATArena, a framework designed to evaluate the evolutionary capabilities of code agents via iterative tournaments.
+Agents engage in multi-turn tournaments and continuously refine their code through self-reflection and peer-learning based on comprehensive execution feedback.
+For evaluation, we propose a dual-metric system to decouple static generation proficiency from evolutionary potential.
+Extensive experiments reveal that an agent’s evolutionary potential is not strictly correlated with its initial proficiency.
+Our analysis further reveals that current agents struggle to concurrently leverage both peer-learning and self-reflection for effective performance gains.
+Furthermore, the results validate CATArena’s high extensibility and resistance to variance tasks, establishing it as a continuous and reliable standard for assessing the evolutionary capability of LLM code agents.
+[^1] The official website of CATArena is https://catarena.ai.
+Code Agents, Evaluation, Evolutionary Capabilities
 1 Introduction
-With the rapid evolution of agents powered by large language models (LLMs), their capabilities have far surpassed simple text generation.
-By actively invoking external tools, LLM agents have significantly expanded the boundaries of artificial general intelligence (AGI).
-These agents are now able to autonomously complete complex, multi-step tasks that are previously considered beyond their reach, such as developing software project (Manish, 2024; Hu et al., 2025b), intelligently performing strategic planning (Belle et al., 2025), and learning user preference (Gao et al., 2024).
-Existing benchmarks mainly focus on end-to-end performance in specific tasks, such as code generation (Yang et al., 2024), AI research (Nathani et al., 2025), and GUI automation (Wang et al., 2024).
-These benchmarks provide detailed observations and analyses of LLM agents’ abilities within particular scenarios and have driven significant progress in the field.
-However, there are important limitations to these approaches.
-First, the scores obtained in these end-to-end benchmarks only reflect performance on specific tasks, whereas an agent’s overall capability is composed of multiple fundamental skills working together.
-Second, the absolute scores in these benchmarks, which are typically based on objective correctness, have an upper bound.
-As agents become increasingly powerful, maintaining and updating these benchmarks requires additional expert-level annotation, and the level of required expertise continues to rise.
-In light of these challenges, there is an urgent need for a quantifiable and continuously evolving benchmark that systematically measures and analyzes the fundamental sub-abilities of agents.
-Previous research has shown that self-learning is an essential ability for agents to achieve human-level intelligence (Gao et al., 2025; Zhu et al., 2025).
-Beyond self-learning, agents, similar to humans, also engage in peer learning, which enables collective evolution through interactions and shared experiences (Liu et al., 2024).
-During this evolutionary process, agents receive feedback from their environment and continually improve themselves.
-This capacity for learning and adaptation is indispensable for LLM agents, as it prepares them for ongoing evolution and more complex challenges.
-To systematically evaluate this crucial ability, we propose an iterative peer-learning-based competitive framework for LLM agents.
-In each iteration, agents are required to revise and update their strategies based on the outcomes and policies observed in previous rounds of competition.
-After every update, the agent policy codes are executed and competed against each other, generating dynamic performance rankings.
-Through this peer-learning architecture, we gain valuable insights into the learning abilities of LLM agents.
-Building on this peer-learning framework, we introduce CATArena (Code Agent Tournament Arena) [^1] Code of CATArena is available in https://github.com/AGI-Eval-Official/CATArena., which utilizes four open-ended, rankable games.
-These games, including both board games and card games, provide LLM agents with a peer-learning environment and unlimited upper bound for improvement.
-They enable agents to continually improve and compete, ensuring that the evaluation framework remains challenging as agent capabilities grow.
-Furthermore, our competitive arena is inherently extensible and can be readily adapted to other types of open-ended, rankable tasks, facilitating the assessment of core agent abilities in new domains.
-As agent capabilities continue to advance, CATArena can evolve by incorporating tasks with greater complexity and discrimination, thereby supporting ongoing evaluation without the need for expert-level human annotation.
-In our experiments, we conduct comparative performance evaluations and data analysis conducted on our self-developed minimal code agent and state-of-the-art commercial code agents.
-CATArena consistently provides stable and reliable benchmarks for assessing both agent capabilities and the agentic potential of the underlying LLMs.
-Within the peer-learning framework, we design general scoring metrics to systematically assess the fundamental abilities of participating agents, including their learning ability.
-Our experiments demonstrate that the strategy coding tasks applied in CATArena are fundamentally different from traditional LLM reasoning tasks.
-This represents a novel evaluation dimension that has not been addressed in previous work.
-Additionally, we analyze characteristics of CATArena, demonstrating its reliability and extensibility as a benchmarking platform.
+The rapid evolution of Large Language Models (LLMs) has catalyzed the emergence of code agents.
+Unlike passive chat assistants, code agents actively interact with dynamic digital workspaces, manipulating file systems and utilizing toolchains to orchestrate complex software development workflows (Dong et al., 2025; Dearing et al., 2025; Xu et al., 2025c).
+To assess these expanding capabilities, the evaluation landscape has evolved from simple method-level generation (Chen et al., 2021) to repository-level engineering scenarios (Xu et al., 2025a, b).
+Performance is predominantly quantified by static metrics such as task success rate (Yang et al., 2024) within a single-turn development process.
+While some emerging benchmarks incorporate non-functional metrics such as security or code complexity, these remain static checks rather than dynamic objectives.
+These static and correctness-centric benchmarks impose two critical limitations on advancing agent intelligence.
+First and foremost, existing protocols suffer from the absence of iterative evolution.
+Real-world software development is inherently incremental and collaborative (Bennett, 1996): human engineers continuously refine their solutions by analyzing execution logs and learning from peer reviews (Liu et al., 2024).
+Crucially, this evolutionary mechanism is equally indispensable for autonomous agents.
+However, by restricting evaluation to a single-turn mode, current frameworks isolate agents from the dynamic feedback loops essential for optimization and evolution.
+Agents are prevented from rectifying runtime errors through self-debugging or optimizing strategies by observing superior peer implementations (Gao et al., 2025; Zhu et al., 2025a).
+Consequently, current benchmarks capture only a static snapshot of the agent’s baseline competency, failing to evaluate its dynamic capacity for self-reflection and peer-learning over time.
+Furthermore, this lack of iterative depth leads to an underemphasis on code quality optimization.
+Since evaluation is confined to a single turn, current frameworks predominantly rely on binary functional correctness metrics like pass@k (Chen et al., 2021; Austin et al., 2021), failing to incentivize agents to refine code for computational efficiency or robustness.
+In large-scale software engineering, a functionally correct but inefficient algorithm can lead to unacceptable system latency or energy consumption (Dearing et al., 2025).
+However, standard benchmarks treat a brute-force solution as equivalent to an optimized expert implementation.
+This equivalence masks the performance gap between agents that merely solve a problem and those that engineer superior solutions, causing metrics to saturate without capturing the potential for unbounded evaluation.
+To bridge this gap, we introduce CATArena (Code Agent Tournament Arena), a dynamic evaluation framework designed to assess the evolutionary capabilities of code agents through iterative refinement.
+Unlike traditional static benchmarks, CATArena immerses agents in a multi-turn development loop, mimicking the continuous improvement process of human engineers.
+Within this interactive workspace, agents must go beyond generating functional code.
+They are required to analyze execution feedback and, crucially, leverage peer solutions and opponent strategies to iteratively optimize their performance.
+To ensure an unbounded evaluation, we instantiate six diverse tasks with high performance ceilings and design scoring mechanisms that reward continuous optimization.
+These tasks are categorized into two types based on the feedback signal: two objective optimization tasks, where agents optimize against fixed metrics (e.g., time complexity, throughput); and four competitive tasks, a novel introduction where agents adaptively optimize their strategies against dynamic opponents (e.g., in strategy games, network attack and defense).
+This dual-track design allows the framework to seamlessly integrate diverse task types while effectively measuring an agent’s potential to evolve from basic correctness to expert-level excellence.
+We conduct extensive experiments using a multi-round iterative process, evaluating both self-developed minimal agents powered by various LLMs and state-of-the-art (SoTA) commercial code agents.
+Our analysis reveals that evolutionary capability is a distinct dimension of intelligence, independent of single-turn generation capability.
+We observe that agents with high initial pass rates do not necessarily possess strong evolving abilities.
+Conversely, some agents demonstrate low-start, high-growth trajectories, effectively leveraging feedback to surpass initially superior competitors.
+Furthermore, by analyzing the optimization trajectories, we identify distinct evolutionary patterns in how agents utilize peer solutions and execution logs.
+Ultimately, these results demonstrate that the iterative framework enables agents to overcome the performance plateaus typical of single-round development, validating the necessity of unbounded evaluation for the next generation of intelligent agents.
 In summary, our contributions are as follows:
-  • Iterative Peer-learning-based Competitive Framework: We propose a novel framework that leverages iterative peer-learning and competition to evaluate the learning abilities of LLM agents.
-Agents continuously revise their strategies based on feedback and outcomes from previous rounds, aligning agent evolution with human evolution.
-  • CATArena Benchmark: We introduce CATArena, a tournament-style benchmark for evaluating the basic capabilities of LLM agents using a diverse set of open-ended games, including board and card games.
-CATArena provides an unlimited upper bound for agent improvement and supports extensible evaluation across diverse, open-ended tasks.
-  • Comprehensive Agent Evaluation: We design general and systematic evaluation matrices and conduct comparative experiments and analyses between our minimal code agent and state-of-the-art commercial agents, demonstrating the reliability, stability, and extensibility of the CATArena.
+  • Evolutionary Capability Evaluation: To the best of our knowledge, we are the first to identify and formalize the evolutionary capability of code agents, aligning with the iterative nature of real-world engineering.
+  • CATArena Framework: We introduce CATArena, a tournament-based evaluation framework that models the development process as a multi-turn evolutionary loop.
+Driven by iterative feedback, it quantifies code refinement via self-reflection and peer-learning, transcending static, single-turn evaluation.
+  • Comprehensive Agent Analysis: We demonstrate that evolutionary capability is orthogonal to static generation performance and analyze its temporal dynamics across tasks.
+Our extensible experiments further highlights CATArena’s extensibility with new environments, revealing that SoTA agents still faces significant challenges on complex tasks.
 2 Related Work
-Learning Ability.
-Learning ability is crucial for LLM agents, as it enables continual adaptation and improvement in dynamic and complex environments.
-Recent studies have shown that self-learning methods, such as self-refinement (Madaan et al., 2023; Shinn et al., 2023), allow models to enhance their outputs through iterative feedback, while environmental feedback further supports continual learning (You et al., 2024).
-In addition to self-learning, peer-learning has also been increasingly recognized, with approaches encouraging agents to learn from others’ reasoning processes and shared experiences (Liang et al., 2024; Luo et al., 2025).
-These diverse learning mechanisms have led to notable advances in tasks such as code generation, complex reasoning, and collaborative problem-solving.
-In the context of LLM-driven agents, learning ability represents a critical capability that supports effective adaptation and enables agents to tackle increasingly complex tasks and evolving challenges (Zhu et al., 2025; Gao et al., 2025).
-Despite the progress, systematic evaluation of how agents learn from each other remains underexplored, highlighting the need for benchmarks that capture both self-learning and peer-learning abilities.
-Evaluation on Agents.
-Recent benchmarks primarily assess LLM-driven agents on end-to-end, task-specific abilities.
-Code-based evaluations such as GitTaskBench (Ni et al., 2025), SUPER (Bogin et al., 2024), ProjectEval (Liu et al., 2025), SWE-PolyBench (Rashid et al., 2025), RedCode (Guo et al., 2024), SWT-Bench (Mündler et al., 2025), InfiAgent-DABench (Hu et al., 2024), and DA-Code (Huang et al., 2024) focus on large-scale software development, code security, bug fixing, and data science tasks.
-Other works extend agent evaluation to research (Du et al., 2025), real-world tool use (Yao et al., 2024), and assistant scenarios (Mialon et al., 2024).
-While some benchmarks explore agent-vs-agent evaluation (Zhuge et al., 2024), most rely heavily on human annotation and objective correctness, leading to upper bounds and saturation as agent capabilities advance.
-Open-ended Tasks.
-To address these limitations, recent work has leveraged open-ended, rankable tasks.
-Benchmarks such as GameBench (Costarelli et al., 2024), lmgame-Bench (Hu et al., 2025a), GAMEBot (Lin et al., 2024), and card game evaluations (Wang et al., 2025) assess LLMs’ strategic reasoning through diverse games.
-Frameworks like Game Reasoning Arena (Cipolina-Kun et al., 2025), GVGAI-LLM (Li et al., 2025), ZeroSumEval (Alyahya et al., 2025), TextArena (Guertler et al., 2025), and MCU (Zheng et al., 2025) further extend evaluation to multi-turn reasoning, spatial adaptability, natural language interaction, and open-ended tasks.
-However, these benchmarks mainly focus on reasoning skills and do not systematically evaluate agents’ learning abilities or coding strategies.
-Our analysis shows that measuring learning ability and coding strategy is fundamentally distinct, and both are essential for advancing agent intelligence.
-Additionally, in human board game competitions, variant rules such as Chess960 (FIDE, 2023) and Six-plus Hold’em (sixplusholdem, 2025) are often introduced to reduce memorization and encourage creativity.
-Notably, these variant rules have received relatively less attention in model evaluation.
-3 CATArena
-Figure 1: Overview of the evaluation framework and CATArena.
-The evaluation framework adopts an iterative peer-learning based competitive process.
-In round 1, LLM agents develop initial strategies via coding.
-These strategies are matched in a tournament arena, producing rankings and logs.
-In each subsequent round, agents analyze previous codes and logs, refine their strategies, and compete again.
-CATArena includes four open-ended and rankable games to cover diverse settings.
-Based on the tournament results from all rounds, a carefully designed scoring matrix and evaluation scheme are used to robustly quantify various agent abilities.
-3.1 Iterative Peer-Learning based Competitive Framework
-As shown in Figure 1, we propose an iterative peer-learning-based competitive framework, where CATArena evaluates code agents through a two-phase workflow: initial strategy development and iterative improvement.
-The initial phase assesses each agent’s ability to independently implement a baseline strategy based on the game code, while the iterative phase focuses on the agent’s learning ability.
+Figure 1: Overview of the CATArena Framework.
+(1) Initial Development: Agents receive task specifications and generate initial solution code $C^{(1)}$.
+(2) Execution Feedback: Submissions are evaluated in either competitive or objective tracks.
+Results are aggregated into a structured set containing execution logs, leaderboards, and peer solutions.
+(3) Iterative Evolution: Agents engage in self-reflection and peer-learning to refine their code $C^{(n)}$ for the subsequent rounds.
+Evolutionary Mechanisms.
+Evolutionary capability is pivotal for agents to adapt in dynamic environments.
+For LLM-based agents, this manifests as in-context evolution, where agents optimize strategies using historical context and execution feedback without parameter updates (Zhu et al., 2025a).
+Recent research highlights two primary mechanisms driving this evolution: self-reflection and peer-learning.
+Self-reflection methods (Madaan et al., 2023; Shinn et al., 2023) enhance outputs through iterative self-correction based on environmental feedback (You et al., 2024).
+Peer-learning approaches encourage agents to distill insights from the reasoning processes and solutions of others (Liang et al., 2024; Luo et al., 2025).
+While these mechanisms have advanced code generation and reasoning, the systematic evaluation of an agent’s evolutionary capability remains underexplored, particularly regarding how agents leverage feedback to evolve over time in competitive coding scenarios.
+CATArena fills this void as the first to assess the evolutionary capability of code agents.
+Evaluation of Code Agents.
+Recent benchmarks have scaled evaluation to simulate complex, real-world software engineering scenarios.
+In the domain of repository-level development, frameworks such as GitTaskBench (Ni et al., 2025), SWE-PolyBench (Rashid et al., 2025), and SWT-Bench (Mündler et al., 2025) challenge agents with large-scale maintenance and bug fixing, complementing established standards like SWE-bench (Yang et al., 2024).
+Beyond general engineering, specialized evaluations have emerged for diverse domains: RedCode (Guo et al., 2024) and SecBench (Lee et al., 2025) focus on code security; InfiAgent-DABench (Hu et al., 2024) and DA-Code (Huang et al., 2024) target data science workflows; while other works assess research automation (Du et al., 2025) and web development (Xu et al., 2025b).
+Despite this extensive coverage, the evaluation methodology remains predominantly static.
+Performance is typically quantified by binary success metrics such as Pass@k (Chen et al., 2021) and task success rate (Yang et al., 2024).
+Although some benchmarks incorporate non-functional metrics like code complexity (Siddeeq et al., 2025) or test coverage (Chen and Jiang, 2025), these are applied as static checks.
+To overcome the limitations of static metrics, research has expanded into open-ended tasks.
+In the realm of strategic reasoning, benchmarks like GameBench (Costarelli et al., 2024) and TextArena (Guertler et al., 2025) assess agents through multi-turn gameplay.
+Similarly, for code optimization, datasets such as OIBench (Zhu et al., 2025b) and SWE-bench (Yang et al., 2024) provide open-ended challenges for algorithmic reasoning and engineering performance.
+However, current evaluations for these tasks are predominantly single-turn, treating them as static problems rather than opportunities for continuous optimization.
+This restriction prevents code agents from demonstrating their full potential through iterative refinement.
+CATArena bridges this gap by incorporating these open-ended tasks into an evolutionary framework, rigorously testing the limits of code agents.
+3 CATArena Framework
+In this section, we detail the CATArena framework, a universal and robust evaluation paradigm designed to shift the assessment of code agents from static correctness to dynamic evolutionary capability.
+3.1 Iterative Evolution via Tournament Feedback
+Unlike traditional benchmarks that treat coding as a static, single-turn task, CATArena abstracts the evaluation process into a continuous loop of development, execution, and evolution.
+As illustrated in Figure 1, this paradigm is structured into two primary phases bridged by an execution feedback mechanism.
+Table 1: Overview of tasks in CATArena.
+The framework instantiates six diverse tasks.
+Crucially, the feedback mechanism provides agents with population-level data, including full tournament logs and peer solutions, rather than just individual execution results, fostering group evolution.
+Category Task Domain Players Feedback Content
+Competitive Gomoku Board Game 2 Adversarial Match Logs & Peer Solutions
+Texas Hold’em Card Game $\geq 8$
+Chess Board Game 2
+Bridge Card Game 4
+Objective OIBench (Zhu et al., 2025b) Algorithm Optimization 1 Deterministic Efficiency Metrics & Peer Solutions
+SWE-Perf (He et al., 2025) Software Optimization 1
 Initial Development (Round 1).
-In this stage, each agent receives the game code and a sample AI implementation.
-Without external guidance, each agent must develop its own strategy to participate in the tournament.
-This phase primarily examines the agent’s strategy coding ability and establishes a baseline for subsequent evaluation.
-Iterative Improvement (Rounds $n>1$).
-After the first round, all strategies submitted are evaluated through a tournament (round-robin format for symmetric games, batch-based competition for asymmetric games).
-Comprehensive competition logs are generated, recording rankings, win counts, and move histories for all matches.
-In subsequent rounds, agents are provided with the game code, previous round submissions from all participants, and these detailed logs.
-Agents must analyze these resources from previous rounds to adapt and improve their own strategies.
-This phase assesses the agent’s learning ability through repeated cycles of analysis and refinement.
-This iterative evaluation framework of CATArena enables a granular assessment of both basic coding skills and advanced learning capabilities, supporting a robust and scalable measurement of code agent performance.
-Table 1: Overview of game arenas and representative variants in CATArena.
-Game Symmetry Type Players Variant
-Gomoku Board 22 Forbidden points; dual three-in-a-row
-Texas Hold’em Card $\geq 8$ Card removal; swapped hand ranks
-Chess Board 22 Chess960; forbidden/special moves
-Bridge Card 44 Card exchange
-    For Bridge, symmetry is defined by assigning identical agent strategies to both teammates.
-3.2 Games and Variants
-Building on the tournament-based evaluation framework, CATArena deploys four distinct game arenas, each selected to test the strategic reasoning and coding capabilities of code agents across varying levels of complexity and interaction patterns.
-These arenas include competitive and cooperative settings, as well as symmetric and asymmetric game structures, thus enabling a diverse analysis of agents’ strategy coding abilities and learning patterns.
-In addition to standard rules, each game is extended with thoughtfully designed variants that introduce novel or altered mechanics, inspired by real-world adaptations such as Fischer Random Chess (Chess960) (FIDE, 2023).
-Like human competition, the variant rules encourage strategy generalization and penalize rote memorization, as most models are trained on card and board game data.
-Table 1 provides an overview of the selected games and their respective variants.
-3.3 Tournament Format and Scoring System
-After the completion of all $N$ development rounds, CATArena conducts a comprehensive tournament to quantitatively evaluate agent strategies and compute performance metrics.
-A total of $T$ agent models participate, each contributing strategies in every round of development.
-Tournament formats are tailored to game types: for symmetric games, all strategies engage in a round-robin cycle, ensuring exhaustive pairwise competition; for asymmetric games such as Texas Hold’em, strategies are grouped into batches and compete in multi-agent matches.
-To mitigate randomness, all matches are repeated multiple times, and results are averaged for robust evaluation.
-Scores are recorded in a scoring matrix $W\in\mathbb{R}^{(TN)\times(TN)}$, where $W_{i,j}^{n,m}\in[0,1]$ denotes the score obtained by agent $i$’s strategy in round $n$ against agent $j$’s strategy in round $m$.
-When $n=m$, the notation simplifies to $W_{i,j}^{n}$; similarly, when $i=j$, it is denoted as $W_{i}^{m,n}$.
-This scoring system enables fine-grained, quantitative analysis of agent performance in both individual and iterative development stages.
-For asymmetric games, pairwise results are not feasible; instead, batch-based tournaments are used and the score matrix records the win rates of multi-agent matches.
-The tournament format is provided in the Appendix A.
-3.4 Evaluation Metrics
-Based on the scoring matrix $W$, we design a set of evaluation metrics to quantitatively assess the key capabilities of code agents.
-Specifically, our metrics are constructed to measure three core capabilities: strategy coding, learning, and generalizability.
-In the following sections, we define these metrics using symmetric games as examples.
-For asymmetric games, the evaluation principles remain consistent.
-The calculation of the scoring matrix $W$ is provided in Appendix B.
-Strategy Coding.
-Strategy coding measures the agent’s fundamental ability to abstract game strategies into reproducible algorithms and implement them as executable code, which is fundamentally different from general reasoning and strategic planning abilities.
-In CATArena, this metric evaluates how effectively an agent can independently develop a baseline strategy for the game environment and compete against other agents in the initial development stage.
-For each agent $i$, strategy coding is quantified by the average score obtained against all other agents in the first round:
-$S_{i}=\mathrm{avg}_{j\neq i}(W_{i,j}^{1}).$
-This metric serves as the foundational benchmark for code agent evaluation in CATArena.
-Learning Ability.
-The learning capability of a code agent captures its ability to leverage historical information and opponent behaviors to improve its own performance.
-Global Learning assesses an agent’s overall improvement in strategy quality.
-This metric evaluates the relative performance of agent $i$’s strategies against all strategies from all agents and rounds, and measures the average progress made compared to its initial baseline.
-It serves as the primary indicator of learning ability.
-Formally, for agent $i$, global learning is defined as:
-$L_{i}=\mathrm{average}_{n=2}^{N}\left(G_{i}^{n}-G_{i}^{1}\right),$
-where $G_{i}^{n}$ represents the global performance of agent $i$’s strategy from round $n$:
-$G_{i}^{n}=\mathrm{average}_{(i,n)\not=(j,m)}\left(W_{i,j}^{n,m}\right).$
-This metric captures the agent’s ability to learn and adapt over multiple rounds, reflecting its progress in a comprehensive competitive landscape.
-Counter-Adaptation measures an agent’s targeted learning ability, reflecting its capacity to achieve improved results against opponents in successive rounds.
-For agent $i$, the counter-adaptation score is defined as the average improvement in scores against other agents from round n1n-1 to round $n$ $(n\geq 2)$:
-$C_{i}=\mathrm{average}_{n=2}^{N}\left(A_{i}^{n}-B_{i}^{n-1}\right),$
-where the advance score $A_{i}^{n}$ and base score $B_{i}^{n-1}$ are defined as:
-$A_{i}^{n}=\mathrm{average}_{j\neq i}\left(W_{i,j}^{n,n-1}\right),B_{i}^{n-1}=\mathrm{average}_{j\neq i}\left(W_{i,j}^{n-1}\right).$
-Here, $A_{i}^{n}$ represents agent $i$’s average performance in round $n$ against the strategies submitted by other agents in the previous in round (n1n-1).
-The base score $B_{i}^{n-1}$ denotes agent $i$’s average performance in round n1n-1 against those same opponents.
-This comparison isolates the agent’s targeted adaptation from one round to the next.
-Self-improvement evaluates an agent’s capacity to genuinely enhance its strategies over successive rounds of development.
-This metric reflects whether newly developed strategies can consistently outperform the agent’s own previous versions.
-We quantify self-improvement by calculating the Pearson correlation Pearson (1896) between the round index and the agent’s average scores across rounds.
-For agent $i$, the self-improvement score is defined as
-$\mathrm{SI}_{i}=\mathrm{Pearson}\left([1,\cdots,N],\ [S_{i}^{1},\cdots,S_{i}^{N}]\right).$
-Here, $S_{i}^{n}$ denotes the average score of agent $i$’s strategy from round $n$ against its own strategies from other rounds
-$S_{i}^{n}=\mathrm{average}_{m\neq n}\left(W_{i}^{n,m}\right).$
-A higher self-improvement score indicates a stronger ability to iteratively refine and upgrade strategies throughout the development process.
-Generalizability.
-Generalizability measures an agent’s ability to comprehend and adapt to novel or altered game rules that differ from those encountered during training or prior experience.
-This metric specifically evaluates the agent’s capacity to generalize beyond previously seen environments, focusing on handling new or modified scenarios.
-For agent $i$, the generalizability score is defined as:
-$U_{i}=B_{i}^{1;\mathrm{Variants}}-B_{i}^{1;\mathrm{Standard}},$
-where $B_{i}^{1;\mathrm{Variants}}$ and $B_{i}^{1;\mathrm{Standard}}$ denote the base scores of agent $i$ in the first round under variant and standard rule settings, respectively.
-A higher value of $U_{i}$ indicates stronger generalizability, reflecting the agent’s ability to effectively develop and apply strategies for previously unseen tasks.
+The process begins with the initial development phase, which establishes the single-turn development capability of the code agent.
+In this stage, the agent receives a task specification $\mathcal{T}$, ranging from game rules to algorithmic problems, along with a minimal functional example.
+Operating without external guidance or historical data, the agent must comprehend the requirements and implement an initial solution $C^{(1)}$.
+Execution Feedback.
+Upon receiving the code submission, CATArena initiates the execution feedback process to drive the transition from static code generation and dynamic evolution.
+The system executes the submitted code within a sandboxed evaluation arena to generate quantitative performance scores.
+This phase aggregates execution results into a structured feedback set containing comprehensive tournament execution logs, a global leaderboard, and peer solutions.
+While the specific metrics and feedback modalities vary depending on the task type (as detailed in Section 3.2), the fundamental function remains consistent in providing the necessary signals for evolution.
+Iterative Evolution (Round $n>1$).
+In the iterative evolution phase, the goal of agents shifts from generation to optimization.
+Agents are provided with the task specification and the execution feedback derived from the previous round.
+By analyzing this information, agents are tasked to identify weaknesses in their previous implementations via self-reflection and assimilate superior approaches through peer-learning.
+The agent then synthesizes these insights to generate an evolved solution $C^{(n)}$.
+This cycle repeats for $N$ rounds, simulating the real-world engineering process where solutions are continuously refined based on testing and analysis of external implementations.
+3.2 Task Categorization and Instantiation
+To comprehensively evaluate the evolutionary potential of code agents, CATArena instantiates six diverse coding tasks.
+As demonstrated in Table 1, these tasks are categorized into two distinct types based on the nature of the feedback source, mirroring the duality of real-world software engineering where development is driven either by adversarial dynamics or distinct system efficiency targets: Competitive Tasks and Objective Optimization Tasks.
+Competitive Tasks.
+This category models scenarios where performance is inherently relative.
+Unlike static benchmarks, the feedback mechanism here is dynamic: a solution’s effectiveness depends entirely on the strength of its opponents.
+As the population evolves, the difficulty of winning increases, creating a non-stationary evaluation target that compels agents to continuously adapt to new counter-strategies.
+To capture this, we employ a Round-Robin tournament structure where the code implementations submitted by agents are executed against each other in pairwise or multi-agent matches.
+Crucially, the feedback provided to agents is not limited to their own match results.
+It encompasses the full tournament logs and peer solutions from the entire population.
+We instantiate four specific arenas: Gomoku, Texas Hold’em, Chess, and Bridge.
+This diversity, covering both symmetric perfect-information games and asymmetric imperfect-information settings, rigorously tests the agent’s ability to analyze complex interaction dynamics and refine strategies against an adaptive field.
+Objective Optimization Tasks.
+This category models scenarios where the primary goal is to optimize a static objective function.
+Here, the feedback is quasi-static.
+While the evaluation metrics are deterministic, the competitive standard is continuously raised by the population.
+Agents must treat functional correctness merely as a starting point, striving to approach theoretical limits to outperform peer solutions.
+We incorporate two specific benchmarks.
+First, OIBench (Zhu et al., 2025b) evaluates algorithmic efficiency using Time AUC, a metric that quantifies the volume of test cases passed within constrained time limits.
+Second, SWE-Perf (He et al., 2025) assesses repository-level engineering by challenging agents to optimize authentic GitHub codebases.
+It measures the performance speedup relative to the original implementation, highlighting the gap between functional code and expert-level optimization.
+3.3 Formalization and Metrics
+To provide a standard of unified evaluation across diverse tasks, we construct a global performance matrix $W$.
+Based on this matrix, we derive two core metrics to comprehensively assess both the agent’s static coding capability and its evolutionary potential.
+Performance Matrix Construction.
+Let $N$ be the total number of rounds and $M$ be the number of participating agents.
+We define the performance matrix $W\in\mathbb{R}^{(M\times N)\times(M\times N)}$, where $W_{i,j}^{n,m}$ denotes the performance score of Agent $i$ (from Round $n$) evaluated against Agent $j$ (from Round $m$).
+  • For competitive tasks, $W_{i,j}^{n,m}$ represents the win rate of agent $i$’s code in round $n$ against agent $j$’s code in round $m$.
+For symmetric settings, this is derived from pairwise round-robin matches.
+For asymmetric settings (e.g., Texas Hold’em), we utilize batch-based tournaments where scores reflect the win probability in the tournament.
+  • For objective optimization tasks, since agents operate independently.
+Thus, we simplify the notation $W_{i,j}^{n,m}$ to $W_{i}^{n}$, representing the scalar performance score of agent $i$ at iteration $n$.
+Specifically, for OIBench, we adopt their Time AUC metric (Zhu et al., 2025b), which quantifies the area under the curve of passed test cases against varying time constraints, normalized by the maximum achievable area.
+For SWE-Perf, we utilize the performance ratio (He et al., 2025), defined as the computational speedup of the optimized code relative to the original baseline, normalized by a target expert reference.
+Detailed formalization of the matrix $W$ in CATArena is provided in Appendix A.
+To mitigate randomness in game outcomes, every entry in the scoring matrix $W$ is estimated by repeated matches for fixed code submissions.
+Detailed repetition settings for each task are provided in Appendix B.
+Basic Performance $S_{\text{base}}$.
+This metric evaluates the agent’s static coding ability in a traditional single-turn setting.
+It quantifies the relative performance of the agent’s initial submission $C^{(1)}$ within the Round 1 population:
+$S_{\text{base},i}=\frac{1}{M-1}\sum_{j\neq i}W_{i,j}^{1,1}.$ (1)
+For objective tasks, this simplifies to the score of the first round $W_{i}^{1}$.
+$S_{\text{base}}$ serves as a baseline to reflect the agent’s zero-shot task-solving proficiency before any feedback is introduced.
+Evolutionary Capability $S_{\text{evo}}$.
+This metric is the primary indicator of evolutionary capability. Unlike static metrics, it quantifies the rate of improvement driven by feedback.
+We first calculate the global performance score $G_{i}^{n}$ for each round $n$.
+For competitive tasks, $G_{i}^{n}$ represents the relative strength of Agent $i$’s Round $n$ code when evaluated against the entire population of code versions ($M\times N$) produced across all rounds:
+$G_{i}^{n}=\frac{1}{MN-1}\sum_{(j,m)\neq(i,n)}W_{i,j}^{n,m}.$ (2)
+We use global average score for $G_{i}^{n}$ and give a brief explanation in Appendix C about the stability of this metric. For objective tasks, $G_{i}^{n}$ is simply the score $W_{i}^{n}$.
+$S_{\text{evo},i}$ is derived as the slope $k_{i}$ of the linear fit $G_{i}^{n}\approx k_{i}\cdot n+b_{i}$ using the Ordinary Least Squares (OLS) method:
+$S_{\text{evo},i}=\frac{\sum_{n=1}^{N}(n-\bar{n})(G_{i}^{n}-\bar{G}_{i})}{\sum_{n=1}^{N}(n-\bar{n})^{2}},$ (3)
+where $\bar{n}$ and $\bar{G}_{i}$ are the average round number and average global performance, respectively.
+A positive $S_{\text{evo}}$ indicates effective adaptation, with higher values signifying a faster rate of evolution.
+We employ OLS method primarily because development failures might occur in any round due to inherent limitations in their robustness.
+Compared to simpler alternatives such as total improvement (i.e., $G^{N}-G^{1}$), which are highly sensitive to single-round outliers, OLS provides a more robust estimation of the overall growth trend by mitigating the impact of transient failures.
+Importantly, we regard development failures as a failure of the evolutionary process itself, and these instances are explicitly factored into the OLS calculation to reflect the agent’s inability to maintain stable progress.
+Furthermore, as our experiments focus on the initial active evolution phase ($N\leq 4$), a linear approximation effectively captures the rate of evolution before performance saturation.
 4 Experiments
+In this section, we conduct comprehensive experiments to validate the effectiveness of the CATArena framework and analyze the evolutionary behaviors of code agents under the proposed iterative paradigm.
+Specifically, our evaluation aims to address the following four research questions:
+  • RQ1: Can CATArena separate evolutionary capability ($S_{evo}$) from static base performance ($S_{base}$)?
+  • RQ2: What dynamics do agents exhibit over iterations?
+  • RQ3: How do peer-learning and self-reflection affect evolution?
+  • RQ4: Is CATArena robust to task variants and extensible to new tracks?
+Table 2: Main Results on CATArena.
+We report the Base Performance ($S_{\text{base}}$) and Evolutionary Capability ($S_{\text{evo}}$) of Minimal and Commercial Agents across six tasks.
+Best results in each column are bolded.
+Note: Commercial agents are excluded from SWE-Perf due to incompatibility with the Docker environment.
+Agents Competitive Objective
+Gomoku Hold’em Bridge Chess OIBench SWE-Perf
+$S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$
+Minimal Claude 0.88 +0.030 0.58 +0.072 0.79 +0.062 0.90 -0.061 0.32 +0.040 0.02 -0.0016
+DeepSeek 0.23 +0.014 0.01 +0.006 0.00 0.000 0.00 0.000 0.10 -0.029 0.02 -0.0011
+Doubao 0.33 -0.046 0.04 -0.011 0.20 +0.013 0.58 -0.214 0.41 +0.041 0.00 +0.0002
+Gemini 0.25 +0.156 0.01 +0.024 0.90 -0.083 0.58 -0.158 0.40 +0.022 0.01 +0.0012
+GPT 0.48 +0.012 0.16 +0.025 0.47 -0.095 0.38 -0.044 0.41 +0.031 0.02 +0.0005
+Qwen3 0.85 -0.196 0.20 +0.019 0.65 +0.003 0.58 -0.187 0.10 +0.099 0.01 +0.0036
+Commercial Best ADK Agent 0.00 +0.176 0.07 +0.057 0.25 +0.191 0.91 -0.072 0.45 +0.049 - -
+Claude Code 0.78 -0.035 0.01 +0.027 1.00 +0.056 0.56 -0.178 0.43 +0.072 - -
+Codex 0.47 +0.103 0.72 -0.008 0.75 +0.062 0.38 +0.011 0.46 +0.186 - -
+Gemini CLI 0.31 -0.030 0.13 -0.017 0.01 -0.056 0.38 +0.099 0.63 -0.026 - -
+Qwen Code 0.94 +0.035 0.07 +0.025 0.49 +0.053 0.28 -0.144 0.35 +0.009 - -
 4.1 Experimental Setups
 Participants.
-In our experiment, we employ three types of agents: (1) Minimal Agents (LLM + ADK Framework) [^2] https://github.com/AGI-Eval-Official/Minimal-CodeAgent: A baseline agent developed with the Agent Development Kit (ADK) Python toolkit.
-We provide essential tools, including file manipulation, bash scripting, and Python execution, for the ADK code agent.
-On this foundation, we integrate state-of-the-art LLMs to systematically compare their core competencies as code agents for strategy implementation.
-(2) Commercial Code Agents: State-of-the-art and commercial CLI-based agents (e.g., Claude Code, CodeX, Gemini-CLI, Qwen-Coder) are included for benchmarking.
-These agents feature advanced integration with various command-line interfaces, tools, and LLMs, resulting in enhanced overall capabilities.
-These agents serve as leading solutions in code agent development and provide valuable reference points for future research.
-(3) LLM-Player: In this control setting, LLMs directly output game moves without generating code.
-For each turn, the LLM receives the game rules, current state, and history, and returns the next action.
-This approach is specifically designed to assess the inherent strategic and reasoning capabilities of LLMs.
-Detailed agent parameter settings and model selections are presented in Appendix C.
-Table 2: Agent Specifications and Open-Source Status.
-Agent Type Agent Framework Model Agent OSS LLM OSS
-Minimal basic code tools with ADK framework DeepSeek-3.1 (DeepSeek-AI, 2024)
-Qwen3-Coder-480B (Team, 2025c)
-Doubao-Seed-1.6 (Team, 2025d)
-GPT-5 (OpenAI, 2025b)
-Claude-4-Sonnet (Anthropic, 2025a)
-Gemini-2.5-pro (Team, 2025a)
-Commercial Gemini-CLI (Google, 2025) Gemini-2.5-pro (Team, 2025a)
-Claude-Code (Anthropic, 2025b) Claude-4/3.7 Hybrid (Anthropic, 2025a)
-CodeX (OpenAI, 2025a) GPT-5 (OpenAI, 2025b)
-Qwen-Coder (Team, 2025b) Qwen3-Coder-480B (Team, 2025c)
-Other LLM-Player Agents’s Corresponding LLM N/A N/A
-Tournaments.
-All experiments are conducted under two main tournament settings: (1) a comparison among minimal agents equipped with different LLMs ($T_{1}=6$), and (2) a comparison between the best-performing minimal agent and a set of commercial code agents ($T_{2}=5$).
-To reduce the impact of randomness on strategy generation, each tournament is repeated for four times, and all reported metrics are averaged over the four runs.
-Each tournament consists of $N=4$ rounds of iterative development.
-To further mitigate stochastic effects in competition outcomes, every entry in the scoring matrix $W$ is estimated by repeated matches.
-Detailed tournament prompts are listed in Appendix K.
-We report the detailed scoring policy, generation configs, and repetition experiments in Appendix D.
-It is noteworthy that agents tend to generate different codes in repeated experiments, but their rankings are relatively stable.
-4.2 Main Results
-(a) Trends of global performance scores $G_{i}^{n}$.
-(b) Quantitative comparison of learning abilities.
-Figure 2: Visualization of agents’ learning patterns and scores.. For clarity, we use family names to represent LLM models instead of their full names. The results for other games are in Appendix E.
-Table 3: Main LeaderBoard of CATArena.
-We conduct two groups of tournaments between minimal agents (rank from 1 to 6) and commercial agents (rank from 1 to 5), and report the average ranking across all four tasks.
-Metrics include S.C.
-(Strategy Coding), G.L.
-(Global Learning), and G.A.
-(Generalizability).
-Agents Standard Variant G.A. $\downarrow$
-S.C. $\downarrow$ G.L. $\downarrow$ S.C. $\downarrow$ G.L. $\downarrow$
-Minimal
-Claude-4-Sonnet 1.25 2.5 1.75 2.75 5.00
-DeepSeek-Chat 5.75 2.75 4.25 2.75 2.75
-Doubao-Seed 3.75 4.75 3.75 4.50 2.75
-Gemini-2.5-Pro 3.25 3.75 3.25 2.75 3.25
-GPT-5 3.75 3.50 3.00 3.75 2.25
-Qwen3-Coder 2.25 3.75 3.00 4.5 4.75
-Commercial
-best ADK 3.25 2.25 2.00 3.75 2.50
-Claude-Code 2.50 3.75 2.50 2.75 3.25
-CodeX 2.25 2.75 3.00 3.00 3.25
-Gemini-CLI 3.50 2.25 3.00 4.00 2.00
-Qwen-Coder 3.00 3.75 4.00 1.25 3.25
-Learning Ability.
-Figure 2(a) visualizes the global performance scores $G_{i}^{n}$ revealing overall trends in agent strategies across iterations.
-Some agents, such as minimal agents driven by claude, exhibit a clear upward trajectory over multiple rounds, demonstrating strong learning capability.
-However, the performance of most agents remains unstable, and no obvious trend is observed.
-To explicitly illustrate the learning ability of each agent, we design a quantitative scoring method for global learning and introduce two additional learning modes: counter-adaption and self-improvement.
-The quantitative results for these three abilities are presented in Figure 2(b).
-Compared to Figure 2(a), these scores offer a more intuitive comparison of the agents’ learning strengths.
-This decomposition provides deeper insights into the mechanisms underlying agent learning.
-When both counter-adaption and self-improvement scores are positive, it indicates that the agent can effectively learn from both its opponents and itself, resulting in a positive global learning score.
-Compared to minimal agents, commercial agents exhibit stronger learning capabilities.
-Learning ability results for other games and a case study of strategy iteration mechanisms are provided in Appendix E.
-Through an analysis of the consistency of agent actions in endgame states, we observe that a majority of agents indeed learn from the code generated in the previous round, leading to increasing consistency.
-This phenomenon is most prominent in the Hold’em environment, possibly because its strategies are relatively simple and easier to learn.
-Ability Evaluation.
-We conduct two sets of tournaments: (1) minimal agents equipped with different LLMs, (2) the best-performing minimal agent against commercial code agents.
-For each tournament, we report the average ranking and scores for three core agent capabilities.
-The main leaderboard and main results of CATArena are summarized in Table 3 and Table 4.
-These results reveal the following key observations:
-Observation 1: The performance gap among LLMs is more pronounced in minimal agents compared to commercial agents.
-Table 3 shows that Claude-4-Sonnet achieves the highest score among minimal agents, while the rankings of other LLMs are more dispersed.
-In contrast, commercial agents driven by the same LLMs exhibit much closer average rankings, with all agents scoring around 2.5 out of 5, indicating a reduced performance gap.
-Moreover, commercial agents demonstrate performance levels similar to the best-performing minimal agent.
-This suggests that the underlying agent framework can significantly influence how effectively an LLM’s capabilities are utilized, as commercial agents are often optimized for specific models.
-Observation 2: The participating agents display different ranking orders across various capabilities.
-The tournament results reveal that the relative rankings of agents change depending on the specific core ability being tested.
-The ranking of these abilities provides a decomposition of end-to-end performance, offering insights for further optimization of both LLMs and agent frameworks.
-Observation 3: Agents exhibit varied performance distributions across different tasks.
-The results indicate that agents’ performance is not uniformly distributed across all tasks, which is mainly attributed to the distinct nature and difficulty of the four tasks.
-In the variant tasks, the performance gap among agents is more pronounced, likely because the game rules and strategies are less familiar to agents.
-[…]
-Games Gomoku Hold’em Bridge Chess
-Standard Variant Standard Variant Standard Variant Standard Variant
-[…]
-Minimal
-Claude-4-Sonnet 0.88 -0.447 0.78 -0.156 -0.14 0.58 0.118 0.13 0.110 -0.45 0.79 0.174 1.0 0.047 0.005 0.90 -0.170 0.65 0.018 -0.55
-Deepseek-Chat 0.23 0.027 0.38 0.077 0.13 0.01 0.010 0.00 -0.022 -0.01 0 0 0.10 0 0.10 0 0 0.10 0.049 0.10
-Doubao-Seed 0.33 -0.192 0.72 -0.302 0.46 0.04 -0.035 0 0 -0.04 0.2 -0.033 0.45 -0.516 0.40 0.58 -0.337 0.10 0.034 -0.46
-Gemini-2.5-Pro 0.25 -0.066 0.00 0.173 -0.12 0.01 0.020 0.00 0.078 -0.01 0.90 -0.195 0.60 -0.049 -0.30 0.58 -0.147 0.90 0.003 0.46
-GPT-5 0.48 0.062 0.76 -0.019 0.18 0.16 0.102 0.87 -0.050 0.71 0.47 -0.095 0.10 0.293 -0.02 0.38 -0.525 0.45 -0 0.24
-Qwen3-Coder 0.85 -0.523 0.36 -0.089 -0.50 0.20 0.038 0.00 0.003 -0.20 0.65 0.032 0.76 -0.230 -0.19 0.58 -0.187 0.80 -0.532 0.21
-Commercial
-best ADK 0 0.075 0.75 -0.022 0.88 0.07 0.073 0.46 0.030 0.39 0.25 0.295 0 0.361 -0.25 0.91 -0.110 1.00 -0.342 -0.47
-Claude-Code 0.78 -0.322 0.66 0.194 -0.28 0.01 0.100 0 0.105 -0.001 1.00 -0.240 0.93 -0.139 -0.04 0.56 -0.158 0.44 -0.226 0.03
-CodeX 0.47 0.454 0.69 -0.095 0.34 0.72 0.050 0.17 0.067 -0.55 0.75 0.098 0.50 0.285 -0.25 0.38 0.033 0.34 0.064 0.09
-Gemini-CLI 0.31 0.260 0.19 0.172 0.0 0.13 0.050 0.37 -0.007 0.24 0.01 0.254 0.83 -0.286 0.79 0.38 0.395 0.38 -0.154 0.16
-Qwen-Coder 0.94 -0.054 0.22 -0.530 -0.94 0.07 0.058 0 0.105 -0.007 0.49 0.157 0.25 0.556 -0.25 0.28 -0.204 0.34 0.039 0.19
-4.3 Effectiveness of CATArena
-We design a series of experiments to demonstrate the effectiveness of CATArena.
-Comparison between Agents and LLM-Players.
-The primary task in CATArena is strategy coding, which relies on the underlying coding capabilities of LLMs.
-We posit that reasoning over code to develop strategies is fundamentally different from direct reasoning during gameplay.
-To validate this distinction, we compare agent-developed strategies with the LLM-Player baseline (see Appendix F).
-Our results show that current agents primarily implement simple rule-based algorithms, indicating substantial room for advancement in agents’ strategy coding abilities.
-CATArena fully leverages this non-saturation, enabling sustainable iterative peer-learning.
-To further analyze the similarities and differences between agent-implemented code strategies and those of LLM-Players, we ask agents’ code and LLM-Player to select the next action on endgame states.
-Figure 3 illustrates the action consistency between agents’ code and LLM-Players in Chess.
-Surprisingly, the strategies encoded in agent code differ significantly from those inferred directly by the LLM, even if they are from the same model.
-Meanwhile, strategies produced by different agents and different LLMs also show notable similarities.
-This indicates that strategy coding and reasoning in LLMs are distinct capabilities.
-We report results of other tasks are demonstrated in Appendix G.
-CATArena evaluates the strategy coding ability of agents rather than their reasoning ability, thereby filling a gap in previous benchmarks.
-The relationship between strategy coding ability and LLM-based strategy reasoning remains unclear and requires further investigation.
-Figure 3: Action consistency between agents’ code and LLM-Players on Chess endgames.
-Collective Learning Trends Among Agents.
-Table 5: Collective learning trends of agents across different tasks in CATArena.
-$\text{DIS}_{\text{range}}$ and $\text{DIS}_{\text{std}}$ represent the Pearson correlation coefficients of the range and standard deviation of agent performance scores over four rounds, reflecting the similarity and dispersion of agent strategies.
-$\text{Trend}_{\text{mean}}$ denotes the Pearson correlation between the mean agent performance and the round number, indicating the overall trend of group improvement.
-Gomoku Hold’em Bridge Chess
-StdV. VarV. StdV. VarV. StdV. VarV. StdV. VarV.
-$\text{DIS}_{\text{std}}$ -0.05 0.15 -0.81 -0.80 -0.82 -0.57 0.55 -0.04
-$\text{DIS}_{\text{range}}$ -0.16 0.44 -0.80 -0.76 -0.54 -0.33 -0.08 0.16
-$\text{Trend}_{\text{mean}}$ 0.42 -0.02 0.75 0.67 0.24 -0.10 -0.74 -0.79
-We analyze the collective learning dynamics of agents across tasks, as presented in Table 5.
-The metrics $\text{DIS}_{\text{range}}$ and $\text{DIS}_{\text{std}}$ report the Pearson correlation between the standard deviation and range of agent performance scores over four rounds ($B_{i}^{n},n=1,2,3,4$).
-The higher similarity in performance scores (i.e., $\text{DIS}>0$) suggests that agents can learn effective strategies more readily, indicating lower task difficulty.
-Based on these results, the relative difficulty ranking of tasks in CATArena is Chess >> Gomoku >> Bridge >> Hold’em.
-[…]
-Additional Results.
-CATArena’s iterative peer-learning framework is easily extensible to new tasks for evaluating other fundamental agent abilities.
-We demonstrate this by introducing a Machine Learning (ML) track and multi-lingual track, with experimental results provided in Appendix H and Appendix I, respectively.
-Experimental results indicate that current agents still exhibit substantial potential for improvement.
-As agents continue to advance, the open-ended task design and peer-learning evaluation framework of CATArena ensure that systematic assessment can be sustained over time.
-We also report agent cost in terms of token usage, time consumption, and generated code statistics for each agent, in Appendix J.
-Notably, Claude-4-Sonnet utilizes the most tools and tokens, and also develops a significantly larger amount of code.
-In contrast, GPT-5 achieves the best balance between token usage and performance.
-Efficient utilization of tokens and external tools remains an important research direction to advance the capabilities of LLM agents.
+We evaluate two categories of agents to disentangle model reasoning from system engineering: (1) Minimal Agents: To isolate the intrinsic evolutionary potential of foundation models, we implement a standardized minimal agent framework using Google Agent Development Kit (ADK) [^2] https://github.com/google/adk-python.
+This minimalist setup provides only core tools (e.g., file I/O and scripting), ensuring that performance differences are primarily attributable to the underlying LLM rather than complex scaffolding (Kim et al., 2023; Shihipar, 2026).
+We select a diverse set of API-based and open-weight models as backbones.
+(2) Commercial Code Agents: We include leading CLI-based commercial agents to establish high-performance industry baselines.
+These systems feature proprietary toolchains and represent the current state-of-the-art in agentic development.
+For brevity, we refer to agents by their model family names; full specifications are provided in Appendix D.
+Evaluation Protocol.
+All experiments are conducted within two distinct cohorts to facilitate intra-group evaluation: (1) A comparison among minimal agents equipped with various LLMs, and (2) A comparison between the best-performing minimal agent and a suite of commercial code agents.
+This setup ensures that agents are evaluated against peers within their respective categories to maintain a fair and reliable assessment of evolutionary capability.
+Repetition and Stability.
+Due to the high computational and token costs of long-term evolution, the main evolutionary experiments are conducted as a single chain.
+To validate the reliability of this approach, we conducted a separate stability analysis on the initial rounds (Rounds 1 and 2), repeating the code generation process four times.
+We observe that while the specific code implementations vary across runs, the relative performance rankings among agents remain highly consistent (see Appendix E for analysis).
+4.2 Results of Evolutionary Evaluation (RQ1)
+Figure 2: Evolutionary Trajectories across Iterations. Solid lines represent the actual $G$ Score, dashed lines indicate the linear regression fits used to calculate $S_{\text{evo}}$.
+Overall Performance Comparison.
+Table 2 summarizes the Base Performance ($S_{\text{base}}$) and Evolutionary Capability ($S_{\text{evo}}$) across all agents.
+We also report statistics on HTTP error rates in Appendix F.
+The results yield three key observations: (1) We observe a weak correlation between $S_{\text{base}}$ and $S_{\text{evo}}$ across all tasks.
+High $S_{\text{base}}$ does not guarantee effective evolution.
+For instance, while Claude-4 dominates the initial round in Chess ($0.90$), it suffers significant regression (0.061-0.061) during evolution.
+Conversely, Gemini-2.5-Pro demonstrates a low-start, high-growth trajectory in Gomoku ($0.25\rightarrow+0.156$).
+While it starts with a deficit, its high $S_{\text{evo}}$ indicates the ability to rapidly absorb feedback.
+(2) Evolutionary dynamics are inherently task-dependent.
+We observe a distinct complexity barrier in the evolution of agents.
+In high-difficulty environments such as Chess and SWE-Perf, most agents exhibit negligible or negative $S_{\text{evo}}$.
+This indicates that when the reasoning gap between the current implementation and the optimal solution exceeds a certain threshold, agents struggle to extract actionable insights from feedback.
+(3) Commercial frameworks primarily enhance static generation rather than evolution ability.
+While commercial agents frequently outperform the Best ADK Agent in $S_{\text{base}}$, they do not exhibit a decisive advantage in $S_{\text{evo}}$.
+This disparity underscores that evolutionary capability is a distinct dimension of intelligence, largely unaddressed by current commercial scaffolding, and represents a critical frontier for future agentic system research.
+Evolutionary Trajectories.
+Figure 2 visualizes the round-wise performance dynamics of global performance $G^{n}$, highlighting three critical phenomena.
+First, we observe significant performance oscillation across iterations.
+This volatility indicates that current agents lack robust self-evaluation mechanisms, often discarding valid strategies for inferior ones during the evolutionary process.
+Second, despite this instability, evolution can drive substantial performance leaps.
+For instance, in OIBench, Qwen (blue line) advances from a last-place initialization ($G_{1}=0.10$) to the top-performing tier ($G_{4}=0.43$) through iterative refinement.
+This trajectory confirms that static zero-shot benchmarks may underestimate an agent’s true potential, which can be significantly amplified through iterative feedback.
+Third, the visualization validates the reliability of our metric design.
+As indicated by the dashed regression lines, $S_{\text{evo}}$ effectively captures the global optimization trend, proving robust to transient performance drops caused by exploratory failures typical of stochastic code generation.
+The results for other tasks are detailed in Appendix G.
+4.3 Temporal Dynamics of Evolution (RQ2)
+We investigate the temporal dynamics of agent evolution by analyzing population statistics across task types and extending the evolutionary horizon to $N=7$ rounds.
+Population Evolution Patterns.
+Table 3 characterizes the evolutionary trends using two metrics derived from Pearson correlation coefficients: $\text{Trend}_{\text{mean}}$, representing the correlation between round number and average population performance, and $\text{DIS}_{\text{std}}$, the correlation between round number and score dispersion (standard deviation).
+The results reveal that the population structure is heavily dictated by the feedback mechanism.
+Objective optimization tasks tend to drive divergence ($\text{DIS}_{\text{std}}>0$), whereas competitive tasks foster convergence ($\text{DIS}_{\text{std}}<0$).
+This disparity arises because objective feedback rewards strong agents capable of approaching theoretical limits, amplifying the gap, while the open-source nature of competitive tournaments allows weaker agents to mimic successful strategies, homogenizing the population.
+Regarding overall performance, the majority of tasks exhibit a positive $\text{Trend}_{\text{mean}}$, confirming that iterative refinement generally drives collective improvement.
+The only exception is Chess ($\text{Trend}_{\textbf{mean}}<0$), where the excessive reasoning complexity overwhelms the feedback loop, leading to a collapse rather than evolution.
+Table 3: Evolutionary Population Dynamics.
+Pearson correlation coefficients between iteration rounds and population performance metrics: $\text{DIS}_{\text{std}}$ (Standard Deviation) and $\text{Trend}_{\text{mean}}$ (Mean Score).
+Task Type Task $\text{DIS}_{\text{std}}$ $\text{Trend}_{\text{mean}}$ Pattern
+Objective OIBench +0.42 +0.79 Divergence
+SWE-Perf +0.20 +0.20 Weak trend
+Competitive Gomoku -0.05 +0.42 Convergence
+Hold’em -0.81 +0.75 Strong convergence
+Bridge -0.82 +0.24 Moderate convergence
+Chess +0.55 -0.74 Oscillation
+Figure 3: Long-term Evolutionary Dynamics ($N=7$). Top: The Peak Performance Gain $\Delta\text{Max}$ across iterations. Bottom: The trajectories of Global Performance $G^{n}$ for individual agents.
+Long-term Evolution.
+We extend the evaluation to $N=7$ for Gomoku and OIBench to investigate long-term evolution, visualizing the results in Figure 3.
+We first analyze the Peak Performance Gain ($\Delta\text{MAX}$), defined as the sum of performance increments across all models between consecutive rounds, which quantifies the evolutionary vitality.
+We observe that $\Delta\text{MAX}$ is significantly higher during the initial rounds ($N<4$) compared to the later stages ($N>4$).
+This decay justifies our selection of a four-round window for the main experiment, as it effectively captures the most active phase of evolution.
+Furthermore, the global performance $G^{n}$ trajectories in the bottom panel reveal a clear pattern of stratification.
+While early rounds exhibit frequent rank crossings, the rankings stabilize in the long term with significantly fewer intersections.
+Most agents eventually trend towards saturation, failing to break through performance ceilings in the later rounds.
+This suggests that while iterative feedback optimizes strategies, the asymptotic performance remains bounded by the intrinsic reasoning limits of the underlying models.
+4.4 Mechanism of Evolution (RQ3)
+To deconstruct the black box of agent evolution, we categorize the evolutionary process into two distinct mechanisms: Self-Reflection, where the agent introspectively analyzes its own execution logs to fix bugs and refine logic; and Peer-Learning, where the agent extrospectively analyzes external solutions to adopt superior strategies or algorithms.
+Table 4: Ablation Study on peer-learning.
+Comparison of performance with full feedback ($G_{2}$) versus self-reflection only ($G_{2}^{\prime}$).
+Bold values indicate improvement over the initial round ($G_{1}$).
+The Pattern column denotes the effective mechanism: S indicates effective Self-Reflection ($G_{2}^{\prime}>G_{1}$), P indicates effective Peer-Learning ($G_{2}>G_{2}^{\prime}$).
+Agent Gomoku OIBench
+$G_{1}$ $G_{2}^{\prime}$ $G_{2}$ Pattern $G_{1}$ $G_{2}^{\prime}$ $G_{2}$ Pattern
+Claude 0.814 0.520 0.951 ✗P 0.318 0.204 0.410 ✗P
+DeepSeek 0.363 0.265 0.333 ✗✗ 0.097 0.105 0.000 S✗
+Doubao 0.716 0.039 0.598 ✗✗ 0.408 0.458 0.000 S✗
+Gemini 0.225 0.794 0.118 S✗ 0.403 0.346 0.301 ✗✗
+GPT 0.500 0.480 0.020 ✗✗ 0.413 0.470 0.423 S✗
+Qwen 0.833 0.814 0.618 ✗✗ 0.101 0.084 0.398 ✗P
+Ablation Study.
+To isolate the contribution of two evolution mechanisms, we compare agent performance with full feedback ($G_{2}$) against a self-reflection-only baseline ($G_{2}^{\prime}$).
+We omit the symmetric ablation (peer-only without logs) as it reduces agents to blind imitation lacking performance grounding.
+As detailed in Table 4, the results reveal a critical limitation: current agents struggle to synergize peer code and execution logs.
+Instead of benefiting from a unified feedback loop, agents typically exhibit a dependency on a single source.
+Specifically, the majority of models still rely predominantly on self-reflection for improvement, often treating peer information as noise or distraction.
+Only select high-capability agents (Claude and Qwen) demonstrate the ability to effectively leverage peer-learning to achieve performance breakthroughs.
+This lack of synthesis capability points to a key direction for future research: developing mechanisms that enable agents to robustly reconcile internal diagnostics with external knowledge.
+Code Consistency Analysis.
+To further dissect the behavioral mechanisms underlying evolution, we quantify the extent to which agents rely on self-refinement versus peer-adoption.
+For each agent’s Round 2 submission, we compute its similarity to its own Round 1 code (X-axis, self-consistency) and the maximum similarity to any peer’s Round 1 code (Y-axis, peer-adoption).
+For competitive tasks, similarity is measured by behavioral alignment in endgame positions; for objective tasks, we employ CodeBLEU (Ren et al., 2020) to assess syntactic and semantic overlap.
+As illustrated in Figure 4, most points cluster near the diagonal, indicating a balanced use of both mechanisms.
+However, three extreme failure modes emerge: (1) Blind Copying (high $y$), where Doubao and DeepSeek in Hold’em completely abandon their logic to mimic peers; (2) Stubborn Stagnation (high $x$), exemplified by DeepSeek in Gomoku, which refuses to adapt despite poor performance; and (3) Chaotic Reconstruction ($x,y<20$), where agents discard both internal and external priors, leading to unstable solutions.
+These extremes highlight a critical inability of current models to dynamically balance exploration and exploitation.
+To provide a more granular view of these behavioral patterns, we present the complete pairwise similarity matrices for all tasks in Appendix H.
+Figure 4: Code Evolution Patterns.
+Each point represents an agent’s Round 2 submission.
+The X-axis measures similarity to its own Round 1 code, while the Y-axis measures the maximum similarity to any peer’s Round 1 code.
+Case Study.
+Development logs reveal a structured refinement process driven by peer-learning and self-reflection (See Appendix I).
+Agents actively utilize tools such as Grep to extract win statistics from tournament reports, identifying top-performing peers to study and synthesize their logic.
+For instance, in Texas Hold’em, Claude transitioned from a basic heuristic formula to a sophisticated opponent-aware strategy after analyzing the aggressive betting patterns of high-ranking competitors.
+This peer-learning is complemented by explicit self-reflection, where agents benchmark their own win rates against rivals and plan targeted improvements in modules such as risk management or search depth.
+These behaviors show that agents evolve through deliberate optimization rather than stochastic mutations.
+4.5 Robustness and Extensibility of CATArena (RQ4)
+Generalization to Unseen Variants.
+To verify that CATArena evaluates genuine adaptive capabilities rather than the retrieval of memorized solutions, we conducted experiments on unseen variants for five tasks.
+These variants introduce rule perturbations, such as inverting hand rankings in Hold’em, that invalidate standard strategies present in the agents’ pre-training data.
+Detailed results are provided in Appendix J.
+As shown in Table 11, while the variance in zero-shot base scores $S_{\text{base}}$ increases significantly due to non-robust agents failing to generate valid strategies, the evolutionary scores $S_{\text{evo}}$ maintain their discriminative power.
+This consistency confirms that CATArena robustly measures the intrinsic evolutionary ability of agents, independent of data contamination or specific task familiarity.
+Extensibility to Diverse Tracks.
+CATArena is designed as a modular framework, allowing seamless integration of new environments and evaluation tracks.
+To demonstrate this extensibility, we incorporated three additional experimental tracks: a Multilingual Track (Python, C++, Java, JS, Go), a Machine Learning (ML) Track, and a Complex Task Track (Pommerman).
+Detailed results are provided in Appendices K, L, and M, respectively.
+We observe that advanced tracks pose significant challenges for SoTA agents.
+Most agents exhibit low $S_{\text{base}}$ and unstable evolutionary trajectories, similar to the oscillation patterns observed in Chess.
+This suggests that while the framework is ready for broader evaluation, current models struggle to generate robust starting points for evolution in these domains.
+As agent capabilities advance, CATArena will scale to these harder tasks to continue evaluating the frontier agents.
 5 Conclusion
-In this work, we address two fundamental challenges in LLM agent evaluation: the need for systematic measurement of learning ability, and the tendency of traditional benchmarks to become saturated as agent capabilities improve.
-To this end, we propose an iterative peer-learning-based competitive framework, enabling agents to continually revise and enhance their strategies through dynamic interaction and feedback.
-Building on this, we introduce CATArena, a tournament-style benchmark featuring open-ended and rankable board and card games.
-CATArena provides an environment with unlimited potential for agent improvement potential and extensible evaluation across new domains.
-Experimental results demonstrate that our framework reliably assesses core agent abilities, particularly learning ability and strategy coding, while ensuring stability and scalability.
-The open and flexible architecture of CATArena supports ongoing research and benchmarking for future intelligent agents.
-Limitations.
-The current evaluation in CATArena is limited to four games, which primarily assess agents’ learning ability and strategy coding.
-These scenarios do not encompass the full spectrum of potential LLM agent capabilities.
-In future work, we plan to introduce a wider variety of more complex tasks to evaluate agents’ learning and other abilities from different perspectives.
-Appendix A Tournament Format and Scoring System
-Table 6: Configs of tournament on four games.
-Environment Code Agents LLM-player
-Gomoku
-Board size: $\times15$
-Number of pairwise matches: 4 $\times$ 2
-Swap black and white pieces after each match
-Maximum time per move: 10 s
-Board size: $\times15$
-Number of pairwise matches: 2 $\times$ 2
-Swap black and white pieces after each match
-Maximum time per move: 600 s
-Texas Hold’em
-Max players: 12
-Rounds: 100
-Random shuffle seat after each round
-Initial chips: 2000
-Blind increase every 24 hands
-Max hands per round: 720 or until winner decided
-Maximum time per move: 3 s
-Max players: 12
-Rounds: 10
-Random shuffle seat after each round
-Initial chips: 2000
-Blind increase every 24 hands
-Max hands per round: 720 or until winner decided
-Maximum time per move: 1000 s
-Bridge
-Number of pairwise matches: 12 $\times$ 2
-Swap directions of open/closed rooms
-Use same deck for each pair of match
-Maximum time per move: 10 s
-Number of pairwise matches: 12 $\times$ 2
-Swap directions of open/closed rooms
-Use same deck for each pair of match
-Maximum time per move: 200 s
-Chess
-Number of pairwise matches: 8 $\times$ 2
-Swap black and white pieces after each match
-Maximum moves per game: 200
-Maximum time per move: 10 s
-Number of pairwise matches: 2 $\times$ 2
-Swap black and white pieces after each match
-Maximum moves per game: 200
-Maximum time per move: 600 s
-We list the basic settings for each games in Table 6.
-For each game, we ensure that the number of pairwise matches among code agents allows the final results to stabilize, i.e., for each game, the L1-norm fluctuation of the scoring matrix $W$ is less than 5%.
-For LLM-players, since their reasoning time is relatively long for most games, we reduce the number of repeated experiments.
-Note that our comparison with LLM-players is only a qualitative analysis of the differences between LLM-player and corresponding coding agent.
-The exploration of LLM-players’ results is not the focus of this paper; refer to the main text for details.
-Appendix B Evaluation Metric Calculation
-Table 7: Scoring rules of tournament on four games.
+In this paper, we present CATArena, a comprehensive framework that shifts the evaluation of code agents from static correctness to dynamic evolutionary capability.
+By simulating the iterative nature of real-world software engineering within a competitive tournament structure, CATArena quantifies the ability of agents to leverage execution feedback and peer knowledge for continuous optimization.
+Through extensive evaluations on CATArena, we benchmark the evolutionary capabilities of state-of-the-art LLM code agents, confirming that evolutionary potential is a distinct dimension from single-turn generation capability.
+Our in-depth analysis further dissects these evolutionary patterns, revealing the specific contributions of self-reflection and peer-learning mechanisms to agent growth.
+Moreover, we demonstrate that evolutionary trajectories are highly task-dependent; CATArena’s diverse spectrum of competitive and objective tasks proves essential for capturing these nuanced behavioral shifts.
+In the future, CATArena will integrate more tasks and serve as a highly extensible testbed.
+Impact Statement
+This paper presents CATArena, a benchmark for evaluating the evolutionary capabilities of code agents.
+Our goal is to advance the reliability and efficiency of automated software engineering.
+There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
+Appendix A Evaluation Metric Calculation
+Table 5: Scoring rules of $W$ matrix calculation on six tasks.
 Environment Scoring Metric
@@
 Chess Pairwise match scoring: Win = 1 point,  Draw = 0.5 point,  Lose = 0 point.
-To evaluate the basic capabilities of LLM agents, we define metrics for each applied game. Let $N$ be the number of rounds and $K$ be the number of participating agents. We construct a matchup matrix
-$W\in\mathbb{R}^{(N\cdot K)\times(N\cdot K)},$
+OIBench TimeAUC per problem: CDF of $\log_{10}(t/T_{\mathrm{ref}})$ AUC over $x_{\min},1]$, normalized by reference AUC.
+SWE-Perf Performance: mean significant speedup across efficiency tests (Mann-Whitney U, p<0.1p{<}0.1) vs baseline.
+To provide a unified evaluation standard across diverse tasks, we construct a global performance matrix $W$.
+Scoring rules for the tasks are summarized in Table 5.
+Let $N$ be the number of rounds and $M$ be the number of participating agents.
+We construct a matchup matrix
+$W\in\mathbb{R}^{(N\cdot M)\times(N\cdot M)},$
 where the generic element $W_{i,j}^{n,m}$ denotes the score when agent $i$ from round $n$ plays against agent $j$ from round $m$.
@@
 Diagonal entries $(n,i)=(m,j)$ are ignored.
-For asymmetric games, pairwise results are not feasible; instead, batch-based tournaments are used and the score matrix records the win rates of multi-agent matches.
+For Hold’em, pairwise results are not feasible; instead, batch-based tournaments are used and the score matrix records the win rates of multi-agent matches.
 For each batch, we obtain a single score group $W_{i_{1},i_{2},\ldots,i_{BS}}^{n_{1},n_{2},\ldots,n_{BS}}$, where $BS$ is the batch size.
 We conduct three types of experiments to accommodate different metric calculations: (1) $W_{i}^{1,2,\ldots,N}$, where the same agent’s strategies from $N$ rounds compete against each other, used to compute self-improvement metrics $S_{i}^{n}$; (2) $W_{i_{1},i_{2},\ldots,i_{T}}^{n}$, where all agents in the same round compete, used to calculate the base score $B$; (3) All $N\times T$ agent strategies are randomly shuffled and grouped for competition (with $BS=12$ in our experiments), used to compute both the global score $G_{i}^{n}$ and advanced score $A_{i}^{n}$.
-Scoring rules for the four games are summarized in Table 7.
-Appendix C Generation Configs
-For all LLMs used in our work, we set temperature to be $0.1$, max token identical to their official APIs’ setting.
-We set top-p to $1.0$, Top-k to be $100$, and presence penalty to be default to the API.
+For objective tasks, we simplify the notation $W_{i,j}^{n,m}$ to $W_{i}^{n}$.
+Appendix B Tournament Format and Repetition
+Table 6: Configs of tournament on six tasks.
+Task Type Environment Code Agents
+Competitive Gomoku
+Board size: $\times15$
+Number of pairwise matches: 4 $\times$ 2
+Swap black and white pieces after each match
+Maximum time per move: 10 s
+Competitive Texas Hold’em
+Max players: 12
+Rounds: 100
+Randomly shuffle seats after each round
+Initial chips: 2000
+Blind increase every 24 hands
+Max hands per round: 720 or until winner decided
+Maximum time per move: 3 s
+Competitive Bridge
+Number of pairwise matches: 12 $\times$ 2
+Swap directions of open/closed rooms
+Use same deck for each pair of matches
+Maximum time per move: 10 s
+Competitive Chess
+Number of pairwise matches: 8 $\times$ 2
+Swap black and white pieces after each match
+Maximum moves per game: 200
+Maximum time per move: 10 s
+Objective OIBench
+Problems per round: 5
+Language: C++
+Per-test timeout: 2s
+Score: mean TimeAUC
+Agent time budget: 60 min
+Objective SWE-Perf
+Instance: repo@base_commit (5 instances per sweep)
+Efficiency tests from dataset
+[…]
+Score: mean significant speedup over baseline
+Figure 5: Cumulative win rate trajectories of Hold’em agents over 300 tournament rounds.
+We list the basic settings for each task in Table 6.
+For competitive tasks, we ensure that the number of pairwise matches among code agents allows the final results to stabilize, i.e., for each task, the L1-norm fluctuation of the scoring matrix $W$ is less than 5%.
+For objective tasks, agents are evaluated independently on the same fixed benchmark each round (no pairwise matches).
+Each agent submits code, the arena runs the official evaluator to compute $W_{i}^{n}$, and a round leaderboard is produced.
+The leaderboard, logs, and top solutions/patches are fed back to all agents for the next round, forming an iterative optimization tournament over $N$ rounds with a diagonal score matrix.
+Specifically, due to the inherent randomness in card dealing in Texas Hold’em, we conduct an additional analysis on the stability of the win rates.
+Figure 5 illustrates the cumulative win rate trends over 10 to 300 rounds for a representative tournament.
+To further assess the reliability of our evaluation, we analyze the win rate trends from different starting rounds to the final round.
+Based on this analysis, starting from round 50, the cumulative win rates of all models become stable, with the absolute value of the linear regression slope dropping below 0.0002 (i.e., per-round changes less than 0.02%).
+This indicates that our tournament design of 100 rounds, with each round consisting of 200 to 720 hands, is sufficient to cover the stochasticity of the game and provides a robust and reliable assessment of model performance.
+The results of the tournament reliably reflect the true capabilities of the evaluated models, ensuring the soundness of subsequent algorithm comparisons and optimizations.
+Appendix C Comparative Analysis of Elo Rating and Global Win Rate
+(a) Elo Rating History
+(b) Win Rate History
+Figure 6: Comparison of performance metrics in a 100-game Texas Hold’em tournament among 6 random selected players.
+(a) The Elo rating fluctuates significantly due to match order sensitivity.
+(b) The global win rate converges to a stable value, providing a more reliable foundation for calculating the evolutionary growth rate.
+In this section, we justify our choice of the global average win rate (the basis for our $G$ score) over the Elo rating system by comparing their stability in a 100-game Texas Hold’em tournament.
+As visualized in Figure 6, the Elo rating (a) exhibits significant volatility and sensitivity to match order, characterized by "zigzag" oscillations that reflect transient outcomes rather than intrinsic capability.
+This "recency bias" makes Elo unsuitable for capturing a stable performance snapshot for each evolutionary round.
+In contrast, the Global Win Rate (b) demonstrates smooth convergence and effectively filters out stochastic noise.
+By reflecting the agent’s expected performance against the entire population, the win rate provides a robust and consistent metric, ensuring that the calculated evolutionary slope represents genuine strategic growth rather than artifacts of tournament scheduling.
+Appendix D Generation Configs
+Table 7: Agent Specifications and Open-Source Status.
+Agent Type Agent Framework Model Agent OSS LLM OSS
+Minimal basic code tools with ADK framework DeepSeek-3.1 (DeepSeek-AI, 2024)
+Qwen3-Coder-480B (Qwen, 2025b)
+Doubao-Seed-1.6 (Seed, 2025)
+GPT-5 (OpenAI, 2025b)
+Claude-4-Sonnet (Anthropic, 2025b)
+Gemini-2.5-Pro (Gemini, 2025)
+Commercial Gemini-CLI (Google, 2025) Gemini-2.5-Pro (Gemini, 2025)
+Claude-Code (Anthropic, 2025a) Claude-4/3.7 Hybrid (Anthropic, 2025b)
+Codex (OpenAI, 2025a) GPT-5 (OpenAI, 2025b)
+Qwen-Code (Qwen, 2025a) Qwen3-Coder-480B (Qwen, 2025b)
+For all LLMs used in our work, we set the temperature to 0.1 and the max tokens to the official API defaults. We set top-p to 1.0, top-k to 100, and the presence penalty to the API default.
 Additionally, both Claude-4-Sonnet and DeepSeek-3.1 occasionally encounter tool call issues that result in no code being generated, as frequently reported by the community.
 If such errors occur three times in a row, we substitute Claude-4-Sonnet with Claude-3.7-Sonnet and DeepSeek-3.1 with DeepSeek v3.
-For LLM-players, considering the uncertainty in model output formats, we allow up to three retries.
-The prompt of LLM-players are in arena’s code and not present in paper considering its excessive length.
-Appendix D Repetition Experiments
-Table 8: Standard deviation of ranking in Round 1 and Round 2 with repeating 4 times.
-Games Average Gomoku Hold’em Bridge Chess
-Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
-Round 1
-Minimal
-Claude-4-Sonnet 0.80 0.91 1.58 0.43 1.12 0.71 0.50 1.64 0.00 0.87
-Deepseek-Chat 0.72 0.81 0.83 0.87 0.83 1.22 1.22 0.71 0.00 0.43
-Doubao-Seed 1.58 0.90 1.87 1.87 1.09 0.43 2.06 0.87 1.30 0.43
-Gemini-2.5-Pro 1.24 1.23 1.50 1.30 1.12 0.83 1.12 1.66 1.22 1.12
-GPT-5 0.75 1.18 0.71 1.50 0.71 1.30 1.09 1.50 0.50 0.43
-Qwen3-Coder 1.16 0.84 1.48 1.50 1.22 0.71 1.09 0.71 0.83 0.43
-Commercial
-Claude-Code 1.27 1.01 1.12 1.66 0.43 0.00 1.79 1.09 1.73 1.30
-CodeX 0.76 0.57 0.83 0.50 0.43 0.50 1.09 1.30 0.71 0.00
-Gemini-CLI 1.14 0.93 0.83 0.43 1.12 1.00 1.30 1.79 1.30 0.50
-Qwen-Coder 1.01 0.95 1.22 1.48 1.12 0.00 0.87 0.83 0.83 1.50
-Round 2
-Minimal
-Claude-4-Sonnet 0.81 0.55 1.48 0.50 0.83 0.43 0.50 0.83 0.43 0.43
-Deepseek-Chat 0.94 1.03 1.30 1.09 0.87 0.83 1.09 1.48 0.50 0.71
-Doubao-Seed 0.79 0.88 1.09 1.12 0.87 0.71 0.71 0.87 0.50 0.83
-Gemini-2.5-Pro 1.30 1.28 1.66 1.92 1.00 1.22 2.06 1.12 0.50 0.87
-GPT-5 0.89 1.22 1.00 1.58 0.43 1.66 1.64 1.22 0.50 0.43
-Qwen3-Coder 0.82 1.02 1.58 1.30 0.83 0.43 0.43 1.50 0.43 0.83
-Commercial
-Claude-Code 1.20 1.04 1.09 1.41 1.22 0.43 1.64 1.09 0.83 1.22
-CodeX 0.71 0.87 0.43 0.87 0.50 1.09 0.83 1.09 1.09 0.43
-Gemini-CLI 0.88 1.10 0.83 1.64 0.83 0.83 1.41 1.48 0.43 0.43
-Qwen-Coder 1.13 0.90 1.66 0.83 0.43 0.83 1.00 1.12 1.41 0.83
-We report the results of repetition experiments($N=4$) on first two tournament round in Table 8.
-From table, we observe that 1.
-The rankings of most agents remain relatively stable ,with ranking standard deviation changes of less than one.
-However, a few agents, such as Gemini-2.5-Pro and Claude-Code, exhibit greater fluctuations; 2.
-The rankings for standard games are more stable than those for variant games; 3.
-The results of the open source model are more stable than those of the closed source model, and commercial agents are more stable than minimal agents; 4.
-Additionally, we observe that agents do not consistently generate runnable code repositories across multiple development attempts.
-Even commercial agents occasionally fail to produce successful builds, which suggests that current code agents still need to improve their development stability.
-Appendix E Learning Ability
-E.1 Global Learning Trend
+Appendix E Repetition Experiments: Stability Analysis
+Table 8: Standard deviation of agent ranking in Round 1 and Round 2 with repeating 4 times.
+Games Gomoku Hold’em Bridge Chess OIBench SWE-Perf
+Round 1 Minimal Claude-4-Sonnet 1.58 1.12 0.50 0.00 0.50 0.82
+Deepseek-Chat 0.83 0.83 1.22 0.00 0.50 1.83
+Doubao-Seed 1.87 1.09 2.06 1.30 0.96 2.00
+Gemini-2.5-Pro 1.50 1.12 1.12 1.22 0.58 0.96
+GPT-5 0.71 0.71 1.09 0.50 0.00 0.50
+Qwen3-Coder 1.48 1.22 1.09 0.83 0.50 0.82
+Comm. Claude-Code 1.12 0.43 1.79 1.73 0.96 -
+Codex 0.83 0.43 1.09 0.71 1.41 -
+Gemini-CLI 0.83 1.12 1.30 1.30 0.82 -
+Qwen-Coder 1.22 1.12 0.87 0.83 0.96 -
+Round 2 Minimal Claude-4-Sonnet 1.48 0.83 0.50 0.43 0.50 1.23
+Deepseek-Chat 1.30 0.87 1.09 0.50 0.58 0.00
+Doubao-Seed 1.09 0.87 0.71 0.50 2.16 0.00
+Gemini-2.5-Pro 1.66 1.00 2.06 0.50 0.82 1.50
+GPT-5 1.00 0.43 1.64 0.50 1.00 1.16
+Qwen3-Coder 1.58 0.83 0.43 0.43 0.50 0.00
+Comm. Claude-Code 1.09 1.22 1.64 0.83 0.58 -
+Codex 0.43 0.50 0.83 1.09 1.91 -
+Gemini-CLI 0.83 0.83 1.41 0.43 1.73 -
+Qwen-Coder 1.66 0.43 1.00 1.41 0.58 -
+Table 8 presents the results of four independent repetitions of the first two tournament rounds.
+The primary objective is to assess whether stochasticity in code generation significantly alters the competitive landscape.
+The data reveals that ranking stability is robust: the majority of agents maintain their relative positions with a standard deviation of less than one rank.
+While specific implementations vary per run, the performance gap between agents is sufficiently large to maintain a stable leaderboard.
+Other minor observations include: (1) standard games exhibit higher stability than variant games; (2) open-source and commercial agents generally outperform minimal agents in terms of consistency.
+Finally, we observe that "runnability" is not guaranteed across runs, suggesting that while the relative capability (ranking) is stable, the absolute success rate of code generation still suffers from inherent randomness.
+Appendix F Error Rate Analysis
+Table 9: Proportion of HTTP errors encountered during code execution for each agent in Round 1.
+Gomoku Hold’em Bridge Chess
+HTTP error(%) Standard Variant Standard Variant Standard Variant Standard Variant
+Minimal Claude-4-Sonnet 7 0 0 0 0 0 0 0
+Deepseek-Chat 21 0 0 48 0 0 100 100
+Doubao-Seed 17 0 49 93 12 0 0 100
+Gemini-2.5-Pro 3 100 8 50 0 1 0 0
+GPT-5 6 0 0 0 0 100 2 21
+Qwen3-Coder 11 0 0 91 2 1 0 0
+Commercial best ADK 0 0 0 0 0 0 0 0
+Claude-Code 0 0 0 0 0 0 0 0
+Codex 0 0 0 0 0 0 0 0
+Gemini-CLI 0 0 0 0 48 0 0 0
+Qwen-Coder 2 0 0 18 0 17 0 0
+During the implementation of strategy code, some agents encounter runtime errors.
+We observe that all agents were able to generate the required files successfully.
+However, most errors occur in interface implementation and action validity.
+These cases were consistently recorded as HTTP errors.
+Whenever an HTTP error occurs, a default action is automatically executed to ensure the continuity of the competition.
+For Texas Hold’em Poker, the default action is to fold for that hand.
+For Chess and Gomoku, the default action is for the player to concede the current game.
+Table 9 reports the proportion of HTTP error actions encountered during the code development of each agent in Round 1.
+The error rate is calculated as the rate of failed actions in total number of actions taken by the strategy code.
+The experimental results indicate that commercial agents exhibit a significantly lower failure rate compared to minimal agents.
+Additionally, the failure rate is higher in variant games than in traditional games.
+Appendix G Detailed Evolutionary Trajectory Analysis
+Figure 7: Trend of global performance score $G_{i}^{n}$ in all tasks.
+In this section, we provide a comprehensive visualization of the evolutionary trajectories across all six tasks.
+Figure 2 plots the round-wise global performance $G^{n}$ for each agent, with dashed lines representing the fitted regression of $S_{\text{evo}}$ to reflect the evolution trend.
+  1. 1.
+    Universality of Performance Oscillation.
+The visualization confirms that performance volatility is not task-specific but a universal phenomenon in agent-based evolution.
+In environments with high strategic complexity, such as Gomoku and Bridge, many agents (e.g., ADK Claude, ADK Gemini) experience significant performance drops in intermediate rounds before potentially recovering.
+This widespread oscillation reinforces our conclusion in the main text: current agents often fail to reliably distinguish between beneficial and detrimental strategy modifications, leading to the accidental discarding of high-performing code versions.
+  2. 2.
+    Latent Potential Discovery via Iteration.
+The trajectories across different tasks suggest that static, zero-shot evaluations often fail to capture an agent’s upper bound.
+Even in the highly challenging SWE-Perf task, despite the low absolute scores, certain agents exhibit clear upward trajectories, suggesting that iterative feedback allows agents to gradually navigate the complex search space of software engineering.
+  3. 3.
+    Reliability of the $S_{\text{evo}}$ Metric.
+The figure highlights the importance of our proposed $S_{\text{evo}}$.
+While the raw $G^{n}$ is highly sensitive and sometimes suffers from development failures in specific rounds, the dashed regression lines remain relatively stable.
+This indicates that our evolutionary score provides a more accurate measure of the agent’s evolutionary progress than any single-round performance snapshot.
+Appendix H Pairwise Code Similarity Heatmaps
+In this section, we provide the detailed pairwise similarity matrices between Round 1 and Round 2 code submissions for all six tasks (Figure 8).
+In Competitive Tasks, similarity is measured by Action Consistency (%).
+A value of 100.0 indicates that the agent’s code produces identical behavioral outputs in endgame positions.
+In objective tasks, similarity is measured by CodeBLEU.
+These matrices generally show lower absolute values and higher dispersion, reflecting the complexity of syntactic and semantic alignment in large-scale code repositories.
+A key observation is that the distribution and range of similarity scores vary significantly by task, reflecting the inherent difficulty of learning and replicating code in different domains.
+In Hold’em (b), the matrix is characterized by exceptionally high similarity values, including several instances of 100.0% alignment.
+This indicates that the logic required for this task is relatively "easy to learn" and replicate.
+Agents can effortlessly adopt a peer’s strategy or maintain their own with perfect consistency, suggesting a lower barrier to entry for strategic imitation in this environment.
+In Gomoku (a), Bridge (c), and Chess (d), the similarity scores are more dispersed.
+In objective tasks, even the highest similarity scores rarely exceed 60-70%.
+This is due to two factors: (1) These tasks involve larger codebases and more intricate dependencies, making it difficult for agents to maintain high syntactic or semantic overlap during refinement.
+(2) The use of CodeBLEU means that even minor functional improvements or refactoring can lead to a significant drop in measured similarity.
 (a) Gomoku.
-(b) Hold’em
+(b) Hold’em.
 (c) Bridge.
 (d) Chess.
-Figure 4: Trend of global performance score $G_{i}^{n}$ in Gomoku, Hold’em, Bridge and Chess.
-As shown in Figure 4, we present the trends of global performance scores $G_{i}^{n}$ across four games, revealing distinct performance patterns for different models.
-In many cases, agents experience a sharp decline in performance during an intermediate round, which we interpret as a learning failure.
-Typically, such failures are recovered in the following round.
-E.2 Detailed Learning Score
-We list the detailed score of global learning, counter-adpatation learning and self-improvement in table 91011 respectively.
-Table 9: Global Learning with Group-wise Average Rankings.
-Models Avg. Ranking Gomoku Hold’em Bridge Chess
-Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
-Minimal
-Claude-4-Sonnet 2.50 2.75 -0.447 -0.156 0.118 0.110 0.174 0.047 -0.170 0.018
-Deepseek-Chat 2.75 2.75 0.027 0.077 0.010 -0.022 0.000 0.000 0.000 0.049
-Doubao-Seed 4.75 4.50 -0.192 -0.302 -0.035 0.000 -0.033 -0.516 -0.337 0.034
-Gemini-2.5-Pro 3.75 2.75 -0.066 0.173 0.020 0.078 -0.195 -0.049 -0.147 0.003
-GPT-5 3.50 3.75 0.062 -0.019 0.102 -0.050 -0.095 0.293 -0.525 -0.000
-Qwen3-Coder 3.75 4.50 -0.523 -0.089 0.038 0.003 0.032 -0.230 -0.187 -0.532
-Commercial
-best ADK 2.25 3.75 0.075 -0.022 0.073 0.030 0.295 0.361 -0.110 -0.342
-Claude-Code 3.75 2.75 -0.322 0.194 0.100 0.105 -0.240 -0.139 -0.158 -0.226
-CodeX 2.75 3.00 0.454 -0.095 0.050 0.067 0.098 0.285 0.033 0.064
-Gemini-CLI 2.25 4.00 0.260 0.172 0.050 -0.007 0.254 -0.286 0.395 -0.154
-Qwen-Coder 3.75 1.25 -0.054 0.536 0.058 0.105 0.157 0.556 -0.204 0.039
-Table 10: Counter-adaptation Score with Group-wise Average Rankings.
-Games Avg. Ranking Gomoku Hold’em Bridge Chess
-Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
-Minimal
-Claude-4-Sonnet 3.75 3.75 -0.096 -0.075 0.001 -0.023 0.005 0.128 -0.042 -0.075
-Deepseek-Chat 2.75 2.00 0.008 0.088 0.004 0.021 0.000 0.000 0.000 0.100
-Doubao-Seed 4.25 4.50 0.063 -0.196 -0.023 -0.061 -0.008 -0.133 -0.196 0.083
-Gemini-2.5-Pro 2.75 2.75 0.354 0.192 0.014 0.097 -0.132 -0.238 -0.038 0.021
-GPT-5 4.75 3.25 0.038 -0.012 -0.086 0.019 -0.098 0.052 -0.154 -0.025
-Qwen3-Coder 2.75 4.75 -0.092 0.029 0.000 -0.080 0.037 -0.008 0.025 -0.167
-Commercial
-best ADK 2.50 3.00 0.260 0.000 0.018 0.039 0.127 0.238 -0.094 -0.104
-Claude-Code 3.25 1.88 -0.042 0.083 0.091 0.194 0.060 0.081 -0.104 0.047
-CodeX 3.50 3.50 0.104 -0.089 -0.034 0.033 0.023 -0.090 0.010 0.052
-Gemini-CLI 2.75 4.25 0.120 -0.021 0.031 -0.083 -0.027 -0.169 0.188 0.010
-Qwen-Coder 3.00 2.38 -0.130 0.193 0.041 0.077 0.092 0.081 -0.062 -0.021
-Table 11: Self-improvement Score with Group-wise Average Rankings.
-Games Avg. Ranking Gomoku Hold’em Bridge Chess
-Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
-Minimal
-Claude-4-Sonnet 2.75 2.50 -0.103 -0.894 0.949 0.517 0.858 0.766 -0.848 0.478
-Deepseek-Chat 3.25 3.38 0.000 0.893 -0.949 -0.747 0.000 0.000 0.000 0.000
-Doubao-Seed 3.75 4.38 0.141 -0.686 0.075 0.000 -0.202 -0.775 -0.894 0.000
-Gemini-2.5-Pro 3.63 2.00 0.400 0.775 -0.758 0.894 -0.598 -0.240 -0.775 0.913
-GPT-5 4.88 4.50 -0.897 -0.949 -0.205 -0.050 -0.767 0.485 -0.775 -0.390
-Qwen3-Coder 2.75 4.25 -0.400 0.161 0.668 0.202 0.473 -0.400 -0.258 -0.775
-Commercial
-best ADK 2.25 3.50 0.400 0.956 0.835 -0.614 0.738 0.546 -0.207 -0.726
-Claude-Code 4.00 2.25 -0.230 0.969 0.346 0.904 0.113 0.537 -0.730 -0.225
-CodeX 2.00 3.25 0.763 -0.763 -0.090 0.602 0.784 0.316 0.424 0.316
-Gemini-CLI 3.75 4.00 -0.183 -0.356 -0.176 0.000 0.000 -0.995 0.811 -0.193
-Qwen-Coder 3.00 2.00 0.632 0.717 0.826 0.705 0.641 0.677 -0.944 0.000
-The trend on four games are rather different.
-In general, the commercial model group consistently demonstrates superior global learning capability, where the advantage is particularly evident in complex strategy games like Chess or Gomoku variant.
-Despite most minimal agents fail to learn well on complex games, we still find that Claude-4-Sonnet significantly surpass the rivals on standard games.
-However, the Claude-4-Sonnet still lacks behind on some cases like Gomoku, indicating that current LLMs agentic ablilities are still limited by the framework, where the commericial agents optimize workflow for their specific models to achieve the best results.
-In simple games such as Hold’em, a larger proportion of agents exhibit positive learning scores, whereas in complex games like Chess, the prevalence of negative scores increases markedly.
-This trend suggests that current agents still face significant limitations in learning complex strategies.
-E.3 Behavioral Changes Induced by Learning
-For each game, we randomly select 80-100 intermediate states from the agents’ rival history and require the agent or LLM-player to choose the next move for each state.
-To ensure clarity in our writing, we uniformly refer to these intermediate states as endgame throughout the paper.
-Please note that endgame here is not limited to the final stages of the game; samples are taken from early, middle, and late stages as well.
-We visualize the action consistency among the first two rounds in four games’s endgame in Figure 5.
-From the matrix, we observe that, in general, agents tend to learn the strategies of other agents in the first round (lower left part vs.
-upper left part).
-Specifically, Doubao-Seed and DeepSeek-Chat simply copy Claude-4-Sonnet’s strategy in Holdem.
-Additionally, the learning trend is more pronounced in simpler games (Holdem vs.
-Chess).
-For simpler games like Holdem, agent strategies in the second round are more similar to those in the first round (lower right part vs.
-upper left part), while for more difficult games like Chess, the trend is reversed.
-This observation is consistent with our findings on $\text{Trend}_{\text{mean}}$ in Table 5.
-(a) Holdem.
-(b) Gomoku.
-(c) Bridge.
-(d) Chess.
-Figure 5: Action consistency between round 1 and round 2 agents’ code on endgames.
-Appendix F Comparison between Agent and LLM-Player
-Table 12: Comparison of match outcomes between each agent and its corresponding LLM-Player. Each value indicates the agent’s win rate when competing against the LLM that powers it.
-Agent VS LLM Gomoku Hold’em Bridge Chess
-Standard Variant Standard Variant Standard Variant Standard Variant
-Claude-4-Sonnet 1.00 1.00 0.00 0.00 0.45 1.00 0.88 0.75
-Deepseek-Chat 0.50 0.00 0.40 0.30 0.00 0.00 0.00 0.00
-Doubao-Seed 0.50 0.25 0.00 0.00 0.00 0.10 0.00 0.00
-Gemini-2.5-Pro 0.00 0.25 0.00 0.00 1.00 1.00 0.13 0.50
-GPT-5 0.00 0.50 0.00 0.50 1.00 0.00 0.00 0.00
-Qwen3-Coder 1.00 0.00 0.50 0.50 1.00 0.00 0.50 0.50
-Claude-Code 1.00 1.00 0.00 0.00 0.70 0.85 0.75 0.63
-CodeX 0.00 0.50 0.20 0.40 1.00 0.38 1.00 1.00
-Gemini-CLI 0.00 0.00 0.10 0.40 0.00 1.00 0.38 0.50
-Qwen-Coder 1.00 1.00 0.50 0.00 1.00 0.00 0.56 0.63
-We compare the strategies of agents’ code and its corresponding LLM-Player on four games in table 12.
-Interestingly, there is no strong correlation between the performance of the agents’ code and that of their underlying models.
-In games with strong strategic elements, such as Gomoku and Chess, some agents’ code significantly outperforms their corresponding LLM-Player, indicating that the code implementation is able to better leverage game rules and strategies.
-For example, the agent developed by claude-4-sonnet achieves a 100% win rate against its LLM-Player in both standard and variant Gomoku, and also demonstrates a high win rate in Chess and Bridge.
-This suggests that the strategies implemented in the code are superior to the large model’s direct reasoning performance as a player in these games.
-In contrast, the agents developed by doubao-seed and deepseek-chat struggle to defeat their respective models.
-However, in Hold’em, agents generally have lower win rates than the LLM-Player, possibly because the LLM-Player performs better in games with more psychological tactics, which are difficult to simulate with code while can be summarized by context learning.
-We further visualize the action consistency between agents’ code and LLM-palyers in Figure 6.
-We find that, in most games, the actions of LLM-players tend to resemble each other, and the strategies implemented in agents’ code also exhibit high similarity among themselves.
-However, the code-based strategies and the plain reasoning of the same model often differ substantially.
-The only exception is Bridge, where we find numerous cases in which both LLM-players and code agents exhibit low consistency with human decisions.
-Considering that Bridge allows for a certain degree of decision freedom and its bidding rules are not strictly unified, we attribute this phenomenon to the intrinsic characteristics of the game.
-Similar observations are also reported in other studies (Kita et al., 2024).
-These findings further demonstrate that the strategies generated by the agents and those employed by the LLM-players are based on different approaches.
-This difference merits additional investigation in subsequent studies.
-Appendix G Case Study on Code
-G.1 Strategy of Agents
-In Gomoku, strategies display clear stratification.
-Gemini and DeepSeek rely on random or near-random moves, while Claude, Doubao, GPT-5, and Qwen3 employ similar pattern-based evaluation with candidate filtering and Minimax search.
-Differences mainly lie in threat recognition, search control, and opening play: Claude and Doubao handle openings and forced moves more effectively, GPT-5 is steadier under time limits, Qwen3 remains balanced, whereas Gemini and DeepSeek are notably weaker.
-The code similarity among agents in Texas Hold’em is relatively high, only DeepSeek employs a fully random strategy, while other models calculate winning probabilities based on the hand.
-On one hand, this is because the available actions in Texas Hold’em are limited to fold, call/raise, and check.
-On the other hand, the strategies for Texas Hold’em are relatively straightforward to implement, as both reasoning and code are primarily based on hand strength.
-As a result, the code can closely simulate the reasoning process.
-In the case of Chess, DeepSeek relies on an external library (Stockfish), but fails to configure it correctly, resulting in unsuccessful development.
-Even after multiple development iterations, DeepSeek continues to use this library without resolving the configuration issues.
-We also find that Claude, Doubao, Gemini, GPT, and Qwen3 utilize a similar combination of heuristic piece and board evaluation, Minimax search, and alpha-beta pruning, which leads to similar behavior.
-There are slight differences in how each model evaluates the value of Chess pieces and the actions in endgame scenarios.
-Notably, Claude incorporates an opening book, which distinguishes it from the others and leads to better performance.
-For Bridge, bidding and play strategies also stratify clearly.
-Qwen3 and Gemini rely on minimal logic, following random choices or only basic rules on High Card Points (HCP).
-By contrast, GPT-5, Doubao, and Claude incorporate structured evaluation, moving from total point counting (GPT-5) to multi-layered systems with suit quality, competitive actions, and signaling (Doubao and Claude).
-Despite these differences, all models share reliance on HCP as a core metric.
-Overall, Claude achieves the most complete integration of evaluation and play, Doubao is comparably advanced, GPT-5 remains simpler, while Gemini and Qwen3 lag behind.
-(a) Holdem.
-(b) Gomoku.
-(c) Bridge.
-(d) Chess.
-Figure 6: Action consistency between agents’ code and LLM-Players on endgames.
-Appendix H Results of ML Track
+(e) OIBench.
+(f) SWE-Perf.
+Figure 8: Pairwise similarity matrices between Round 1 and 2 code submissions across six tasks.
+For competitive games (a-d), similarity is measured by Action Consistency in endgame positions.
+For programming tasks (e-f), similarity is measured by CodeBLEU.
+Appendix I Case Study of Agent Learning
+To demonstrate how agents actively improve their strategies by analyzing other agents’ code and tournament reports, we present a detailed case study focusing on Claude-Code_A in Texas Hold’em Poker tournaments.
+This case clearly shows that score changes across rounds reflect the agent’s understanding and adaptation to opponent strategies.
+Development Planning and Evidence of Learning
+Analysis of the Round 3 development log shows that the agent explicitly planned to study other agents’ strategies. The planning log has the following entries:
+1. "Read game project documentation and requirements" - completed
+2. "Analyze tournament rules and mode" - ongoing
+3. "Study last round tournament reports and AI strategies" - pending
+4. "Examine existing AI code from round 2" - pending
+Tool Use Summary
+Table 10 summarizes all the learning-related tool use operations performed by claude-code_A during the development of its Round 3 code.
+The tool use logs show that the agent first located and read the tournament report file, extracted win statistics using Grep (codex_A: 35 wins, gemini-cli_B: 30 wins, claude-code_A: 10 wins), then listed all agents, and finally read the code of the top performers.
+Table 10: Summary of learning-related tool use operations by claude-code_A (Round 3 development).
+Tool Operation Purpose
+Glob Search **/tournament_report_tourney_*.json Locate tournament report files
+Glob Search **/history_tourney_*.json Locate tournament history files
+Read Read tournament_report_tourney_075455.json Analyze agent performance
+Read Read first 100 lines of report Quickly understand report structure
+Grep Search for "wins": field Extract win counts for each agent
+Glob List round_2/**/* Discover all agents in previous round
+Read Read own previous solutions Review existing implementation
+Read Read codex_A code Peer-strategy extraction (Top performer, 35 wins)
+Read Read gemini-cli_B code Cross-agent logic synthesis (second performer, 30 wins)
+Strategy Adaptation in Code
+Based on the discovery that top-performing agents (codex_A) utilized aggressive betting patterns, the agent transitioned from a static probability-based model to a dynamic opponent-aware strategy.
+Comparing the code between rounds, it is evident that the agent made targeted improvements based on its analysis:
+  • Round 2: Basic Chen formula evaluation, simple tournament strategy, no opponent modeling.
+  • Round 3: Added opponent analysis (analyze_opponents()), improved decision formula incorporating opponent aggression, refined stack and position factors.
+Example of newly added code in Round 3:
+1def analyze_opponents(self, game_state: Dict[str, Any]) $\rightarrow$ Dict[str, float]:
+2 """Analyze opponent tendencies based on action history"""
+3 action_history = game_state.get('action_history', [])
+4 # ... analyze aggression for each opponent ...
+5 return opponent_stats
+6
+7# Use opponent analysis in decision making
+8opponent_stats = self.analyze_opponents(game_state)
+9avg_opponent_aggression = sum(opponent_stats.values()) / max(len(opponent_stats), 1)
+10decision_strength = (
+11 hand_strength *
+12 (1 + position_factor * 0.4) *
+13 stack_factor *
+14 bubble_factor *
+15 (1 + (1 - avg_opponent_aggression) * 0.2)
+16)
+These code modifications are targeted responses to the strategies employed by top-performing agents, enabling more effective adaptation and strategic improvement.
+Reflections in Development Summary
+After completing development, the agent’s summary explicitly compares its performance and highlights improvements:
+Agent Self-Reflection Log Performance Analysis from Round 2 Based on the tournament results: - codex_A: 35 wins (best performer) - gemini-cli_B: 30 wins - claude-code_A: 10 wins (my previous version) My new AI improves on the previous version with: - Better preflop ranges and position play - Enhanced tournament strategy - More sophisticated opponent modeling - Improved risk management
+Similar learning behavior is observed for codex_A and other agents.
+This case study clearly demonstrates that agents actively analyze tournament reports and top-performing code via explicit tool use operations.
+The sequence and content of these operations, combined with targeted code improvements and reflective summaries, confirm that score changes across rounds are a reliable indicator of the agent’s growing understanding and adaptation to competitor strategies.
+This micro-level behavioral loop consists of performance benchmarking, code synthesis, and self-reflection.
+It provides a mechanistic explanation for the macro-level performance leaps observed in our evolutionary trajectories, as agents successfully internalize superior logic from their competitive environment.
+Appendix J Robustness Analysis on Variant Tasks
+This section details the variant settings and analyzes the impact of rule changes on agent performance.
+The primary objective is to demonstrate the robustness of the CATArena framework: even when agents cannot rely on memorized opening books or standard algorithms, our metric for evolutionary capability $S_{\text{evo}}$ remains a reliable indicator of their learning capability.
+Variant Configurations
+We introduce specific rule perturbations designed to disrupt established priors while preserving the core reasoning logic:
+  • Gomoku: Introduce forbidden points (e.g., double-three restrictions) to constrain the search space.
+  • Texas Hold’em: Alter probability landscapes by removing specific cards and inverting hand rank hierarchies.
+  • Chess: Adopt Chess960 (Fischer Random Chess) and modified movement rules to neutralize opening databases.
+  • Bridge: Implement a "Card Exchange" mechanism between teammates to disrupt standard signaling conventions.
+  • OIBench: Shift the optimization objective from time complexity to space complexity, requiring memory-efficient refactoring rather than speed optimization.
+Table 11: Results of variant tasks.
+We report the Base Performance ($S_{\text{base}}$) and Evolutionary Capability ($S_{\text{evo}}$) of Minimal and Commercial Agents across five variants of standard tasks.
+Best results in each column are bolded.
+Agents Competitive Objective
+Gomoku Hold’em Bridge Chess OIBench
+$S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$ $S_{\text{base}}$ $S_{\text{evo}}$
+Minimal Claude 0.78 -0.100 0.13 +0.017 1.00 +0.031 0.65 -0.014 0.23 +0.111
+DeepSeek 0.38 +0.063 0.00 -0.008 0.10 0.000 0.10 +0.020 0.07 -0.021
+Doubao 0.73 -0.201 0.00 0.000 0.45 -0.155 0.10 +0.018 0.22 +0.181
+Gemini 0.00 +0.159 0.00 +0.043 0.60 -0.011 0.90 -0.016 0.56 -0.028
+GPT 0.76 -0.008 0.87 +0.014 0.10 +0.014 0.45 +0.004 0.66 +0.003
+Qwen 0.36 +0.041 0.00 -0.008 0.76 -0.021 0.80 -0.158 0.41 -0.011
+Commercial Best ADK Agent 0.75 +0.012 0.46 +0.010 0.00 +0.109 1.00 -0.182 0.75 +0.036
+Claude Code 0.66 +0.093 0.00 +0.038 0.93 +0.024 0.44 -0.058 0.74 +0.019
+Codex 0.69 -0.087 0.17 +0.030 0.50 +0.013 0.34 +0.029 0.00 0.000
+Gemini CLI 0.19 -0.058 0.37 0.000 0.83 -0.191 0.38 -0.012 0.66 +0.160
+Qwen Code 0.22 +0.146 0.00 +0.049 0.25 +0.144 0.34 +0.011 0.70 -0.042
+Results and Discussion
+Table 11 compares the performance metrics between standard and variant settings. We observe two critical phenomena:
+  1. 1.
+    Widening Gap in Baseline Performance $S_{\text{base}}$: Unlike standard tasks where scores are often tightly clustered, variant tasks exhibit higher variance in zero-shot rankings.
+This is because rule perturbations cause less capable agents to fail in generating valid or robust strategies, thereby amplifying the performance distinction between agents that rely on memorization versus those with genuine reasoning capabilities.
+  2. 2.
+    Consistency in Evolutionary Capability: Crucially, the evolutionary metric retains its effectiveness in distinguishing agent capabilities.
+Despite the altered baselines, the relative evolutionary trends align with those in standard tasks (where most capable agents achieve $S_{\text{evo}}>0$).
+This demonstrates that CATArena effectively captures the gradient of learning ability: it can reliably differentiate between agents that can adapt to new rules and those that cannot, proving that the evaluation of evolutionary potential is robust to task variations.
+Appendix K Results of Multilingual Track
+Table 12: Scores of agents on games across programming languages, with variance analysis.
+[…]
+[…]
+Minimal Claude-4-Sonnet 1.000 0.250 0.250 0.125 0.360 0.640 0.000 0.069 1.000 0.250 0.250 0.125 0.106
+DeepSeek-Chat 1.000 0.250 0.250 0.125 1.000 0.000 0.000 0.222 0.500 0.500 0.500 0.000 0.116
+Doubao-Seed 1.000 0.500 0.000 0.167 1.000 0.000 0.000 0.222 0.500 1.000 0.000 0.167 0.185
+Gemini-2.5-Pro 0.750 0.750 0.000 0.125 0.010 0.990 0.000 0.216 1.000 0.250 0.250 0.125 0.155
+GPT-5 0.500 0.000 1.000 0.167 1.000 0.000 0.000 0.222 1.000 0.500 0.000 0.167 0.185
+Qwen3-Coder 1.000 0.250 0.250 0.125 0.610 0.290 0.100 0.044 0.688 0.000 0.812 0.128 0.099
+Commercial Claude-Code 1.000 0.250 0.250 0.125 0.200 0.020 0.780 0.105 1.000 0.250 0.250 0.125 0.118
+Codex 0.000 0.500 1.000 0.167 0.200 0.030 0.770 0.100 1.000 0.000 0.500 0.167 0.145
+Gemini-CLI 0.750 0.750 0.000 0.125 1.000 0.000 0.000 0.222 1.000 0.250 0.250 0.125 0.157
+Qwen-Coder 1.000 0.250 0.250 0.125 1.000 0.000 0.000 0.222 0.975 0.000 0.525 0.159 0.169
+Table 12 details the performance of agents across five programming languages: Python, C++, Java, JavaScript, and Go.
+The primary goal of this track is to assess whether agents can abstract strategic logic independently of implementation syntax.
+  1. 1.
+    Domain-Specific Language Bias: We observe a strong correlation between task type and language performance.
+In competitive tasks, agents achieve peak performance in Python, consistent with the prevalence of Python in AI and game scripting.
+Conversely, in OIBench, agents implementing algorithms in C++ significantly outperform those using Python.
+This suggests that agents rely heavily on the statistical distributions of their pre-training data (where competitive programming is dominated by C++) rather than possessing a generalized coding capability.
+  2. 2.
+    Generalization Gaps: Significant performance degradation is observed in languages that are less aligned with the specific domain (e.g., JavaScript/Go for games).
+Qwen3-Coder demonstrates the most consistent results across all languages with the lowest average variance, indicating strong cross-language adaptability.
+In contrast, models such as GPT-5 and Doubao-Seed show considerable fluctuations, reflecting limited generalization ability.
+Considering that the core logic for both games and algorithms is inherently language-agnostic, the observed performance gaps indicate a critical failure in algorithmic abstraction.
+Current agents struggle to decouple strategic reasoning from syntactic implementation, failing to transfer a known solution (e.g., a C++ algorithm) to a different language context (e.g., Python) without performance loss.
+Appendix L Results of ML Track
 Table 13: ML ability scores and average rankings of agents.
 […]
-Minimal
-Claude-4-Sonnet 0.787 0.360 0.600 0.700 2.25
+Minimal Claude-4-Sonnet 0.787 0.360 0.600 0.700 2.25
 DeepSeek-Chat 0.612 0.000 0.170 0.100 4.25
@@
 Qwen3-Coder 0.600 0.000 0.900 0.725 2.50
-Commercial
-best ADK 0.750 0.190 0.700 0.656 1.25
+Commercial best ADK 0.750 0.190 0.700 0.656 1.25
 Claude-Code 0.578 0.170 0.000 0.406 4.00
@@
 Qwen-Coder 0.500 0.170 0.700 0.531 2.50
-The detailed results of agents’ performance on machine learning devlopments is shown in table 13.
-In ML track, agents autonomously generate data, design code, train models, and deliver ML-based strategies in a GPU-enabled environment.
-The results of ML ability is provided in Appendix H.
-Most agents only implement basic models with limited training, resulting in smaller performance gaps and different rankings compared to the strategy track.
-Appendix I Results of Multi-lingual Track
-We list the detailed results on different languages in  14.
-Table 14: Scores of agents on games with different languages, with variance analysis.
-[…]
-[…]
-Minimal
-Claude-4-Sonnet 1.000 0.250 0.250 0.125 0.360 0.640 0.000 0.069 1.000 0.250 0.250 0.125 0.106
-DeepSeek-Chat 1.000 0.250 0.250 0.125 1.000 0.000 0.000 0.222 0.500 0.500 0.500 0.000 0.116
-Doubao-Seed 1.000 0.500 0.000 0.167 1.000 0.000 0.000 0.222 0.500 1.000 0.000 0.167 0.185
-Gemini-2.5-Pro 0.750 0.750 0.000 0.125 0.010 0.990 0.000 0.216 1.000 0.250 0.250 0.125 0.155
-GPT-5 0.500 0.000 1.000 0.167 1.000 0.000 0.000 0.222 1.000 0.500 0.000 0.167 0.185
-Qwen3-Coder 1.000 0.250 0.250 0.125 0.610 0.290 0.100 0.044 0.688 0.000 0.812 0.128 0.099
-Commercial
-Claude-Code 1.000 0.250 0.250 0.125 0.200 0.020 0.780 0.105 1.000 0.250 0.250 0.125 0.118
-CodeX 0.000 0.500 1.000 0.167 0.200 0.030 0.770 0.100 1.000 0.000 0.500 0.167 0.145
-Gemini-CLI 0.750 0.750 0.000 0.125 1.000 0.000 0.000 0.222 1.000 0.250 0.250 0.125 0.157
-Qwen-Coder 1.000 0.250 0.250 0.125 1.000 0.000 0.000 0.222 0.975 0.000 0.525 0.159 0.169
-Most agents achieve their highest scores in Python, while several models exhibit significant performance fluctuations in JS and Go.
-Qwen3-Coder demonstrates the most consistent results across all languages, with the lowest average variance, indicating strong cross-language adaptability.
-In contrast, models such as GPT-5 and Doubao-Seed show considerable differences between languages, reflecting limited generalization ability.
-Commercial agents also exhibit score differences across different programming languages.
-Considering that board game strategies are inherently language-agnostic tasks, the performance gaps observed in the multi-language track of CATArena indicate that current agents are not yet able to effectively abstract strategies into unified algorithms and implement them consistently across languages.
-Such algorithmic abstraction should be a key direction for the future development of agents.
-Appendix J Cost and Code Complexity of Participants
-Table 15: Comparison of model cost on four games of 1st round.
+In the ML track, agents are tasked with an end-to-end machine learning pipeline: autonomously generating data, designing architectures, training models in a GPU-enabled environment, and deploying the model for inference.
+The results are summarized in Table 13.
+Unlike the Strategy Track where logic is explicit, the ML Track requires agents to manage a complex, multi-stage workflow.
+  1. 1.
+    Limited Technical Depth.
+Despite the availability of GPU resources, we observe that agents generally fail to leverage advanced deep learning techniques.
+Most agents only managed to implement basic models (e.g., simple MLPs or decision trees) with limited training epochs.
+Additionally, a subset of agents fail to implement a valid training loop entirely, submitting models with randomized parameters instead of learned weights.
+  2. 2.
+    Commercial vs.
+ADK Agent.
+A distinct capability gap exists in end-to-end development stability.
+All commercial agents successfully completed the full lifecycle—from data generation to model deployment.
+In contrast, ADK agents struggled with pipeline integration, frequently failing to produce a runnable submission due to environment configuration errors or logical breaks in the workflow.
+  3. 3.
+    Performance vs.
+Rule-Based Track.
+Currently, the performance of agents in the ML track significantly lags behind the rule-based strategies observed in the standard Strategy Track.
+This underperformance is attributed to the agents’ inability to autonomously debug and optimize complex ML workflows.
+Consequently, due to the low baseline quality and high failure rate, we did not conduct multi-round evolutionary experiments for this track.
+However, as agent capabilities in coding and scientific modeling improve, the ML track will become a critical frontier for evaluating self-evolving intelligence in CATArena.
+Appendix M Agent performance on Hard Games: Pommerman
+(a) Start Status (step=0)
+(b) End Status (step=800)
+Figure 9: One Pommerman round demo. All agents failed to generate functional strategies, resulting in aimless movement or inactivity.
+Beyond the fundamental board games discussed in previous sections, we further investigate more sophisticated competitive environments that necessitate the ability to perceive complex scenarios and synthesize highly strategic code.
+Specifically, we utilize the Pommerman environment (Resnick et al., 2018), which was originally designed for multi-agent AI research and featured in the NeurIPS 2018 Competition Track [^3] https://neurips.cc/Conferences/2018/CompetitionTrack.
+Inspired by the classic Nintendo game Bomberman, Pommerman is played on a randomly generated symmetric $11\times 11$ grid.
+The environment contains both rigid walls (indestructible and impassable) and wooden walls (destructible by bombs).
+Each agent can execute one of six actions per turn: Stop, Up, Left, Down, Right, and Bomb, with the ultimate objective of being the sole survivor.
+Zero-Shot Performance Failure $S_{\text{base}}$
+The environment’s infrastructure comprises approximately 4,500 lines of Python code, while a proficient human-authored solution typically requires around 1,000 lines of strategic logic to achieve plausible gameplay.
+We task the agents with developing autonomous AI strategies for Pommerman; however, almost all agent-generated strategies proved ineffective during live evaluation.
+Our observations indicate that these results typically fall into three failure paradigms:
+  • Random Walk: The agent moves aimlessly across the field without ever deploying bombs.
+  • Self-Destruction: The agent successfully deploys bombs but fails to execute evasion maneuvers, resulting in self-elimination.
+  • Inactivity: The agent fails to produce valid operations within the turn limit, remaining stationary throughout the match.
+Code Analysis and Implications
+The generated code ranged from 100 to 400 lines, relying mostly on random actions or rudimentary Breadth-First Search (BFS).
+While Claude attempted to implement modular functions for tactical value estimation, its code was riddled with logical bugs that led to runtime errors or inactivity.
+These results highlight a critical cold start problem in complex environments: current LLMs cannot bridge the gap between understanding game rules and implementing a minimum viable strategy (MVP).Without a functional baseline, the evolutionary feedback loop cannot be initiated, validating our decision to exclude such high-complexity tasks from the main evolutionary experiments.
+Figure 9 illustrates a typical match where agents fail to engage meaningfully.
+The blue(Claude’s) and green(OpenAI’s) pommerman stay still, the red(Gemini’s) one find a more valuable place and then become inactive, the yellow(Qwen’s) pommerman walk in random.
+Appendix N Cost and Code Complexity of Participants
+We list the agents’ cost and code statistics in Table 14 for first round development of standard games,  15 for second round development of standard games, 16 for first round development of variant games and  17 for second round development of variant games.
+We can see that game development token costs show minimal variation, while differences are significant due to model changes.
+Claude (both minimal and code-based agents) consumes significantly more input tokens than competitors, exceeding the average by over 2 times, while Gemini generates notably more output tokens compared to other models.
+GPT-5 offers the best trade-off between cost and performance.
+Among all agents, second-round game development requires more input tokens, while output token growth remains marginal.
+In addition, commercial agents consistently use fewer tokens than their minimal-agent counterparts.
+In terms of code complexity, agents driven by Claude-4 model consistently surpass other agents in both the number of effective lines of code developed and the time spent considering development strategies.
+We observe that its development strategies are more sophisticated.
+Additionally, the complexity of its code increases with each iteration, which indirectly demonstrates the model’s exceptional learning capabilities.
+Table 14: Model cost statistics on standard tasks in Round 1.
 Metric Claude-4-Sonnet Deepseek-Chat Doubao-Seed Gemini-2.5-Pro GPT-5 Qwen3-Coder Claude-Code Codex Gemini-CLI Qwen-Coder
@@
 Avg thinking time (s) 1.3500 0.0000 0.0030 0.0030 0.2690 0.0020 1.9730 0.0030 0.0030 0.0050
-Average across games
-Total input tokens 1208689.625 36707.75 87115.25 464059.75 370266.5 239983.25 637555.875 - 342248.875 496871.125
-Session tokens 38442.75 10668.5 18443.0 41083.0 31757.25 19460.0 30771.125 - 30445.25 27170.375
-Total output tokens 17093.5 1193.25 13365.0 40106.0 17752.75 5707.0 11396.0 - 8021.375 11981.75
-Total time (s) 568.989 21.203 270.910 471.917 332.034 185.803 256.603 - 308.907 459.155
-Tools used 77.5 10.0 15.5 51.25 33.5 26.5 26.75 20.25 14.375 30.875
-Valid lines of code 642.75 22.5 354.5 199.0 362.25 305.0 332.25 327.5 253.0 357.25
-Avg thinking time (s) 0.3489 0.0012 0.7503 0.0044 0.5867 0.0260 0.4975 0.1792 0.0324 0.0050
-Table 16: Comparison of model cost on four games of 2nd round.
+OIBench
+Total input tokens 1101801 127528 1056516 8320878 193403 472132 1591886 - - -
+Session tokens 1117511 130209 1135372 8387950 216516 479643 1603116 - - -
+Total output tokens 15710 2681 78856 67072 23113 7511 11230 - - -
+Total time (s) 663.28 144.24 1982.12 3603.76 659.12 346.27 529.58 - - -
+Tools used 39 19 22 95 17 25 40 - - -
+Valid lines of code 312 149 242 375 167 273 381 - 548 135
+Avg thinking time (s) - - - - - - - - - -
+SWE-Perf
+Total input tokens 12566767 1233619 2867580 1109026 5945407 12246895 - - - -
+Session tokens 12622336 1240596 2977490 1155080 6049146 12308897 - - - -
+Total output tokens 55569 6977 109910 46054 103739 62002 - - - -
+Total time (s) 5028.39 2187.02 3284.32 2661.64 4082.8 4318.23 - - - -
+Tools used 123 23 31 14 76 138 - - - -
+Valid lines of code 257 0 31 100 125 133 - - - -
+Avg thinking time (s) - - - - - - - - - -
+Table 15: Model cost statistics on standard tasks in Round 2.
 Metric Claude-4-Sonnet Deepseek-Chat Doubao-Seed Gemini-2.5-Pro GPT-5 Qwen3-Coder Claude-Code Codex Gemini-CLI Qwen-Coder
@@
 Avg thinking time (s) 0.8240 0.0000 0.0020 0.0000 0.0030 0.0020 0.0030 0.0040 1.8470 0.0050
-Average across games
-Total input tokens 1570364.75 72464.75 297566.75 842630.5 992826.5 2566477.75 1743925.25 - 536505.75 445568.75
-Session tokens 93916.75 8146.5 43761.5 109111.0 86704.0 143595.25 50515.0 - 44458.25 43107.75
-Total output tokens 12182.5 1087.38 13830.88 21083.38 11294.13 8876.13 15005.63 - 8060.25 11583.75
-Total time (s) 622.63 334.81 346.27 865.02 415.34 1025.73 601.21 - 536.66 487.98
-Tools used 41.5 26.75 24.5 36.25 42.75 57.75 42.25 17.75 18.0 29.0
-Valid lines of code 549.25 22.5 418.25 269.75 408.25 362.0 548.25 334.75 298.25 534.5
-Avg thinking time (s) 0.7000 0.0013 0.5367 0.0029 0.4231 0.0911 0.1925 0.0585 0.4669 0.0057
-Table 17: Comparison of model cost on four variant games of 1st round.
+OIBench
+Total input tokens 689268 56536 734892 445268 331248 434818 1667173 - - -
+Session tokens 697796 57894 817803 491666 372662 441152 1674074 - - -
+Total output tokens 8528 1358 82911 46398 41414 6334 6901 - - -
+Total time (s) 499.51 68.95 2024.68 1868.17 1204.64 249.76 471.25 - - -
+Tools used 29 14 34 23 22 25 46 - - -
+Valid lines of code 278 0 398 271 339 306 419 - 780 420
+Avg thinking time (s) - - - - - - - - - -
+SWE-Perf
+Total input tokens 6720993 633967 1735119 435355 4530952 16977184 - - - -
+Session tokens 6738614 635214 1803480 506936 4593913 17070695 - - - -
+Total output tokens 17621 1247 68361 71581 62961 93511 - - - -
+Total time (s) 3939.86 263.16 2344.74 2377.89 3652.3 2546.19 - - - -
+Tools used 79 15 18 17 48 156 - - - -
+Valid lines of code 34 0 13 39 31 62 - - - -
+Avg thinking time (s) - - - - - - - - - -
+Table 16: Model cost statistics on variant tasks in Round 1.
 Metric Claude-4-Sonnet Deepseek-Chat Doubao-Seed Gemini-2.5-Pro GPT-5 Qwen3-Coder Claude-Code Codex Gemini-CLI Qwen-Coder
@@
 Avg thinking time (s) 0.7520 0.0000 0.0000 0.2560 1.1190 0.0020 1.9660 0.0040 0.0040 0.0050
-Average across games
-Total input tokens 873125.5 32763.25 94909.5 897345.25 267624.75 340063.0 886002.0 - 369823.25 235447.25
-Complete tokens 34555.25 9821.5 23031.75 31329.25 17193.5 12377.75 34906.0 - 31263.5 24084.0
-Total output tokens 14267.25 1154.5 9934.5 27090.5 13740.0 8697.5 11363.25 - 6911.75 7056.75
-Total time (s) 369.3838 18.7773 218.2230 839.8785 377.5413 1863.3930 274.1518 - 190.5898 227.2668
-Tools used 53.5 10.25 15.25 48.5 35.25 97.5 31.5 15.0 15.5 23.25
-Valid lines of code 668.5 31.5 410.0 280.75 302.75 237.5 325.0 341.0 188.0 322.5
-Avg thinking time (s) 0.2022 0.0012 0.0103 1.7729 0.3052 0.0036 0.4970 0.0245 0.0063 0.2025
-Table 18: Comparison of model cost on four variant games of 2nd round.
+OIBench
+Total input tokens 1360571 33086 7669553 339500 325721 721710 2035052 27754 - -
+Complete tokens 1380136 35679 7786790 406925 373777 732945 2052561 28256 - -
+Total output tokens 19565 2593 117237 67425 48056 11235 17509 502 - -
+Total time (s) 2433.550 80.090 3607.680 3600.980 1344.090 834.490 541.540 56.040 - -
+Tools used 44 8 127 27 23 34 46 7 - -
+Valid lines of code 224 82 421 248 276 271 248 0 405 371
+Avg thinking time (s) - - - - - - - - - -
+Table 17: Model cost statistics on variant tasks in Round 2.
 Metric Claude-4-Sonnet Deepseek-Chat Doubao-Seed Gemini-2.5-Pro GPT-5 Qwen3-Coder Claude-Code Codex Gemini-CLI Qwen-Coder
@@
 Avg thinking time (s) 1.6260 0.0000 0.0000 0.0440 0.7560 0.1780 0.0030 0.0040 1.7130 0.0050
-Average across games
-Total input tokens 1277124.75 53501.75 385533.5 599261.75 741507.75 1478472.25 1144616.25 - 597675.0 876130.5
-Complete tokens 93389.25 11325.25 70819.5 76283.5 57610.5 42925.75 43872.0 - 39733.25 39306.5
-Total output tokens 13538.5 1423.25 16999.0 28075.5 8760.0 14102.5 14683.0 - 8596.75 8802.0
-Total time (s) 421.4060 35.8547 363.4863 553.4572 511.2015 7458.3587 480.9153 - 266.3653 441.8328
-Tools used 49.5 15.5 21.5 49.5 54.0 209.75 37.75 20.75 19.0 33.0
-Valid lines of code 575.5 39.5 400.25 239.25 318.0 283.25 412.25 321.75 267.5 403.0
-Avg thinking time (s) 0.4141 0.0012 0.0247 0.0134 0.2137 0.0477 0.0258 0.0273 0.4330 0.0059
-We list the agents’ cost and code statistics in table 15 for first round development of standard games,  16 for second round development of standard games, 17 for first round development of vairant games and  18 for second round development of variant games.
-We can see that game development token costs show minimal variation, while differences is significant due to model changes.
-Claude (both minimal and code-based agents) consumes significantly more input tokens than competitors, exceeding the average by over 2 times, while Gemini generates notably more output tokens compared to other models.
-GPT-5 offers the best trade-off between cost.
-Among all agents, second-round game development require more input tokens, while output token growth remains marginal.
-In addition, commercial agents consistently use fewer tokens than their minimal-agent counterparts.
-In terms of code complexity, agents driven by Claude-4 model consistently surpasses other agents in both the number of effective lines of code developed and the time spent considering development strategies.
-We observe that its development strategies are more sophisticated.
-Additionally, the complexity of its code increases with each iteration, which indirectly demonstrates the model’s exceptional learning capabilities.
-Appendix K Full Prompts
+OIBench
+Total input tokens 1604531 40210 8364987 620122 665089 1081920 1027936 46854 - -
+Complete tokens 1610748 42112 8493081 653688 704924 1095605 1033256 78278 - -
+Total output tokens 6217 1902 128094 33566 39835 13685 5320 31424 - -
+Total time (s) 214.430 60.390 3604.330 3708.590 1235.390 411.730 259.790 810.060 - -
+Tools used 20 8 135 34 33 39 34 8 - -
+Valid lines of code 256 43 391 396 262 222 241 0 600 357
+Avg thinking time (s) - - - - - - - - - -
+Appendix O Full Prompts
 The agent is instructed to develop a competitive game AI based on the provided game environment.
 The AI must be deployed as an HTTP service with a single-port startup script, follow the official development instructions, and be named with the model prefix.
-And the agent is encouraged to iteratively improve its strategy based on the tournament report and the previous codes.
-Full prompt of details are in Table 19 for main leaderboard,  LABEL:tab:ml_game_ai_prompt for ML track and  21 for multilingual track.
+The agent is encouraged to iteratively improve its strategy based on the tournament report and the previous solutions.
+Full prompt details are in Tables 18 for the main leaderboard,  LABEL:tab:ml_game_ai_prompt for ML track and  20 for multilingual track.
 Table 18: Game AI Development Prompt
@@
 The final AI should be provided as an HTTP service.
-You can refer to the guides in game env path / README.md and game env path / develop instruction.md for development instructions.
-The content in game env path / develop instruction.md is very important, please read it carefully!
+You can refer to the guides in game env path / README.md and game env path / develop_instruction.md for development instructions.
+The content in game env path / develop_instruction.md is very important, please read it carefully!
 Please develop your AI service directly under dir path.
@@
 Access the main server
-You can play game of game env pathin at game server. You can play the games with your own AI or any other AI to improve your strategy. You can use bash tools to improve yourself.
-Final Remind
+You can play the game at game server using game env path. You can play the games with your own AI or any other AI to improve your strategy. You can use bash tools to improve yourself.
+Final Reminder
 You should write start_ai.sh in dir path and implement the AI service in dir path. DO NOT MODIFY THE CODE IN game env path
 Condition (if round_num>1round\_num>1):
-Tournament report of last round is in last round log dir.
+The tournament report from the last round is in last round log dir.
 The historical records are quite large, please use tools start_interactive_shell and run_interactive_shell to analyze the data efficiently.
@@
 Develop directly in dir path without repeatedly asking for the next step. Report to me only after you have completed the development.
-Final Remind
+Final Reminder
 You should write start_ai.sh and train_ai.sh in dir path and implement the REAL ML-based AI service with actual training capabilities in dir path.
@@
 Zero Tolerance for Pseudo-ML: Any submission without real training will be rejected.
-Table 21: Multi-language Game AI Development Prompt
+Table 20: Multilingual Game AI Development Prompt
 Game Development
@@
 The final AI should be provided as an HTTP service.
-You can refer to the guides in game env path / README.md and game env path / develop instruction.md for development instructions.
-The content in game env path / develop instruction.md is very important, please read it carefully!
+You can refer to the guides in game env path / README.md and game env path / develop_instruction.md for development instructions.
+The content in game env path / develop_instruction.md is very important, please read it carefully!
 Please develop your AI service directly under dir path.
@@
 Access the main server
-You can play game of game env pathin at game server. You can play the games with your own AI or any other AI to improve your strategy. You can use bash tools to improve yourself.
-Final Remind
+You can play the game at game server using game env path. You can play the games with your own AI or any other AI to improve your strategy. You can use bash tools to improve yourself.
+Final Reminder
 You should write start_ai.sh in dir path and implement the AI service in dir path. DO NOT MODIFY THE CODE IN game env path